### v1

# 3cm3cm3cm3cm

Abstract.
We study the joint action of the Bethe algebra $\mathcal{B}_{d}^{\bf w}$ for the general linear Lie algebra ${\mathfrak{gl}}_{d}$ with respect to ${\bf w}\in{\mathbb{C}}^{d}$ and the Bethe algebra ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ for the general linear Lie superalgebra ${\mathfrak{gl}}_{p+m|q+n}$ with respect to ${\bf z}\in{\mathbb{C}}^{p+q+m+n}$ on the Fock space of $d(p+m)$ bosonic and $d(q+n)$ fermionic oscillators. We establish a duality, called the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$, which is an equivalence between the actions of the algebras $\mathcal{B}_{d}^{\bf w}$ and ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on the Fock space. As an application, we show that the action of ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on each weight space of the evaluation module $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})$, where $\underline{M\mkern-4.0mu}\mkern 4.0mu$ is a $d$-fold tensor product of certain classes of infinite-dimensional unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-modules, is a cyclic ${\mathcal{B}_{p+m|q+n}^{\bf z}}$-module, and that ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ is diagonalizable with a simple spectrum on the weight space for generic ${\bf w}$ and ${\bf z}$.
1. Introduction
Let $\mathcal{F}:={\mathbb{C}}[{\bf x},{\bf y},{\boldsymbol{\eta}},{\boldsymbol{\zeta}}]$ be the polynomial superalgebra generated by $x^{a}_{i}$, $\eta^{a}_{j}$, $y^{a}_{r}$ and $\zeta^{a}_{s}$, for $i=1,\ldots,m$, $j=1,\ldots,n$, $r=1,\ldots,p$, $s=1,\ldots,q$ and $a=1,\ldots,d$. Here $x^{a}_{i}$ and $y^{a}_{r}$ are even variables while $\eta^{a}_{j}$ and $\zeta^{a}_{s}$ are odd. The superalgebra $\mathcal{F}$ can be realized as the Fock space of $d(p+m)$ bosonic and $d(q+n)$ fermionic oscillators. Let $\mathcal{D}:={\mathbb{D}}[{\bf x},{\bf y},{\boldsymbol{\eta}},{\boldsymbol{\zeta}}]$ be the corresponding Weyl superalgebra, i.e., $\mathcal{D}$ is the superalgebra generated by $x^{a}_{i}$, $\eta^{a}_{j}$, $y^{a}_{r}$ and $\zeta^{a}_{s}$ as well as their derivatives $\frac{\partial}{\partial x_{i}^{a}}$, $\frac{\partial}{\partial\eta_{j}^{a}}$, $\frac{\partial}{\partial y_{r}^{a}}$ and $\frac{\partial}{\partial\zeta_{r}^{a}}$. There is a natural action of $\mathcal{D}$ on $\mathcal{F}$, and hence $\mathcal{D}$ is a subalgebra of the endomorphism algebra ${\rm End}(\mathcal{F})$ of $\mathcal{F}$.
Moreover, there exist superalgebra homomorphisms $\phi:U({\mathfrak{gl}}_{d})\longrightarrow\mathcal{D}$ and $\overline{\phi}:U({\mathfrak{gl}}_{p+m|q+n})\longrightarrow\mathcal{D}$ from the universal enveloping algebras of ${\mathfrak{gl}}_{d}$ and ${\mathfrak{gl}}_{p+m|q+n}$ to $\mathcal{D}$ defined by (4.2) and (4.3), respectively. These maps induce a ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{p+m|q+n}$-action on $\mathcal{F}$. It is well-known that $\mathcal{F}$ decomposes into a direct sum of unitarizable ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{p+m|q+n}$-modules. This is called the Howe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ ([CLZ]). Such unitarizable modules are, in general, infinite-dimensional. Taking $p=q=0$, the maps $\phi$ and $\overline{\phi}$ specialize to the superalgebra homomorphisms $U({\mathfrak{gl}}_{d})\longrightarrow{\mathbb{D}}[{\bf x},{\boldsymbol{\eta}}]$ and $U({\mathfrak{gl}}_{m|n})\longrightarrow{\mathbb{D}}[{\bf x},{\boldsymbol{\eta}}]$, respectively, and the duality reduces to the Howe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m|n})$, which yields a decomposition of ${\mathbb{C}}[{\bf x},{\boldsymbol{\eta}}]$ into a direct sum of finite-dimensional ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{m|n}$-modules ([CW, Se1, Se2]). Assuming further that $n=0$, it gives the Howe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m})$ ([H]).
Fix ${\bf w}\in{\mathbb{C}}^{\ell}$ and ${\bf z}\in{\mathbb{C}}^{p+q+m+n}$. There are natural extensions of $\phi$ and $\overline{\phi}$ to the universal enveloping algebras of current algebras, and we have superalgebra homomorphisms $\phi_{\bf z}:U({\mathfrak{gl}}_{d}[t])\longrightarrow\mathcal{D}$ and $\overline{\phi}_{\bf w}:U({\mathfrak{gl}}_{p+m|q+n}[t])\longrightarrow\mathcal{D}$; see (4.4) and (4.5). It is natural to ask if dualities exist in the context of current algebras. When $p=q=n=0$, a duality phenomenon is noticed in [TL] (see also [TV]). In [MTV3], a precise statement of the duality between $U({\mathfrak{gl}}_{d}[t])$ and $U({\mathfrak{gl}}_{m}[t])$ is made and proved. So far, various generalizations have emerged (see [HM, TU, VY]). All of these are related to the Gaudin models for general linear Lie (super)algebras.
We will establish a duality between $U({\mathfrak{gl}}_{d}[t])$ and $U({\mathfrak{gl}}_{p+m|q+n}[t])$ in this paper. We consider the Bethe algebra $\mathcal{B}_{d}^{\bf w}$ for ${\mathfrak{gl}}_{d}$ with respect to ${\bf w}$ and the Bethe algebra ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ for ${\mathfrak{gl}}_{p+m|q+n}$ with respect to ${\bf z}$; see Section 3.3 and Section 3.4. As the algebras $\mathcal{B}_{d}^{\bf w}$ and ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ are respectively subalgebras of $U({\mathfrak{gl}}_{d}[t])$ and $U({\mathfrak{gl}}_{p+m|q+n}[t])$, the maps $\phi_{\bf z}$ and $\overline{\phi}_{\bf w}$ induce an action of $\mathcal{B}_{d}^{\bf w}$ and an action of ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on the polynomial superalgebra $\mathcal{F}$, respectively. We show that these two actions are equivalent. More precisely, we have:
Theorem 1.1 (Theorem 4.7).
$\phi_{\bf z}\big{(}\mathcal{B}_{d}^{\bf w}\big{)}=\overline{\phi}_{\bf w}\big{(}{\mathcal{B}_{p+m|q+n}^{\bf z}}\big{)}$.
We name Theorem 1.1 the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$. It recovers several dualities established earlier. After specializing to $p=q=0$, it recovers the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m|n})$ due to Huang and Mukhin [HM]. If we assume further that $n=0$ (resp., $m=0$), Theorem 1.1 gives the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m})$ obtained by Mukhin, Tarasov, and Varchenko [MTV3] (resp., a variant of the duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{n})$ obtained by Tarasov and Uvarov [TU]). Furthermore, we will show in [ChL3] that there is another duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ for Gaudin models with irregular singularities, which specializes to a variant of Theorem 1.1.
The Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ enables us to study the action of ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on $\mathcal{F}$ via the action of $\mathcal{B}_{d}^{\bf w}$ on $\mathcal{F}$, and vice versa. Using the properties of the action of $\mathcal{B}_{d}^{\bf w}$ on finite-dimensional ${\mathfrak{gl}}_{d}$-modules, we will better understand the action of ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on unitarizable ${\mathfrak{gl}}_{p+m|q+n}$-modules. Let $\underline{M\mkern-4.0mu}\mkern 4.0mu:=M_{1}\otimes\cdots\otimes M_{d}$, where each $M_{i}$ is an infinite-dimensional unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-module with highest weight corresponding to a generalized partition of depth 1 (see Section 5.2 for a detailed description). We will see in Theorem 5.1 that each weight space of the evaluation ${\mathfrak{gl}}_{p+m|q+n}[t]$-module $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})$ is a cyclic ${\mathcal{B}_{p+m|q+n}^{\bf z}}$-module, and that ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ is diagonalizable with a simple spectrum on the weight space for generic ${\bf w}$ and ${\bf z}$.
Let us explain a little how we prove Theorem 1.1. The Bethe algebras $\mathcal{B}_{d}^{\bf w}$ and ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ are determined by the column determinant ${\rm cdet}\,({\mathcal{L}}_{d}^{\bf w})$ and the Berezinian ${\rm Ber}^{\bf s}\,({\mathcal{L}}_{{p+m|q+n}}^{\bf z})$ of type ${\bf s}:=(0^{p},1^{q},0^{m},1^{n})$, respectively; see (3.6) and (3.9). In the special case where $p=q=0$, the specializations of the series $\phi_{\bf z}({\rm cdet}\,({\mathcal{L}}_{d}^{\bf w}))$ and $\overline{\phi}_{\bf w}({\rm Ber}^{\bf s}\,({\mathcal{L}}_{{p+m|q+n}}^{\bf z}))$ admit Capelli-type identities [HM, Proposition 5.4 and Proposition 5.6], and the duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m|n})$ is obtained by matching the identities. However, the situation is somewhat complicated in our case as no such identities are known to exist. Based on the properties of Berezinians, we can relate $\overline{\phi}_{\bf w}({\rm Ber}^{\bf s}\,({\mathcal{L}}_{{p+m|q+n}}^{\bf z}))$ to $\phi_{\bf z}({\rm cdet}\,({\mathcal{L}}_{d}^{\bf w}))$ (for general $p,q,m,n$) by examining an anti-involution on the Weyl superalgebra $\mathcal{D}$ and an anti-involution on the Bethe algebra $\mathcal{B}_{d}^{\bf w}$ (see Proposition 4.5). Theorem 1.1 will then follow.
We organize the paper as follows. In Section 2, we set up our notations. We recall the notions of the universal affine vertex algebra at the critical level, the Feigin–Frenkel center and the Bethe algebra. In Section 3, we discuss the Bethe algebras for general linear Lie (super)algebras and their basic properties. In Section 4, we introduce the joint action of the Bethe algebras $\mathcal{B}_{d}^{\bf w}$ and ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on the polynomial superalgebra $\mathcal{F}$ for any sequences ${\bf w}\in{\mathbb{C}}^{\ell}$ and ${\bf z}\in{\mathbb{C}}^{p+q+m+n}$. We prove Theorem 1.1, i.e., the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$. In Section 5, we apply the Bethe duality to the action of ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on unitarizable ${\mathfrak{gl}}_{p+m|q+n}$-modules and obtain Theorem 5.1.
Notations. Throughout the paper, the symbol ${\mathbb{Z}}$ (resp., ${\mathbb{N}}$ and ${\mathbb{Z}}_{+}$) stands for the set of all (resp., positive and non-negative) integers, the symbol ${\mathbb{C}}$ for the field of complex numbers, and the symbol ${\mathbb{Z}}_{2}:=\{\bar{0},\bar{1}\}$ for the field of integers modulo 2. All vector spaces, algebras, tensor products, etc., are over ${\mathbb{C}}$. We fix $d\in{\mathbb{N}}$ and $p,q,m,n\in{\mathbb{Z}}_{+}$.
2. Preliminaries
In this section, we set up our notations for the general linear Lie (super)algebras and their related concepts. We review several notions needed in this paper, including current algebras, the Feigin–Frenkel centers and the Bethe algebras.
2.1. The general linear Lie (super)algebras
For $p,q,m,n\in{\mathbb{Z}}_{+}$ that are not all zero, let
[…]
For $i\in\mathbb{I}$, define $|i|\in{\mathbb{Z}}_{2}$ by
$|i|=\begin{cases}\,\bar{0}&\,\,\mbox{if}\ \ i\in\{1,\ldots,p\}\cup\{p+q+1,\ldots,p+q+m\},\\ \,\bar{1}&\,\,\mbox{otherwise}.\end{cases}$
[…]
For any $i,j\in\mathbb{I}$, let $E^{i}_{j}$ denote the ${\mathbb{C}}$-linear endomorphism on ${\mathbb{C}}^{p|q}\oplus{\mathbb{C}}^{m|n}$ defined by
$E^{i}_{j}(e_{k})=\delta_{j,k}e_{i}\quad\mbox{ for $k\in\mathbb{I}$,}$
where $\delta$ is the Kronecker delta. The parity of $E^{i}_{j}$ is given by $|E^{i}_{j}|=|i|+|j|$. The superspace of ${\mathbb{C}}$-linear endomorphisms on ${\mathbb{C}}^{p|q}\oplus{\mathbb{C}}^{m|n}$ has a natural structure of a Lie superalgebra, called a general linear Lie (super)algebra and denoted by ${\mathfrak{gl}}_{p+m|q+n}$. The commutation relations of ${\mathfrak{gl}}_{p+m|q+n}$ are given by
$[E^{i}_{j},E^{k}_{\ell}]=\delta_{j,k}E^{i}_{\ell}-(-1)^{(|i|+|j|)(|k|+|\ell|)}\delta_{i,\ell}E^{k}_{j}\qquad\mbox{for $i,j,k,\ell\in\mathbb{I}$.}$
Note that $\{E^{i}_{j}\,|\,i,j\in\mathbb{I}\}$ is a homogeneous basis for ${\mathfrak{gl}}_{p+m|q+n}$.
Let $\mathfrak{b}_{{p+m|q+n}}=\bigoplus_{\begin{subarray}{c}i,j\in\mathbb{I},i\leq j\end{subarray}}{\mathbb{C}}E^{i}_{j}$ be a Borel subalgebra of ${\mathfrak{gl}}_{p+m|q+n}$. The corresponding Cartan subalgebra ${\mathfrak{h}}_{{p+m|q+n}}$ has a basis $\{E^{i}_{i}\,|\,i\in\mathbb{I}\}$, and the dual basis in ${\mathfrak{h}}_{{p+m|q+n}}^{*}$ is denoted by $\{\epsilon_{i}\,|\,i\in\mathbb{I}\}$, where the parity of $\epsilon_{i}$ is given by $|\epsilon_{i}|=|i|$. Note that the Borel subalgebra $\mathfrak{b}_{{p+m|q+n}}$ is not the standard one unless some of $p,q,m,n$ are set to 0.
For $m=d$ and $p=q=n=0$, we write
${\mathfrak{gl}}_{d}={\mathfrak{gl}}_{d|0},\quad{\mathfrak{b}}_{d}={\mathfrak{b}}_{d|0},\quad{\mathfrak{h}}_{d}={\mathfrak{h}}_{d|0},$
and
$\hskip 28.45274pte_{ij}=E^{i}_{j}\qquad\mbox{for $i,j=1,\ldots,d$.}$
2.2. Current algebras and evaluation modules
For any Lie (super)algebra ${\mathfrak{g}}$ and an even variable $t$, let ${\mathfrak{g}}[t,t^{-1}]:={\mathfrak{g}}\otimes{\mathbb{C}}[t,t^{-1}]$ denote the loop algebra of ${\mathfrak{g}}$, which is a Lie superalgebra with commutation relations
$\hskip 28.45274pt\big{[}A\otimes t^{r},B\otimes t^{s}\big{]}=[A,B]\otimes t^{r+s}\qquad\mbox{for $A,B\in{\mathfrak{g}}$ and $r,s\in{\mathbb{Z}}$.}$
Here $[A,B]$ is the supercommutator of $A$ and $B$. We will use the notation
$\hskip 28.45274ptA[r]:=A\otimes t^{r}\qquad\mbox{for $A\in{\mathfrak{g}}$ and $r\in{\mathbb{Z}}$.}$
The current algebra ${\mathfrak{g}}[t]:={\mathfrak{g}}\otimes{\mathbb{C}}[t]$ is a subalgebra of ${\mathfrak{g}}[t,t^{-1}]$ defined in an obvious way. We identify the Lie (super)algebra ${\mathfrak{g}}$ with the subalgebra ${\mathfrak{g}}\otimes 1$ of constant polynomials in ${\mathfrak{g}}[t]$ and hence $U({\mathfrak{g}})\subseteq U({\mathfrak{g}}[t])$. Hereafter $U(\mathcal{G})$ denotes the universal enveloping algebra of $\mathcal{G}$ for any Lie (super)algebra $\mathcal{G}$.
For $\alpha\in{\mathbb{C}}$, we have an evaluation homomorphism ${\rm ev}_{\alpha}:U({\mathfrak{g}}[t])\longrightarrow U({\mathfrak{g}})$ given by
${\rm ev}_{\alpha}(A[r])=\alpha^{r}A\qquad\mbox{for $A\in{\mathfrak{g}}$ and $r\in{\mathbb{Z}}_{+}$}.$
Any ${\mathfrak{g}}$-module $M$ is a ${\mathfrak{g}}[t]$-module via ${\rm ev}_{\alpha}$, which is called the evaluation module and denoted by $M(\alpha)$. For any $A\in{\mathfrak{g}}$ and any even variable $z$, we set
$A(z)=\sum_{r=0}^{\infty}A[r]z^{-r-1}.$
It is a formal power series in $z^{-1}$ with coefficients in ${\mathfrak{g}}[t]$ and acts on $M(\alpha)$ by
$A(z)v=\frac{Av}{z-\alpha}\qquad\mbox{for $v\in M(\alpha)$.}$
Here every rational function in $z$ represents its power series expression at $\infty$.
Let $\ell\in{\mathbb{N}}$ and ${\bf z}=(z_{1},\ldots,z_{\ell})\in{\mathbb{C}}^{\ell}$. Let $\Delta^{(\ell-1)}:U({\mathfrak{g}}[t])\longrightarrow U({\mathfrak{g}}[t])^{\otimes\ell}$ be the $(\ell-1)$-fold coproduct on $U({\mathfrak{g}}[t])$. We have the homomorphism
$\underline{{\rm ev}}_{\bf z}:=({\rm ev}_{z_{1}}\otimes\ldots\otimes{\rm ev}_{z_{\ell}})\circ\Delta^{(\ell-1)}:U({\mathfrak{g}}[t])\longrightarrow U({\mathfrak{g}})^{\otimes{\ell}},$
given by
$\underline{{\rm ev}}_{\bf z}(A[r])=\sum_{i=1}^{\ell}z_{i}^{r}A^{(i)}\qquad\mbox{for $A\in{\mathfrak{g}}$ and $r\in{\mathbb{Z}}_{+}$}.$
Here $A^{(i)}:=\underbrace{1\otimes\cdots\otimes 1\otimes\stackrel{{\scriptstyle i}}{{A}}\otimes 1\otimes\cdots\otimes 1}_{\ell}$ for $i=1,\ldots,\ell$. We call $\underline{{\rm ev}}_{\bf z}$ the evaluation map at ${\bf z}$.
For any ${\mathfrak{g}}$-modules $M_{1},\ldots,M_{\ell}$, let
$\underline{M\mkern-4.0mu}\mkern 4.0mu=M_{1}\otimes\cdots\otimes M_{\ell}.$ (2.1)
The tensor product $\underline{M\mkern-4.0mu}\mkern 4.0mu$ is a ${\mathfrak{g}}[t]$-module via $\underline{{\rm ev}}_{\bf z}$, which coincides with the tensor product of evaluation ${\mathfrak{g}}[t]$-modules
$\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z}):=M_{1}(z_{1})\otimes\cdots\otimes M_{\ell}(z_{\ell}).$ (2.2)
As ${\mathfrak{g}}$-modules, $\underline{M\mkern-4.0mu}\mkern 4.0mu$ and $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})$ are isomorphic via the identity map. For any $A\in{\mathfrak{g}}$, the series $A(z)$ acts on $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})$ by
$\hskip 42.67912ptA(z)v=\sum_{i=1}^{\ell}\frac{A^{(i)}v}{z-z_{i}}\qquad\mbox{for $v\in\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})$}.$
2.3. The Feigin–Frenkel center
The Lie algebra ${\mathfrak{gl}}_{d}$ admits a nondegenerate invariant symmetric bilinear form
$(A,B)={\rm tr}(AB)-\frac{1}{d}{\rm tr}(A){\rm tr}(B),\quad\mbox{for $A,B\in{\mathfrak{gl}}_{d}$},$
where ${\rm tr}$ denotes the trace. The affine Lie superalgebra
$\widehat{{\mathfrak{gl}}_{d}}:={\mathfrak{gl}}_{d}[t,t^{-1}]\oplus{\mathbb{C}}K,$
where $K$ is even and central, is a Lie superalgebra with commutation relations
$\big{[}A[r],B[s]\big{]}=[A,B][r+s]+r\delta_{r+s,0}(A,B)K$
for $A,B\in{\mathfrak{gl}}_{d}$ and $r,s\in{\mathbb{Z}}$.
The vacuum module $V_{{\rm crit}}({\mathfrak{gl}}_{d})$ at the critical level is
$V_{{\rm crit}}({\mathfrak{gl}}_{d}):=U(\widehat{{\mathfrak{gl}}_{d}})/I,$
where $I$ is the left ideal of $U(\widehat{{\mathfrak{gl}}_{d}})$ generated by ${\mathfrak{gl}}_{d}[t]$ and $K+d$. It possesses a unique vertex algebra structure, called the universal affine vertex algebra associated to ${\mathfrak{gl}}_{d}$ at the critical level; see [FBZ, Theorem 4.4.1]. The center of the vertex algebra $V_{{\rm crit}}({\mathfrak{gl}}_{d})$ is given by
${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d}):=\left\{\,\!v\in V_{{\rm crit}}({\mathfrak{gl}}_{d})\;\middle|\;{\mathfrak{gl}}_{d}[t]v=0\!\,\right\},$
called the Feigin–Frenkel center. Any element of ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d})$ is called a Segal–Sugawara vector.
We also have the universal affine vertex algebra $V_{{\rm crit}}({\mathfrak{gl}}_{p+m|q+n})$ at the critical level and the Feigin–Frenkel center ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{p+m|q+n})$ for ${\mathfrak{gl}}_{p+m|q+n}$. We skip details here; see [MR, Section 1.3] (and also [ChL2, Section 3.1]).
Let ${\mathfrak{g}}$ be a general linear Lie (super)algebra. The algebra $U(t^{-1}{\mathfrak{g}}[t^{-1}])$ is equipped with the (even) derivation ${\rm T}:=-d/dt$ defined by
${\rm T}(1)=0\qquad\text{and}\qquad{\rm T}(A[-r])=rA[-r-1]$
for $A\in{\mathfrak{g}}$ and $r\in{\mathbb{N}}$. Note that $U(t^{-1}{\mathfrak{g}}[t^{-1}])$ is isomorphic to $V_{{\rm crit}}({\mathfrak{g}})$ as superspaces by the Poincaré–Birkhoff–Witt theorem, and that the derivation ${\rm T}$ corresponds to the translator operator on $V_{{\rm crit}}({\mathfrak{g}})$. It is known that ${\mathfrak{z}}(\widehat{{\mathfrak{g}}})$ can be viewed as a commutative subalgebra of $U(t^{-1}{\mathfrak{g}}[t^{-1}])$ and is ${\rm T}$-invariant. We refer the reader to [Fr, FBZ, Mo, MR] for further details.
2.4. Bethe algebras
For $\mu\in{\mathfrak{gl}}_{d}^{*}$, there is an algebra homomorphism
$\Psi^{\mu}:U(t^{-1}{\mathfrak{gl}}_{d}[t^{-1}])\longrightarrow U({\mathfrak{gl}}_{d}[t])[\kern-1.1pt[z^{-1}]\kern-1.1pt]$
given by
$\hskip 28.45274pt\Psi^{\mu}(A[-r])=A\otimes(t-z)^{-r}+\delta_{r,1}\mu(A),\quad\mbox{for $A\in{\mathfrak{gl}}_{d}\,$ and $\,r\in{\mathbb{N}}$.}$ (2.3)
The Bethe algebra $\mathcal{B}_{d}^{\mu}$ of ${\mathfrak{gl}}_{d}$ with respect to $\mu$ is defined to be the subalgebra of $U({\mathfrak{gl}}_{d}[t])$ generated by the coefficients of $\Psi^{\mu}(S)$ for $S\in{\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d})$. Since ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d})$ is commutative, so is $\mathcal{B}_{d}^{\mu}$ ([FFTL, Ry]).
For any $\mathcal{B}_{d}^{\mu}$-module $V$, let $(\mathcal{B}_{d}^{\mu})_{V}$ denote the image of the Bethe algebra $\mathcal{B}_{d}^{\mu}$ in the endomorphism algebra ${\rm End}(V)$ of $V$. We call $(\mathcal{B}_{d}^{\mu})_{V}$ the Bethe algebra of $V$ with respect to $\mu$. For $\alpha\in{\mathbb{C}}$, any ${\mathfrak{gl}}_{d}$-module $M$ is the evaluation module $M(\alpha)$ over ${\mathfrak{gl}}_{d}[t]$, and so $\mathcal{B}_{d}^{\mu}$ acts on $M(\alpha)$. If the centralizer ${\mathfrak{gl}}_{d}^{\mu}$ of $\mu$ in ${\mathfrak{gl}}_{d}$ contains the Cartan subalgebra ${\mathfrak{h}}_{d}$, then $\mathcal{B}_{d}^{\mu}$ commutes with ${\mathfrak{h}}_{d}$ and hence acts on any $\gamma$-weight space $M(\alpha)_{\gamma}$ of $M(\alpha)$, where $\gamma\in{\mathfrak{h}}_{d}^{*}$ ([Ry, Proposition 4]).
Let
${\boldsymbol{X}}_{\!\ell}=\left\{\,\!(z_{1},\ldots,z_{\ell})\in{\mathbb{C}}^{\ell}\;\middle|\;z_{i}\not=z_{j}\,\,\mbox{for any $i\not=j$}\!\,\right\}$
be the configuration space of $\ell$ distinct points on ${\mathbb{C}}^{\ell}$, and let ${\bf z}\in{\boldsymbol{X}}_{\!\ell}$. Let $V_{1},\ldots,V_{\ell}$ be finite-dimensional irreducible ${\mathfrak{gl}}_{d}$-modules. The Bethe algebra $\mathcal{B}_{d}^{\mu}\subseteq U({\mathfrak{gl}}_{d}[t])$ acts on $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$ (for the notation, see (2.2)) via the evaluation map $\underline{{\rm ev}}_{\bf z}$.
An element $\mu\in{\mathfrak{gl}}_{d}^{*}$ is called regular if the dimension of ${\mathfrak{gl}}_{d}^{\mu}$ is $d$. The following remarkable results for $\mathcal{B}_{d}^{\mu}$, due to Feigin, Frenkel, and Rybnikov, will be used in this paper.
Theorem 2.1 (cf. [FFRy, Corollary 5]).
For any regular $\mu\in{\mathfrak{gl}}_{d}^{*}$ and ${\bf z}\in{\boldsymbol{X}}_{\!\ell}$, $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$ is a cyclic $\mathcal{B}_{d}^{\mu}$-module.
Theorem 2.2 (cf. [FFRy, Corollary 6]).
For generic $\mu\in{\mathfrak{gl}}_{d}^{*}$ and generic ${\bf z}\in{\boldsymbol{X}}_{\!\ell}$, the Bethe algebra $(\mathcal{B}_{d}^{\mu})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is diagonalizable with a simple spectrum.
3. Bethe algebras for general linear Lie (super)algebras
In this section, we recall column determinants, Berezinians and pseudo-differential operators. We discuss the Bethe algebras of the general linear Lie (super)algebras and their properties.
3.1. Column determinants and Berezinians
Let $\mathcal{A}$ be an associative unital superalgebra over ${\mathbb{C}}$. The parity of a homogeneous element $a\in\mathcal{A}$ is denoted by $|a|$, which lies in ${\mathbb{Z}}_{2}$. Fix $k\in{\mathbb{N}}$. For any $k\times k$ matrix $A=\big{[}a_{i,j}\big{]}_{i,j=1,\ldots,k}$ over $\mathcal{A}$, the column determinant of $A$ is defined to be
${\rm cdet}\,A=\sum_{\sigma\in\mathfrak{S}_{k}}(-1)^{l(\sigma)}\,a_{\sigma(1),1}\ldots a_{\sigma(k),k}.$
Here $\mathfrak{S}_{k}$ denotes the symmetric group on $\{1,\ldots,k\}$ and $l(\sigma)$ denotes the length of $\sigma$. For any nonempty subset $P=\{i_{1}<\ldots<i_{\ell}\}$ of $\{1,\ldots,k\}$, the matrix $A_{P}:=\big{[}a_{i,j}\big{]}_{i,j\in P}$ is called a standard submatrix of $A$.
Assume that $A$ has a two-sided inverse $A^{-1}=\big{[}\widetilde{a}_{i,j}\big{]}$. For all $i,j=1,\ldots,k$, the $(i,j)$th quasideterminant of $A$ is defined to be $|A|_{ij}:=\widetilde{a}_{j,i}^{-1}$ provided that $\widetilde{a}_{j,i}$ has an inverse in $\mathcal{A}$. According to [GGRW, Proposition 1.2.6], if $k\geq 2$, then
$|A|_{ij}=a_{i,j}-r_{i}^{j}(A^{ij})^{-1}c_{j}^{i},$
where $A^{ij}$ is the $(k-1)\times(k-1)$ submatrix of $A$ obtained by deleting the $i$th row and the $j$th column of $A$ and is assumed to be invertible, $r_{i}^{j}$ is the $i$th row of $A$ with $a_{i,j}$ removed, and $c_{j}^{i}$ is the $j$th column of $A$ with $a_{i,j}$ removed. Following the notation of [GGRW], it is convenient to write
$|A|_{ij}=\begin{vmatrix}a_{1,1}&\ldots&a_{1,j}&\ldots&a_{1,k}\\ \ldots&\ldots&\ldots&\ldots&\ldots\\ a_{i,1}&\ldots&\framebox{$a_{i,j}$}&\ldots&a_{i,k}\\ \ldots&\ldots&\ldots&\ldots&\ldots\\ a_{k,1}&\ldots&a_{k,j}&\ldots&a_{k,k}\end{vmatrix}.$
For $i=1,\ldots,k$, we define
$d_{i}(A)=\begin{vmatrix}a_{1,1}&\ldots&a_{1,i}\\ \ldots&\ldots&\ldots\\ a_{i,1}&\ldots&\framebox{$a_{i,i}$}\end{vmatrix},$
which are called the principal quasiminors of $A$.
A $k\times k$ matrix $A$ over $\mathcal{A}$ is called sufficiently invertible if every principal quasiminor of $A$ is well defined, and $A$ is called amply invertible if each of its standard submatrices is sufficiently invertible.
Let ${{\bf s}}=(s_{1},\ldots,s_{m+n})$ be a sequence of 0’s and 1’s such that exactly $m$ of the $s_{i}$’s are 0 and the others are 1. We call such a sequence a $0^{m}1^{n}$-sequence. Every $0^{m}1^{n}$-sequence can be written in the form $(0^{m_{1}},1^{n_{1}},\ldots,0^{m_{r}},1^{n_{r}}),$ where the sequence begins with $m_{1}$ copies of 00’s, followed by $n_{1}$ copies of $1$’s, and so on. The set of all $0^{m}1^{n}$-sequences is denoted by $\mathcal{S}_{{m|n}}$.
Let ${\bf s}\in\mathcal{S}_{{m|n}}$. For any $\sigma\in\mathfrak{S}_{m+n}$ and any $(m+n)\times(m+n)$ matrix $A=[a_{i,j}]$ over $\mathcal{A}$, we define ${\bf s}^{\sigma}=\left(s_{\sigma^{-1}(1)},s_{\sigma^{-1}(2)},\ldots,s_{\sigma^{-1}(m+n)}\right)$ and $A^{\sigma}=\big{[}a_{\sigma^{-1}(i),\sigma^{-1}(j)}\big{]}$. We say that $A$ is of type ${\bf s}$ if $a_{i,j}$ is a homogeneous element of parity $|a_{i,j}|=\bar{s}_{i}+\bar{s}_{j}$ for any $i,j=1,\ldots,m+n$.
For any $(m+n)\times(m+n)$ sufficiently invertible matrix $A$ of type ${\bf s}$ over $\mathcal{A}$, the Berezinian of type ${\bf s}$ of $A$ is defined to be
${\rm Ber}^{\bf s}\,A=d_{1}(A)^{\hat{s}_{1}}\ldots d_{m+n}(A)^{\hat{s}_{m+n}},$
where $\hat{s}_{i}:=(-1)^{s_{i}}$ for $1\leq i\leq m+n$ (see [HM, (3.3)]). We refer the reader to [Ber, Na, MR] for earlier definitions of Berezinians. We have the following proposition.
Proposition 3.1 (cf. [HM, Proposition 3.5]).
Let $A$ be an $(m+n)\times(m+n)$ amply invertible matrix of type ${\bf s}$ over $\mathcal{A}$. Fix $k\in\{1,\ldots,m+n-1\}$. We write
$A=\begin{bmatrix}W&X\\ Y&Z\end{bmatrix},$ (3.1)
where $W,X,Y,Z$ are respectively $k\times k$, $k\times(m+n-k)$, $(m+n-k)\times k$, and $(m+n-k)\times(m+n-k)$ matrices. Then $W$ and $Z-YW^{-1}X$ are sufficiently invertible matrices of types ${\bf s}^{\prime}:=(s_{1},\ldots,s_{k})$ and ${\bf s}^{\prime\prime}:=(s_{k+1},\ldots,s_{m+n})$, respectively. Moreover,
${\rm Ber}^{\bf s}\,A={\rm Ber}^{{\bf s}^{\prime}}W\cdot{\rm Ber}^{{\bf s}^{\prime\prime}}\big{(}Z-YW^{-1}X\big{)}.$
An $(m+n)\times(m+n)$ matrix $A=\big{[}a_{i,j}\big{]}$ over $\mathcal{A}$ is called a Manin matrix of type ${\bf s}$ if $A$ is a matrix of type ${\bf s}$ satisfying the following relations
$[a_{i,j},a_{k,l}]=(-1)^{s_{i}s_{j}+s_{i}s_{k}+s_{j}s_{k}}[a_{k,j},a_{i,l}]\qquad\text{for all}\quad i,j,k,l=1,\ldots,m+n,$
where $[a,b]:=ab-(-1)^{|a||b|}ba$ for any homogeneous elements $a,b\in\mathcal{A}$. We recall some basic facts about Manin matrices.
Proposition 3.2 (cf. [HM, Section 3]).
Let $A$ be an $(m+n)\times(m+n)$ Manin matrix of type ${\bf s}$ over $\mathcal{A}$. Then
  1. (i)
    $A^{-1}$ is an $(m+n)\times(m+n)$ Manin matrix of type ${\bf s}$ over $\mathcal{A}$.
  2. (ii)
    If $P=\{i_{1}<\ldots<i_{\ell}\}$ is a nonempty subset of $\{1,\ldots,m+n\}$, then the standard submatrix $A_{P}$ of $A$ is a Manin matrix of type ${{\bf s}}_{P}:=(s_{i_{1}},\ldots s_{i_{\ell}})$.
  3. (iii)
    For any $\sigma\in\mathfrak{S}_{m+n}$, $A^{\sigma}$ is a Manin matrix of type ${\bf s}^{\sigma}$.
The following proposition asserts that the Berezinian of a Manin matrix is invariant under the action of a permutation.
Proposition 3.3 ([HM, Proposition 3.6]).
Let $A$ be an $(m+n)\times(m+n)$ amply invertible Manin matrix of type ${\bf s}$ over $\mathcal{A}$. Then
$\hskip 28.45274pt{\rm Ber}^{{\bf s}^{\sigma}}A^{\sigma}={\rm Ber}^{\bf s}\,A\qquad\mbox{for any $\sigma\in\mathfrak{S}_{m+n}$.}$
We may decompose ${\rm Ber}^{\bf s}\,A$ in Proposition 3.1 in a different way if $A$ is a Manin matrix.
Proposition 3.4.
Let $A$ be an $(m+n)\times(m+n)$ amply invertible Manin matrix of type ${\bf s}$ over $\mathcal{A}$. Write $A$ as in (3.1). Then $W-XZ^{-1}Y$ and $Z$ are sufficiently invertible Manin matrices of types ${\bf s}^{\prime}:=(s_{1},\ldots,s_{k})$ and ${\bf s}^{\prime\prime}:=(s_{k+1},\ldots,s_{m+n})$, respectively. Moreover,
${\rm Ber}^{\bf s}\,A={\rm Ber}^{{\bf s}^{\prime\prime}}Z\cdot{\rm Ber}^{{\bf s}^{\prime}}\big{(}W-XZ^{-1}Y\big{)}.$ (3.2)
Proof.
The first assertion is clear from Proposition 3.2. Let $\sigma\in\mathfrak{S}_{m+n}$ be the permutation defined by $\sigma(k+i)=i$ for $i=1,\ldots,m+n-k$ and $\sigma(j)=m+n-k+j$ for $j=1,\ldots,k$. Then
$A^{\sigma}=\begin{bmatrix}Z&Y\\ X&W\end{bmatrix}.$
The equality (3.2) is now a direct consequence of Proposition 3.1 and Proposition 3.3. ∎
Proposition 3.4 will be used in Section 4.2. We also have the following proposition.
Proposition 3.5 ([CFR, Lemma 8]).
Let ${\bf s}=(0^{m})$. For any $m\times m$ sufficiently invertible Manin matrix $A$ of type ${\bf s}$ over $\mathcal{A}$, we have ${\rm Ber}^{\bf s}\,A={\rm cdet}\,\,A$.
3.2. Pseudo-differential operators
Let $\mathcal{A}$ be an associative unital superalgebra over ${\mathbb{C}}$ and $z$ an even variable. We denote by $\mathcal{A}[\kern-1.1pt[z]\kern-1.1pt]$ and $\mathcal{A}(\!(z)\!)$ the superalgebras of formal power series and Laurent series in $z$ with coefficients in $\mathcal{A}$, respectively. Let $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ denote the set of all formal series of the form
$\sum_{\ell=-\infty}^{s}\sum_{k=-\infty}^{r}a_{k\ell}z^{k}\partial_{z}^{\ell},$
where $r,s\in{\mathbb{Z}}$ and $a_{k\ell}\in\mathcal{A}$. We may endow $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ with a superalgebra structure using the rules:
$\partial_{z}\partial_{z}^{-1}=\partial_{z}^{-1}\partial_{z}=1,\quad\partial_{z}^{k}z^{\ell}=\sum_{i=0}^{\infty}\binom{k}{i}\binom{\ell}{i}i!\,z^{\ell-i}\,\partial_{z}^{k-i},\quad\mbox{for $k,\ell\in{\mathbb{Z}}$.}$ (3.3)
Here, e.g., ${\binom{k}{i}:=\frac{k(k-1)\ldots(k-i+1)}{i!}}$. The superalgebra $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ is called the superalgebra of pseudo-differential operators over $\mathcal{A}$.
Let $\omega$ be an anti-involution on $\mathcal{A}$. In particular, $\omega^{2}=1$ and $\omega(ab)=\omega(b)\omega(a)$ for $a,b\in\mathcal{A}$. We readily see that $\omega$ extends to a ${\mathbb{C}}$-linear automorphism $\omega:\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ defined by
$\omega(az^{k}\partial_{z}^{\ell})=\omega(a)z^{\ell}\partial_{z}^{k}$ (3.4)
for $a\in\mathcal{A}$ and $k,\ell\in{\mathbb{Z}}$.
Proposition 3.6.
The map $\omega$ is an anti-involution on $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$.
Proof.
Clearly, $\omega^{2}=1$. It remains to show that $\omega$ is an anti-homomorphism. For $a,b\in\mathcal{A}$ and $k,\ell,r,s\in{\mathbb{Z}}$, we have
$\omega((az^{k}\partial_{z}^{\ell})(bz^{r}\partial_{z}^{s}))$ $=$ $\omega(abz^{k}\partial_{z}^{\ell}z^{r}\partial_{z}^{s})$
$=$ $\sum_{i=0}^{\infty}\binom{\ell}{i}\binom{r}{i}i!\,\omega(abz^{k+r-i}\partial_{z}^{\ell+s-i})$
$=$ $\sum_{i=0}^{\infty}\binom{\ell}{i}\binom{r}{i}i!\,\omega(ab)z^{\ell+s-i}\partial_{z}^{k+r-i}$
$=$ $\omega(b)\omega(a)z^{s}\partial_{z}^{r}z^{\ell}\partial_{z}^{k}$
$=$ $\left(\omega(b)z^{s}\partial_{z}^{r}\right)\left(\omega(a)z^{\ell}\partial_{z}^{k}\right)$
$=$ $\omega(bz^{r}\partial_{z}^{s})\,\omega(az^{k}\partial_{z}^{\ell}),$
and the proposition follows. ∎
3.3. The Bethe algebra for ${\mathfrak{gl}}_{d}$ and Frobenius algebras
Let
${\bf w}=(w_{1},\ldots,w_{d})\in{\mathbb{C}}^{\ell}.$
Define $\mu_{\bf w}\in{\mathfrak{gl}}_{d}^{*}$ by
$\hskip 28.45274pt\mu_{\bf w}(e_{ab})=-\delta_{a,b}w_{a}\qquad\mbox{for $1\leq a\leq d$.}$
The choice of the minus sign will be clear below.
Recall the map $\Psi^{\mu_{\bf w}}$ defined in (2.3). We write
$\Psi^{\bf w}=\Psi^{\mu_{\bf w}}.$
Let $t$ be an even variable and
$\tau=-\partial_{t}.$
Here $t$ and $\partial_{t}$ satisfy the rules similar to (3.3). The map $\Psi^{\bf w}$ extends to an algebra homomorphism
$\Psi^{\bf w}:U(t^{-1}{\mathfrak{gl}}_{d}[t^{-1}])[\tau]\longrightarrow U({\mathfrak{gl}}_{d}[t])[\kern-1.1pt[z^{-1}]\kern-1.1pt][\partial_{z}]$
such that $\Psi^{\bf w}(\tau)=\partial_{z}$. Let
$\mathcal{T}_{d}=\Big{[}\delta_{a,b}\tau+e_{ab}[-1]\Big{]}_{a,b=1,\ldots,d},$
which is a $d\times d$ Manin matrix. We have
${\rm cdet}\,(\mathcal{T}_{d})=\tau^{d}+\sum_{i=1}^{d}S_{i}\tau^{i-1}.$ (3.5)
By [CM, Theorem 3.1] (see also [CT]), $S_{1},\ldots,S_{d}$ form a complete set of Segal–Sugawara vectors for the Feigin–Frenkel center ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d})$. That is, the set $\{{\rm T}^{r}S_{i}\,|\,i=1,\ldots,d,\,r\in{\mathbb{Z}}_{+}\}$ is algebraically independent, and
${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d})={\mathbb{C}}[{\rm T}^{r}S_{i}\,|\,i=1,\ldots,d,\,r\in{\mathbb{Z}}_{+}].$
Now consider the $d\times d$ Manin matrix
${\mathcal{L}}_{d}^{\bf w}:=\Big{[}\Psi^{\bf w}\big{(}\delta_{a,b}\tau+e_{ab}[-1]\big{)}\Big{]}_{a,b=1,\ldots,d}=\Big{[}\delta_{a,b}(\partial_{z}-w_{a})-e_{ab}(z)\Big{]}_{a,b=1,\ldots,d}.$
Then in view of (3.5),
${\rm cdet}\,({\mathcal{L}}_{d}^{\bf w})=\partial_{z}^{d}+\sum_{i=1}^{d}S_{i}^{{\bf w}}(z)\partial_{z}^{i-1},$ (3.6)
where $S_{i}^{{\bf w}}(z):=\Psi^{\bf w}(S_{i})$.
The following is well-known and matches the definition of the Bethe algebra in [MTV1].
Proposition 3.7.
The Bethe algebra $\mathcal{B}_{d}^{\bf w}$ is generated by the coefficients of $S_{i}^{{\bf w}}(z)$ for $i=1,\ldots,d$.
A Frobenius algebra $\mathcal{A}$ is a finite-dimensional commutative associative unital algebra which admits a nondegenerate symmetric bilinear form (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) such that
$\hskip 28.45274pt(a_{1}a_{2},a_{3})=(a_{1},a_{2}a_{3})\quad\mbox{for all $a_{1},a_{2},a_{3}\in\mathcal{A}$.}$
Recall the notations (2.1) and (2.2). Let $V_{1},\ldots,V_{\ell}$ be finite-dimensional irreducible ${\mathfrak{gl}}_{d}$-modules, and let ${\bf z}\in{\boldsymbol{X}}_{\!\ell}$. As ${\mathfrak{h}}_{d}\subseteq{\mathfrak{gl}}_{d}^{\mu_{\bf w}}$, the Bethe algebra $\mathcal{B}_{d}^{\bf w}$ acts on any weight space of $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$. We will explain why $(\mathcal{B}_{d}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is a Frobenius algebra for ${\bf w}\in{\boldsymbol{X}}_{\!d}$.
Define an anti-involution $\boldsymbol{\theta}:{\mathfrak{gl}}_{d}\longrightarrow{\mathfrak{gl}}_{d}$ by
$\boldsymbol{\theta}(e_{ab})=e_{ba}\qquad\mbox{for $1\leq a,b\leq d$.}$ (3.7)
The map $\boldsymbol{\theta}$ is called the Cartan anti-involution on ${\mathfrak{gl}}_{d}$.
For $i=1,\ldots,\ell$, $V_{i}$ is an irreducible highest weight ${\mathfrak{gl}}_{d}$-module with a dominant integral highest weight. Let $v_{i}$ be a highest weight vector of $V_{i}$. There is a unique nondegenerate symmetric bilinear form $(\cdot,\cdot)_{i}$ on $V_{i}$ defined by
$(v_{i},v_{i})_{i}=1\quad\text{and}\quad(Av,v^{\prime})_{i}=(v,\boldsymbol{\theta}(A)v^{\prime})_{i}$
for all $v,v^{\prime}\in V_{i}$ and $A\in{\mathfrak{gl}}_{d}$ ([Sh]). The bilinear form $(\cdot,\cdot)_{i}$ is called the Shapovalov form on $V_{i}$. It is easy to see that if $v$ and $v^{\prime}$ are weight vectors of $V_{i}$ of distinct weights, then $(v,v^{\prime})_{i}=0$.
Let (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) denote the tensor form of the bilinear forms $(\cdot,\cdot)_{1},\ldots,(\cdot,\cdot)_{\ell}$, which is called the tensor Shapovalov form. It is a nondegenerate symmetric bilinear form on $\underline{V\mkern-4.0mu}\mkern 4.0mu$. Let ${\bf w}\in{\mathbb{C}}^{d}$. By [MTV1, Theorem 9.1] and [Lu, Lemma 2.6], the Bethe algebra $\mathcal{B}_{d}^{\bf w}$ is symmetric with respect to (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ), i.e.,
$\hskip 42.67912pt({\boldsymbol{b}}v,v^{\prime})=(v,{\boldsymbol{b}}v^{\prime})\quad\mbox{for ${\boldsymbol{b}}\in\mathcal{B}_{d}^{\bf w}$ and $v,v^{\prime}\in\underline{V\mkern-4.0mu}\mkern 4.0mu$}.$
Theorem 3.8.
For any ${\bf w}\in{\boldsymbol{X}}_{\!d}$ and ${\bf z}\in{\boldsymbol{X}}_{\!\ell}$, the Bethe algebra $(\mathcal{B}_{d}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is a Frobenius algebra. Moreover, $(\mathcal{B}_{d}^{\bf w})_{{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}_{\gamma}}$ is a Frobenius algebra for any weight $\gamma$ of $\underline{V\mkern-4.0mu}\mkern 4.0mu$.
Proof.
The first statement follows from the above discussion on the tensor Shapovalov form (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) on $\underline{V\mkern-4.0mu}\mkern 4.0mu$ in view of Theorem 2.1 and [Lu, Lemma 2.7]. As (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) restricts to a nondegenerate symmetric bilinear form on ${\underline{V\mkern-4.0mu}\mkern 4.0mu}_{\!\gamma}$ for any weight $\gamma$ of $\underline{V\mkern-4.0mu}\mkern 4.0mu$, the second statement follows. ∎
The following is a consequence of [Lu, Lemma 1.3].
Theorem 3.9.
For any ${\bf w}\in{\boldsymbol{X}}_{\!d}$ and ${\bf z}\in{\boldsymbol{X}}_{\!\ell}$, the following properties hold:
  1. (i)
    The Bethe algebra $(\mathcal{B}_{d}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is a maximal commutative subalgebra of ${\rm End}(\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z}))$ of dimension $\dim\!\left({\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}\right)$.
  2. (ii)
    Every eigenspace of $(\mathcal{B}_{d}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is one-dimensional, and the set of eigenspaces of $(\mathcal{B}_{d}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is in bijective correspondence with the set of maximal ideals of $(\mathcal{B}_{d}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$.
  3. (iii)
    Every generalized eigenspace of $(\mathcal{B}_{d}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is a cyclic $\mathcal{B}_{d}^{\bf w}$-module.
Remark 3.10.
Each statement of Theorem 3.9 holds if $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$ is replaced with ${\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}_{\gamma}$ for any weight $\gamma$ of $\underline{V\mkern-4.0mu}\mkern 4.0mu$.
3.4. The Bethe algebra for ${\mathfrak{gl}}_{p+m|q+n}$
Fix ${\bf s}:=(0^{p},1^{q},0^{m},1^{n})\in\mathcal{S}_{p+m|q+n}$. Let
$\mathcal{T}_{p+m|q+n}=\Big{[}\delta_{i,j}\tau+(-1)^{|i|}E^{i}_{j}[-1]\Big{]}_{i,j\in\mathbb{I}}.$
It is straightforward to verify that $\mathcal{T}_{p+m|q+n}$ is an amply invertible $(p+q+m+n)\times(p+q+m+n)$ Manin matrix of type ${\bf s}$ over $U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])(\!(\tau^{-1})\!)$. The verification is analogous to [MR, Lemma 3.1]. The following is an immediate consequence of Proposition 3.3.
Proposition 3.11.
For each $\sigma\in\mathfrak{S}_{p+q+m+n}$,
${\rm Ber}\,^{{\bf s}^{\sigma}}\!\left(\mathcal{T}_{p+m|q+n}^{\sigma}\right)={\rm Ber}^{\bf s}\,\!\left(\mathcal{T}_{p+m|q+n}\right).$
We have the expansion
${\rm Ber}^{\bf s}\,\!\left(\mathcal{T}_{p+m|q+n}\right)=\sum_{i=-\infty}^{p+m-q-n}b_{i}\tau^{i},$ (3.8)
for some $b_{i}\in U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])$. According to [MR, Corollary 3.3] (see also [ChL2, Proposition 3.3]),
$b_{i}\in{\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{p+m|q+n})$
for all $i\in{\mathbb{Z}}$ with $i\leq p+m-q-n$.
Now let
${\bf z}=(z_{1},\ldots,z_{p+q+m+n})\in{\mathbb{C}}^{p+q+m+n}.$
Define $\mu_{\bf z}\in{\mathfrak{gl}}_{p+m|q+n}^{*}$ by
$\hskip 28.45274pt\mu_{\bf z}(E^{i}_{j})=(-1)^{|i|+1}\delta_{i,j}z_{i}\qquad\mbox{for $i,j\in\mathbb{I}$.}$
Similar to (2.3), we have a superalgebra homomorphism
$\Psi^{\bf z}:U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])\longrightarrow U({\mathfrak{gl}}_{p+m|q+n}[t])(\!(z^{-1})\!)$
given by
$\Psi^{\bf z}(A[-r])=A\otimes(t-z)^{-r}+\delta_{r,1}\mu_{\bf z}(A),\quad\mbox{for $A\in{\mathfrak{gl}}_{p+m|q+n}\,$ and $\,r\in{\mathbb{N}}$.}$
The map $\Psi^{\bf z}$ extends to a superalgebra homomorphism
$\Psi^{\bf z}:U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])(\!(\tau^{-1})\!)\longrightarrow U({\mathfrak{gl}}_{p+m|q+n}[t])(\!(z^{-1},\partial_{z}^{-1})\!)$
given by
$\Psi^{\bf z}\left(\sum_{i=-\infty}^{s}a_{i}\tau^{i}\right)=\sum_{i=-\infty}^{s}\Psi^{\bf z}(a_{i})\partial_{z}^{i}$
for $a_{i}\in U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])$ and $s\in{\mathbb{Z}}$.
Now consider the $(p+q+m+n)\times(p+q+m+n)$ matrix
${\mathcal{L}}_{{p+m|q+n}}^{\bf z}:=\Big{[}\Psi^{\bf z}\big{(}\delta_{i,j}\tau+(-1)^{|i|}E^{i}_{j}[-1]\big{)}\Big{]}_{i,j\in\mathbb{I}}=\Big{[}\delta_{i,j}(\partial_{z}-z_{i})-(-1)^{|i|}E^{i}_{j}(z)\Big{]}_{i,j\in\mathbb{I}}.$
Evidently, ${\mathcal{L}}_{{p+m|q+n}}^{\bf z}$ is an amply invertible Manin matrix of type ${\bf s}$ over the superalgebra $U({\mathfrak{gl}}_{p+m|q+n}[t])(\!(z^{-1},\partial_{z}^{-1})\!)$ of pseudo-differential operators over $U({\mathfrak{gl}}_{p+m|q+n}[t])$. Just like Proposition 3.11, we have the following.
Proposition 3.12.
For each $\sigma\in\mathfrak{S}_{p+q+m+n}$,
${\rm Ber}\,^{{\bf s}^{\sigma}}\big{(}\big{(}{\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big{)}^{\sigma}\big{)}={\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right).$
By (3.8), we have
${\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right)=\sum_{i=-\infty}^{p+m-q-n}b_{i}^{{\bf z}}(z)\partial_{z}^{i},$ (3.9)
where $b_{i}^{{\bf z}}(z):=\Psi^{\bf z}(b_{i})$. Let ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ be the subalgebra of $U({\mathfrak{gl}}_{p+m|q+n}[t])$ generated by the coefficients of the series $b_{i}^{{\bf z}}(z)$, for $i\in{\mathbb{Z}}$ with $i\leq p+m-q-n$. It is called the Bethe algebra of ${\mathfrak{gl}}_{p+m|q+n}$ with respect to ${\bf z}$. If we take $q=n=0$, we recover the Bethe algebra of ${\mathfrak{gl}}_{p+m}$ by virtue of Proposition 3.5. According to [MR, Corallary 3.7], the algebra ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ is a commutative subalgebra of $U({\mathfrak{gl}}_{p+m|q+n}[t])$. It also commutes with the Cartan subalgebra ${\mathfrak{h}}_{{p+m|q+n}}$ of ${\mathfrak{gl}}_{p+m|q+n}$ and hence acts on any weight space of a ${\mathfrak{gl}}_{p+m|q+n}$-module.
For any ${\mathcal{B}_{p+m|q+n}^{\bf z}}$-module $V$, let $\big{(}{\mathcal{B}_{p+m|q+n}^{\bf z}}\big{)}_{V}$ denote the image of the Bethe algebra ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ in ${\rm End}(V)$. We call $\big{(}{\mathcal{B}_{p+m|q+n}^{\bf z}}\big{)}_{V}$ the Bethe algebra of $V$ with respect to ${\bf z}$.
4. Duality for Bethe algebras
This section is devoted to establishing a duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ between the Bethe algebras $\mathcal{B}_{d}^{\bf w}\subseteq U({\mathfrak{gl}}_{d}[t])$ and ${\mathcal{B}_{p+m|q+n}^{\bf z}}\subseteq U({\mathfrak{gl}}_{p+m|q+n}[t])$ for any ${\bf w}\in{\mathbb{C}}^{d}$ and ${\bf z}\in{\mathbb{C}}^{p+q+m+n}$.
4.1. The actions of $U({\mathfrak{gl}}_{d}[t])$ and $U({\mathfrak{gl}}_{p+m|q+n}[t])$
Let $\mathcal{D}$ be the associative unital superalgebra generated by
$x^{a}_{i},\,\eta^{a}_{j},\,y^{a}_{r},\,\zeta^{a}_{s}$
as well as their derivatives
$\partial_{x^{a}_{i}}:=\frac{\partial}{\partial x_{i}^{a}},\quad\partial_{\eta^{a}_{j}}:=\frac{\partial}{\partial\eta_{j}^{a}},\quad\partial_{y^{a}_{r}}:=\frac{\partial}{\partial y_{r}^{a}},\quad\partial_{\zeta^{a}_{s}}:=\frac{\partial}{\partial\zeta_{r}^{a}}$
for $i=1,\ldots,m$, $j=1,\ldots,n$, $r=1,\ldots,p$, $s=1,\ldots,q$ and $a=1,\ldots,d$. Here $x^{a}_{i}$, $y^{a}_{r}$, $\partial_{x^{a}_{i}}$ and $\partial_{y^{a}_{r}}$ are even while $\eta^{a}_{j}$, $\zeta^{a}_{s}$, $\partial_{\eta^{a}_{j}}$, and $\partial_{\zeta^{a}_{s}}$ are odd. The superalgebra $\mathcal{D}$ is called the Weyl superalgebra.
There is an anti-involution $\omega:\mathcal{D}\longrightarrow\mathcal{D}$ defined by
${\begin{array}[]{lll}x^{a}_{i}\mapsto\partial_{x^{a}_{i}},&\quad\partial_{x^{a}_{i}}\mapsto x^{a}_{i},\quad\eta^{a}_{j}\mapsto\partial_{\eta^{a}_{j}},&\quad\partial_{\eta^{a}_{j}}\mapsto\eta^{a}_{j},\\ y^{a}_{r}\mapsto\partial_{y^{a}_{r}},&\quad\partial_{y^{a}_{r}}\mapsto y^{a}_{r},\quad\zeta^{a}_{s}\mapsto\partial_{\zeta^{a}_{s}},&\quad\partial_{\zeta^{a}_{s}}\mapsto\zeta^{a}_{s},\end{array}}$ (4.1)
for $1\leq i\leq m$, $1\leq j\leq n$, $1\leq r\leq p$, and $1\leq s\leq q$. In light of Proposition 3.6, the map $\omega$ extends to an anti-involution
$\omega:\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!).$
According to [CLZ, pp. 789–790], there is a superalgebra homomorphism $\phi:U({\mathfrak{gl}}_{d})\longrightarrow\mathcal{D}$, given by
$e_{ab}\mapsto\sum_{i=1}^{m}x_{i}^{a}\partial_{x^{b}_{i}}+\sum_{j=1}^{n}\eta_{j}^{a}\partial_{\eta^{b}_{j}}-\sum_{r=1}^{p}y_{r}^{b}\partial_{y^{a}_{r}}-\sum_{s=1}^{q}\zeta_{s}^{b}\partial_{\zeta^{a}_{s}},$ (4.2)
for $1\leq a,b\leq d$, and a superalgebra homomorphism $\overline{\phi}:U({\mathfrak{gl}}_{p+m|q+n})\longrightarrow\mathcal{D}$, given by
${\begin{array}[]{cc}E^{r}_{r^{\prime}}\mapsto-\sum_{a=1}^{d}\partial_{y^{a}_{r}}y_{r^{\prime}}^{a},&E^{r}_{p+s}\mapsto\sum_{a=1}^{d}\partial_{y^{a}_{r}}\zeta_{s}^{a},\\ E^{p+s}_{r}\mapsto-\sum_{a=1}^{d}\partial_{\zeta^{a}_{s}}y_{r}^{a},&E^{p+s}_{p+s^{\prime}}\mapsto\sum_{a=1}^{d}\partial_{\zeta^{a}_{s}}\zeta_{s^{\prime}}^{a},\\ E^{r}_{p+q+i}\mapsto\sum_{a=1}^{d}\partial_{y^{a}_{r}}\partial_{x^{a}_{i}},&E^{r}_{p+q+m+j}\mapsto\sum_{a=1}^{d}\partial_{y^{a}_{r}}\partial_{\eta^{a}_{j}},\\ E^{p+s}_{p+q+i}\mapsto\sum_{a=1}^{d}\partial_{\zeta^{a}_{s}}\partial_{x^{a}_{i}},&E^{p+s}_{p+q+m+j}\mapsto\sum_{a=1}^{d}\partial_{\zeta^{a}_{s}}\partial_{\eta^{a}_{j}},\\ E^{p+q+i}_{r}\mapsto-\sum_{a=1}^{d}x^{a}_{i}y_{r}^{a},&E^{p+q+i}_{p+s}\mapsto\sum_{a=1}^{d}x^{a}_{i}\zeta_{s}^{a},\\ E^{p+q+m+j}_{r}\mapsto-\sum_{a=1}^{d}\eta^{a}_{j}y_{r}^{a},&E^{p+q+m+j}_{p+s}\mapsto\sum_{a=1}^{d}\eta^{a}_{j}\zeta_{s}^{a},\\ E^{p+q+i}_{p+q+i^{\prime}}\mapsto\sum_{a=1}^{d}x_{i}^{a}\partial_{x^{a}_{i^{\prime}}},&E^{p+q+i}_{p+q+m+j}\mapsto\sum_{a=1}^{d}x_{i}^{a}\partial_{\eta^{a}_{j}},\\ E^{p+q+m+j}_{p+q+i}\mapsto\sum_{a=1}^{d}\eta_{j}^{a}\partial_{x^{a}_{i}},&E^{p+q+m+j}_{p+q+m+j^{\prime}}\mapsto\sum_{a=1}^{d}\eta_{j}^{a}\partial_{\eta^{a}_{j^{\prime}}}.\end{array}}$ (4.3)
where $1\leq i,i^{\prime}\leq m$, $1\leq j,j^{\prime}\leq n$, $1\leq r,r^{\prime}\leq p$, and $1\leq s,s^{\prime}\leq q$.
Let $\mathcal{F}$ be the polynomial superalgebra generated by $x^{a}_{i}$, $\eta^{a}_{j}$, $y^{a}_{r}$ and $\zeta^{a}_{s}$, for $1\leq i\leq m$, $1\leq j\leq n$, $1\leq r\leq p$, and $1\leq s\leq q$. With an appropriate inner product, $\mathcal{F}$ may be regarded as the Fock space of $d(p+m)$ bosonic and $d(q+n)$ fermionic oscillators. The Weyl superalgebra $\mathcal{D}$ acts naturally on $\mathcal{F}$, and so it is a subalgebra of ${\rm End}(\mathcal{F})$.
We simplify notations for the rest of this section. Let
$x^{a}_{m+j}=\eta^{a}_{j},\quad\partial_{x^{a}_{m+j}}=\partial_{\eta^{a}_{j}},\quad y^{a}_{p+s}=\xi^{a}_{s},\quad\partial_{y^{a}_{p+s}}=\partial_{\zeta^{a}_{s}},$
where $1\leq j\leq n$ and $1\leq s\leq q$. Fix two sequences
${\bf w}:=(w_{1},\ldots,w_{d})\in{\mathbb{C}}^{d}\qquad\text{and}\qquad{\bf z}:=(z_{1},\ldots,z_{p+q+m+n})\in{\mathbb{C}}^{p+q+m+n}.$
The map $\phi$ induces a superalgebra homomorphism $\phi_{\bf z}:U({\mathfrak{gl}}_{d}[t])\longrightarrow\mathcal{D}$ defined by
$e_{ab}[k]\mapsto\sum_{i=1}^{m+n}z_{p+q+i}^{k}x^{a}_{i}\partial_{x^{b}_{i}}-\sum_{r=1}^{p+q}z_{r}^{k}y^{b}_{r}\partial_{y^{a}_{r}},$ (4.4)
for $1\leq a,b\leq d$ and $k\in{\mathbb{Z}}_{+}$, and $\overline{\phi}$ induces a superalgebra homomorphism $\overline{\phi}_{\bf w}:U({\mathfrak{gl}}_{p+m|q+n}[t])\longrightarrow\mathcal{D}$ defined by
${\begin{array}[]{cll}E^{r}_{s}[k]&\mapsto\sum_{a=1}^{d}(-1)^{|s|+1}w_{a}^{k}\partial_{y^{a}_{r}}y^{a}_{s},\quad E^{r}_{p+q+j}[k]&\mapsto\sum_{a=1}^{d}w_{a}^{k}\partial_{y^{a}_{r}}\partial_{x^{a}_{j}},\\ E^{p+q+i}_{s}[k]&\mapsto\sum_{a=1}^{d}(-1)^{|s|+1}w_{a}^{k}x^{a}_{i}y^{a}_{s},\quad E^{p+q+i}_{p+q+j}[k]&\mapsto\sum_{a=1}^{d}w_{a}^{k}x^{a}_{i}\partial_{x^{a}_{j}},\end{array}}$ (4.5)
for $1\leq r,s\leq p+q$ and $1\leq i,j\leq m+n$.
The Bethe algebras $\mathcal{B}_{d}^{\bf w}$ and ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ are respectively commutative subalgebras of $U({\mathfrak{gl}}_{d}[t])$ and $U({\mathfrak{gl}}_{p+m|q+n}[t])$. Thus, the maps $\phi_{\bf z}$ and $\overline{\phi}_{\bf w}$ induce an action of $\mathcal{B}_{d}^{\bf w}$ and an action of ${\mathcal{B}_{p+m|q+n}^{\bf z}}$, respectively, on the polynomial superalgebra $\mathcal{F}$.
The Cartan anti-involution $\boldsymbol{\theta}$ on ${\mathfrak{gl}}_{d}$, defined in (3.7), gives rise to an anti-involution $\boldsymbol{\theta}:U({\mathfrak{gl}}_{d}[t])\longrightarrow U({\mathfrak{gl}}_{d}[t])$, defined by
$\hskip 42.67912pt\boldsymbol{\theta}(A[r])=\boldsymbol{\theta}(A)[r]\qquad\mbox{for $A\in{\mathfrak{gl}}_{d}$ and $r\in{\mathbb{Z}}_{+}$}.$
Proposition 4.1 ([MTV1, Proposition 8.4]).
The anti-involution $\boldsymbol{\theta}:U({\mathfrak{gl}}_{d}[t])\longrightarrow U({\mathfrak{gl}}_{d}[t])$ fixes each element of $\mathcal{B}_{d}^{\bf w}$.
Recall the anti-involution $\omega$ on $\mathcal{D}$ defined in (4.1). It is straightforward to verify that the diagram
[…]
commutes.
Proposition 4.2.
For any ${\boldsymbol{b}}\in\mathcal{B}_{d}^{\bf w}$, $\omega\big{(}\phi_{\bf z}({\boldsymbol{b}})\big{)}=\phi_{\bf z}({\boldsymbol{b}})$.
Proof.
This follows from (4.6) and Proposition 4.1. ∎
4.2. From Berezinians to column determinants
The map $\phi_{\bf z}$ and $\overline{\phi}_{\bf w}$, defined in (4.4) and (4.5), extend naturally to the superalgebra homomorphisms
$\phi_{\bf z}:U({\mathfrak{gl}}_{d}[t])(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!)$
and
$\overline{\phi}_{\bf w}:U({\mathfrak{gl}}_{p+m|q+n}[t])(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!),$
respectively.
Let
$\mathcal{C}=\left[\delta_{a,b}(z-w_{a})-\left(\sum_{i=1}^{m+n}\frac{x^{b}_{i}\partial_{x^{a}_{i}}}{\partial_{z}-z_{p+q+i}}-\sum_{r=1}^{p+q}\frac{y^{a}_{r}\partial_{y^{b}_{r}}}{\partial_{z}-z_{r}}\right)\right]_{a,b=1,\ldots,d}.$
Lemma 4.3.
We have
$\omega\big{(}\phi_{\bf z}{\rm cdet}\,({\mathcal{L}}_{d}^{\bf w})\big{)}={\rm cdet}\,(\mathcal{C}).$
Proof.
For $1\leq a,b\leq d$,
$\phi_{\bf z}\big{(}e_{ab}(z)\big{)}=\sum_{i=1}^{m+n}\frac{x^{a}_{i}\partial_{x^{b}_{i}}}{z-z_{p+q+i}}-\sum_{r=1}^{p+q}\frac{y^{b}_{r}\partial_{y^{a}_{r}}}{z-z_{r}}.$ (4.7)
We define
$F^{i}_{a,b}=\begin{cases}\,\,\delta_{a,b}(\partial_{z}-w_{a})&\mbox{if }\,i=0,\\ \displaystyle{\,\,\frac{y^{b}_{i}\partial_{y^{a}_{i}}}{z-z_{i}}}&\mbox{if }\,1\leq i\leq p+q,\\ \displaystyle{\,\frac{-x^{a}_{i-p-q}\partial_{x^{b}_{i-p-q}}}{z-z_{i}}}&\mbox{if }\,p+q+1\leq i\leq p+q+m+n,\end{cases}$
and
$G^{i}_{a,b}=\begin{cases}\,\,\delta_{a,b}(z-w_{a})&\mbox{if }\,i=0,\\ \displaystyle{\,\,\frac{y^{a}_{i}\partial_{y^{b}_{i}}}{\partial_{z}-z_{i}}}&\mbox{if }\,1\leq i\leq p+q,\\ \displaystyle{\,\frac{-x^{b}_{i-p-q}\partial_{x^{a}_{i-p-q}}}{\partial_{z}-z_{i}}}&\mbox{if }\,p+q+1\leq i\leq p+q+m+n.\end{cases}$
We have
$\omega(F^{i}_{a,b})=G^{i}_{a,b}$ (4.8)
for all $0\leq i\leq{p+q+m+n}$ and $1\leq a,b\leq d$.
Note that ${\mathcal{L}}_{d}^{\bf w}$ is an amply invertible Manin matrix of type $(0^{d})$. Let $\pi\in\mathfrak{S}_{d}$ be the permutation given by $\pi(i)=d+1-i$ for $i=1,\ldots,d$. Clearly, $\pi^{2}=1$. By (4.7), Proposition 3.5 and Proposition 3.12, we obtain
$\phi_{\bf z}{\rm cdet}\,({\mathcal{L}}_{d}^{\bf w})$ $=$ $\phi_{\bf z}{\rm cdet}\,\big{(}({\mathcal{L}}_{d}^{\bf w})^{\pi}\big{)}$
$=$ $\sum_{\sigma\in\mathfrak{S}_{d}}\sum_{i_{1},\ldots,i_{d}=0}^{p+q+m+n}(-1)^{l(\sigma)}F^{i_{d}}_{\pi(\sigma(1)),\pi(1)}\ldots F^{i_{1}}_{\pi(\sigma(d)),\pi(d)}$
$=$ $\sum_{\sigma\in\mathfrak{S}_{d}}\sum_{i_{1},\ldots,i_{d}=0}^{p+q+m+n}(-1)^{l(\sigma)}F^{i_{d}}_{\pi\sigma(1),d}\ldots F^{i_{1}}_{\pi\sigma(d),1}$
$=$ $\sum_{\sigma\in\mathfrak{S}_{d}}\sum_{i_{1},\ldots,i_{d}=0}^{p+q+m+n}(-1)^{l(\sigma)}F^{i_{d}}_{\sigma(d),d}\ldots F^{i_{1}}_{\sigma(1),1}.$
Here the last equality holds since $\pi^{-1}\mathfrak{S}_{d}\pi=\mathfrak{S}_{d}$ and $l(\pi^{-1}\sigma\pi)=l(\sigma)$ for all $\sigma\in\mathfrak{S}_{d}$. Thus,
$\omega\big{(}\phi_{\bf z}{\rm cdet}\,({\mathcal{L}}_{d}^{\bf w})\big{)}=\sum_{\sigma\in\mathfrak{S}_{d}}\sum_{i_{1},\ldots,i_{d}=0}^{p+q+m+n}(-1)^{l(\sigma)}\omega\!\left(F^{i_{1}}_{\sigma(1),1}\right)\ldots\omega\!\left(F^{i_{d}}_{\sigma(d),d}\right).$
In view of (4.8), we deduce that $\omega\big{(}{\rm cdet}\,\phi_{\bf z}\,{\mathcal{L}}_{d}^{\bf w}\big{)}={\rm cdet}\,(\mathcal{C})$. ∎
Fix ${\bf s}:=(0^{p},1^{q},0^{m},1^{n})\in\mathcal{S}_{p+m|q+n}$. For $u=z$ or $\partial_{z}$, let
$R_{p|q}(u)=\frac{(u-z_{1})\ldots(u-z_{p})}{(u-z_{p+1})\ldots(u-z_{p+q})}$
and
$R^{\prime}_{m|n}(u)=\frac{(u-z_{p+q+1})\ldots(u-z_{p+q+m})}{(u-z_{p+q+m+1})\ldots(u-z_{p+q+m+n})}.$
Proposition 4.4.
$(z-w_{1})\ldots(z-w_{d})\,\overline{\phi}_{\bf w}\,{\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right)=\omega\left(\!R^{\prime}_{m|n}(z)\cdot\phi_{\bf z}{\rm cdet}\,({\mathcal{L}}_{d}^{\bf w})\cdot R_{p|q}(z)\!\right).$
Proof.
For $1\leq r,s\leq p+q$ and $1\leq i,j\leq m+n$, we have
${\begin{array}[]{cll}\overline{\phi}_{\bf w}\big{(}E^{r}_{s}(z)\big{)}&=\sum_{a=1}^{d}\frac{(-1)^{|s|+1}\partial_{y^{a}_{r}}y^{a}_{s}}{z-w_{a}},\quad\overline{\phi}_{\bf w}\big{(}E^{r}_{p+q+j}(z)\big{)}&=\sum_{a=1}^{d}\frac{\partial_{y^{a}_{r}}\partial_{x^{a}_{j}}}{z-w_{a}},\\ \overline{\phi}_{\bf w}\big{(}E^{p+q+i}_{s}(z)\big{)}&=\sum_{a=1}^{d}\frac{(-1)^{|s|+1}x^{a}_{i}y^{a}_{s}}{z-w_{a}},\quad\overline{\phi}_{\bf w}\big{(}E^{p+q+i}_{p+q+j}(z)\big{)}&=\sum_{a=1}^{d}\frac{x^{a}_{i}\partial_{x^{a}_{j}}}{z-w_{a}}.\end{array}}$ (4.9)
Let
$D=\Big{[}\delta_{a,b}(z-w_{a})\Big{]}_{a,b=1,\ldots,d}\quad\text{and}\quad D^{\prime}=\Big{[}\delta_{i,j}(\partial_{z}-z_{i})\Big{]}_{i,j\in\mathbb{I}}.$
Consider the $(m+n)\times d$ matrices
$X:=\big{[}(-1)^{|i|}x^{a}_{i}\big{]}^{a=1,\ldots,d}_{i=1,\ldots,m+n}\quad\text{and}\quad P_{X}:=\big{[}\partial_{x^{a}_{i}}\big{]}^{a=1,\ldots,d}_{i=1,\ldots,m+n}$
and the $(p+q)\times d$ matrices
$Y:=\big{[}(-1)^{|r|+1}y^{a}_{r}\big{]}^{a=1,\ldots,d}_{r=1,\ldots,p+q}\qquad\text{and}\qquad P_{Y}:=\big{[}(-1)^{|r|}\partial_{y^{a}_{r}}\big{]}^{a=1,\ldots,d}_{r=1,\ldots,p+q}.$
Write
$T=\begin{bmatrix}Y^{t}\hskip 5.69054ptP_{X}^{t}\end{bmatrix}\quad\text{and}\quad T^{\prime}=\begin{bmatrix}P_{Y}\vspace{3mm}\\ X\end{bmatrix}.$
By (4.9), we get
$\overline{\phi}_{\bf w}\,{\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right)={\rm Ber}^{\bf s}\,(D^{\prime}-T^{\prime}D^{-1}T).$
Let
$\mathfrak{L}=\begin{bmatrix}D&T\vspace{2mm}\\ T^{\prime}&D^{\prime}\end{bmatrix}.$
By Proposition 3.1, we have
${\rm Ber}\,^{\!\hat{{\bf s}}}\,(\mathfrak{L})=(z-w_{1})\ldots(z-w_{d})\,\overline{\phi}_{\bf w}\,{\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right),$
where $\hat{{\bf s}}:=(0^{d},{\bf s})\in\mathcal{S}_{d+p+m|q+n}$. As $D-T(D^{\prime})^{-1}T^{\prime}$ is a sufficiently invertible Manin matrix of type $(0^{d})$, Proposition 3.4 and Proposition 3.5 imply that
${\rm Ber}\,^{\!\hat{{\bf s}}}\,(\mathfrak{L})$ $=$ ${\rm Ber}^{\bf s}\,(D^{\prime})\cdot{\rm cdet}\,(D-T(D^{\prime})^{-1}T^{\prime})$
$=$ $R_{p|q}(\partial_{z})\cdot R^{\prime}_{m|n}(\partial_{z})\cdot{\rm cdet}\,(\mathcal{C}^{\prime}),$
where
$\mathcal{C}^{\prime}:=\left[\delta_{a,b}(z-w_{a})-\left(\sum_{i=1}^{m+n}\frac{(-1)^{|i|}\partial_{x^{a}_{i}}x^{b}_{i}}{\partial_{z}-z_{p+q+i}}-\sum_{r=1}^{p+q}\frac{y^{a}_{r}\partial_{y^{b}_{r}}}{\partial_{z}-z_{r}}\right)\right]_{a,b=1,\ldots,d}.$
We claim that
$R^{\prime}_{m|n}(\partial_{z})\cdot{\rm cdet}\,(\mathcal{C}^{\prime})={\rm cdet}\,(\mathcal{C})\cdot R^{\prime}_{m|n}(\partial_{z}).$
This yields
${\rm Ber}\,^{\!\hat{{\bf s}}}\,(\mathfrak{L})=R_{p|q}(\partial_{z})\cdot{\rm cdet}\,(\mathcal{C})\cdot R^{\prime}_{m|n}(\partial_{z}),$
which proves the proposition by Lemma 4.3.
To prove the claim, it suffices to show that
$R^{\prime}_{m|n}(\partial_{z})\cdot\mathcal{C}^{\prime}_{a,b}=\mathcal{C}_{a,b}\cdot R^{\prime}_{m|n}(\partial_{z}),\qquad\mbox{for $1\leq a,b\leq d$.}$ (4.10)
Here, for instance, $\mathcal{C}_{a,b}$ denotes the $(a,b)$ entry of $\mathcal{C}$. Since
$\partial_{x^{a}_{i}}x^{b}_{i}=(-1)^{|i|}x^{b}_{i}\partial_{x^{a}_{i}}+\delta_{a,b},\qquad\mbox{for $1\leq a,b\leq d$},$ (4.11)
the off-diagonal entries of $\mathcal{C}^{\prime}$ coincide with those of $\mathcal{C}$. For $1\leq i\leq m+n$, the term $\partial_{z}-z_{p+q+i}$ commutes with $\partial_{x^{a}_{j}}x^{b}_{j}$ and $y^{a}_{r}\partial_{y^{b}_{r}}$ for all $1\leq j\leq m+n$, $1\leq r\leq p+q$ and $1\leq a,b\leq d$. In particular, $R^{\prime}_{m|n}(\partial_{z})$ commutes with any off-diagonal entry of $\mathcal{C}^{\prime}$. This proves (4.10) for $a\not=b$. Moreover, for $1\leq i\leq m+n$ and $1\leq a\leq d$,
$(\partial_{z}-z_{p+q+i})(z-w_{a})=(z-w_{a})(\partial_{z}-z_{p+q+i})+1.$ (4.12)
If $1\leq i\leq m$, we see from (4.11) and (4.12) that
${(\partial_{z}-z_{p+q+i})}\left(z-w_{a}-\frac{\partial_{x^{a}_{i}}x^{a}_{i}}{\partial_{z}-z_{p+q+i}}\right)=\left(z-w_{a}-\frac{x^{a}_{i}\partial_{x^{a}_{i}}}{\partial_{z}-z_{p+q+i}}\right)(\partial_{z}-z_{p+q+i}).$
Similarly, if $m+1\leq i\leq m+n$, we have
${(\partial_{z}-z_{p+q+i})^{-1}}\left(z-w_{a}+\frac{\partial_{x^{a}_{i}}x^{a}_{i}}{\partial_{z}-z_{p+q+i}}\right)=\left(z-w_{a}-\frac{x^{a}_{i}\partial_{x^{a}_{i}}}{\partial_{z}-z_{p+q+i}}\right)(\partial_{z}-z_{p+q+i})^{-1}.$
The above discussion proves (4.10) for $a=b$. Thus, the proof of (4.10) is now complete. ∎
We relate ${\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right)$ to ${\rm cdet}\,({\mathcal{L}}_{d}^{\bf w})$ as follows.
Proposition 4.5.
We may write
$R_{p|q}(z)\cdot R^{\prime}_{m|n}(z)\cdot\phi_{\bf z}{\rm cdet}\,({\mathcal{L}}_{d}^{\bf w})=\sum_{\ell=0}^{d}\sum_{k=-\infty}^{p+m-q-n}b_{k,\ell}^{{\bf w},{\bf z}}z^{k}\partial_{z}^{\ell},$
where $b_{k,\ell}^{{\bf w},{\bf z}}:=\phi_{\bf z}(b_{k,\ell}^{\bf w})$ for some $b_{k,\ell}^{\bf w}\in\mathcal{B}_{d}^{\bf w}$. Then
$(z-w_{1})\ldots(z-w_{d})\ \overline{\phi}_{\bf w}\,{\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right)$
$\hskip 85.35826pt=R_{p|q}(\partial_{z})\left(\sum_{k=-\infty}^{p+m-q-n}\sum_{\ell=0}^{d}b_{k,\ell}^{{\bf w},{\bf z}}z^{\ell}\partial_{z}^{k}\right)R_{p|q}(\partial_{z})^{-1}.$
Proof.
By Proposition 4.4,
$(z-w_{1})\ldots(z-w_{d})\,\overline{\phi}_{\bf w}{\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right)$
$\hskip 42.67912pt=R_{p|q}(\partial_{z})\cdot\omega\left(R^{\prime}_{m|n}(z)\cdot\phi_{\bf z}{\rm cdet}\,({\mathcal{L}}_{d}^{\bf w})\right)\!$
$\hskip 42.67912pt=R_{p|q}(\partial_{z})\cdot\omega\left(R_{p|q}(z)\cdot R^{\prime}_{m|n}(z)\cdot\phi_{\bf z}{\rm cdet}\,({\mathcal{L}}_{d}^{\bf w})\right)\cdot R_{p|q}(\partial_{z})^{-1}.$
Therefore,
$(z-w_{1})\ldots(z-w_{d})\ \overline{\phi}_{\bf w}\,{\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right)$
$\hskip 85.35826pt=R_{p|q}(\partial_{z})\left(\sum_{k=-\infty}^{p+m-q-n}\sum_{\ell=0}^{d}\omega\big{(}b_{k,\ell}^{{\bf w},{\bf z}}\big{)}z^{\ell}\partial_{z}^{k}\right)R_{p|q}(\partial_{z})^{-1}.$
By Proposition 4.2,
$\omega\big{(}b_{k,\ell}^{{\bf w},{\bf z}}\big{)}=b_{k,\ell}^{{\bf w},{\bf z}}$
for all $k\leq p+m-q-n$ and $0\leq\ell\leq d$. This completes the proof. ∎
4.3. Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$
Lemma 4.6.
Let $c\in\{z_{1},\ldots,z_{p+q}\}$. Given $\alpha_{r,s}^{{\bf w},{\bf z}}\in\mathcal{D}$ for $r,s\in{\mathbb{Z}}$ with $0\leq r\leq d$ and $s\leq p+m-q-n$. Let $\mathcal{D}_{1}$ be the subalgebra of $\mathcal{D}$ generated by all $\alpha_{r,s}^{{\bf w},{\bf z}}$’s, and let $\mathcal{D}_{2}$ be the subalgebra of $\mathcal{D}$ generated by the coefficients of the series
$(\partial_{z}-c)\left(\sum_{s=-\infty}^{p+m-q-n}\sum_{r=0}^{d}\alpha_{r,s}^{{\bf w},{\bf z}}z^{r}\partial_{z}^{s}\right)(\partial_{z}-c)^{-1}.$ (4.13)
Then $\mathcal{D}_{1}=\mathcal{D}_{2}$.
Proof.
For any $r\in{\mathbb{Z}}_{+}$,
$(\partial_{z}-c)z^{r}(\partial_{z}-c)^{-1}=z^{r}+rz^{r-1}(\partial_{z}-c)^{-1}.$
Let $f$ denote the series (4.13). We find that
$f=\sum_{s=-\infty}^{p+m-q-n}\sum_{r=0}^{d}\left(\alpha_{r,s}^{{\bf w},{\bf z}}z^{r}\partial_{z}^{s}+r\alpha_{r,s}^{{\bf w},{\bf z}}z^{r-1}\partial_{z}^{s}(\partial_{z}-c)^{-1}\right)\!,$
and hence
$f=\sum_{\ell=-\infty}^{p+m-q-n}\sum_{k=0}^{d}\beta_{k,\ell}^{{\bf w},{\bf z}}z^{k}\partial_{z}^{\ell},$
where
$\beta_{d,\ell}^{{\bf w},{\bf z}}=\alpha_{d,\ell}^{{\bf w},{\bf z}},$ (4.14)
$\beta_{k,\ell}^{{\bf w},{\bf z}}=\alpha_{k,\ell}^{{\bf w},{\bf z}}+\sum_{i=0}^{\infty}(k+1)\alpha_{k+1,\ell+i+1}^{{\bf w},{\bf z}}c^{i},$ (4.15)
for $10\leq k\leq d-$ 1 and $\ell\leq p+m-q-n$. Note that the right-hand side of (4.15) has only finitely many nonzero terms.
The algebra $\mathcal{D}_{2}$ is, by definition, generated by $\beta_{k,\ell}^{{\bf w},{\bf z}}$’s. So, it is a subalgebra of $\mathcal{D}_{1}$ by (4.14) and (4.15). Also, (4.14) means that $\alpha_{d,\ell}^{{\bf w},{\bf z}}\in\mathcal{D}_{2}$ for all $\ell\leq p+m-q-n$. Using (4.15), we see, by induction, that $\alpha_{r,\ell}^{{\bf w},{\bf z}}\in\mathcal{D}_{2}$ for all $0\leq r\leq d$ and $\ell\leq p+m-q-n$. It follows that $\mathcal{D}_{1}\subseteq\mathcal{D}_{2}$, and thus $\mathcal{D}_{1}=\mathcal{D}_{2}$. ∎
We have the following duality between the Bethe algebras $\mathcal{B}_{d}^{\bf w}$ and ${\mathcal{B}_{p+m|q+n}^{\bf z}}$. We call it the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$.
Theorem 4.7.
$\phi_{\bf z}(\mathcal{B}_{d}^{\bf w})=\overline{\phi}_{\bf w}\big{(}{\mathcal{B}_{p+m|q+n}^{\bf z}}\big{)}$.
Proof.
This follows from Proposition 4.5 and repeated use of Lemma 4.6. ∎
Remark 4.8.
We will demonstrate another duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ for Gaudin models with irregular singularities, which specializes to a variant of Theorem 4.7, in [ChL3]. The classical version of the duality will also be investigated there.
If some of $p,q,m,n$ are 0, we will suppress (most of) them in the symbol ${\mathcal{B}_{p+m|q+n}^{\bf z}}$. For instance,
$\mathcal{B}^{\bf z}_{p+m|n}:=\mathcal{B}^{\bf z}_{p+m|0+n},\quad\mathcal{B}^{\bf z}_{m|n}:=\mathcal{B}^{\bf z}_{0+m|0+n},\quad\mathcal{B}^{\bf z}_{0|n}:=\mathcal{B}^{\bf z}_{0+0|0+n},\quad\text{etc.}$. (4.16)
Theorem 4.7 gives the Bethe dualities of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|n})$ and $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m})$.
Corollary 4.9.
  1. (i)
    Suppose $q=0$. Then $\phi_{\bf z}(\mathcal{B}_{d}^{\bf w})=\overline{\phi}_{\bf w}\big{(}\mathcal{B}^{\bf z}_{p+m|n}\big{)}$.
  2. (ii)
    Suppose $q=n=0$. Then $\phi_{\bf z}(\mathcal{B}_{d}^{\bf w})=\overline{\phi}_{\bf w}\big{(}\mathcal{B}^{\bf z}_{p+m}\big{)}$.
Theorem 4.7 recovers the duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m|n})$ due to Huang and Mukhin as well as the duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m})$ due to Mukhin, Tarasov, and Varchenko.
Corollary 4.10 ([HM, Theorem 5.2]).
Suppose $p=q=0$. Then $\phi_{\bf z}(\mathcal{B}_{d}^{\bf w})=\overline{\phi}_{\bf w}\big{(}\mathcal{B}^{\bf z}_{m|n}\big{)}$.
Corollary 4.11 ([MTV3, Theorem 6]).
Suppose $p=q=n=0$. Then $\phi_{\bf z}(\mathcal{B}_{d}^{\bf w})=\overline{\phi}_{\bf w}\big{(}\mathcal{B}^{\bf z}_{m}\big{)}$.
We also obtain a variant of the duality of Bethe algebras established by Tarasov and Uvarov (cf. [TU, Theorem 5.2]).
Corollary 4.12.
Suppose $p=q=m=0$. Then $\phi_{\bf z}(\mathcal{B}_{d}^{\bf w})=\overline{\phi}_{\bf w}\big{(}\mathcal{B}^{\bf z}_{0|n}\big{)}$.
5. An application of Bethe duality
In this section, we recall *-structures and unitarizable modules. We apply the Bethe duality to the Bethe algebra of ${\mathfrak{gl}}_{p+m|q+n}$ acting on a tensor product of infinite-dimensional unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-modules.
5.1. *-structures and unitarizable modules
A *-superalgebra is an associative superalgebra $A$ with an even anti-linear anti-involution ${\boldsymbol{\sigma}}:A\longrightarrow A$. A homomorphism $f:(A,{\boldsymbol{\sigma}})\longrightarrow(A^{\prime},{\boldsymbol{\sigma}}^{\prime})$ of *-superalgebras is a homomorphism of superalgebras satisfying ${\boldsymbol{\sigma}}^{\prime}\circ f=f\circ{\boldsymbol{\sigma}}$.
[…]
$\langle av|v^{\prime}\rangle=\langle v|{\boldsymbol{\sigma}}(a)v^{\prime}\rangle,$
for all $a\in A$ and $v,v^{\prime}\in V$. An $A$-module equipped with a positive definite contravariant Hermitian form is called a unitarizable $A$-module (with respect to ${\boldsymbol{\sigma}}$).
Let ${\mathfrak{g}}$ be a Lie superalgebra. An anti-linear anti-involution ${\boldsymbol{\sigma}}:{\mathfrak{g}}\longrightarrow{\mathfrak{g}}$ is called a *-structure on ${\mathfrak{g}}$. A homomorphism $f:({\mathfrak{g}},{\boldsymbol{\sigma}})\longrightarrow({\mathfrak{g}}^{\prime},{\boldsymbol{\sigma}}^{\prime})$ of Lie superalgebras with *-structures is a homomorphism of Lie superalgebras satisfying ${\boldsymbol{\sigma}}^{\prime}\circ f=f\circ{\boldsymbol{\sigma}}$. Moreover, ${\boldsymbol{\sigma}}$ is a *-structure on ${\mathfrak{g}}$ if and only if the natural extension of ${\boldsymbol{\sigma}}$ to $U({\mathfrak{g}})$ is an anti-linear anti-involution.
[…]
The Lie superalgebra ${\mathfrak{gl}}_{p+m|q+n}$ admits a *-structure ${\boldsymbol{\sigma}}:{\mathfrak{gl}}_{p+m|q+n}\longrightarrow{\mathfrak{gl}}_{p+m|q+n}$ (see [CLZ, Section 3.2]), defined by
$\hskip 42.67912pt\sum_{i,j\in\mathbb{I}}a_{ij}E^{i}_{j}\mapsto\sum_{i,j\in\mathbb{I}}(-1)^{[i]+[j]}\bar{a}_{ij}E^{j}_{i}\qquad\mbox{for $1\leq i,j\leq p+q+m+n$},$
where
$[i]:=\begin{cases}\,0&\,\,\mbox{if}\ \ 1\leq i\leq p,\\ \,1&\,\,\mbox{otherwise},\end{cases}$
and $\bar{a}_{ij}$ denotes the complex conjugate of $a_{ij}\in{\mathbb{C}}$. Let $\pi:{\mathfrak{gl}}_{d}\longrightarrow{\mathfrak{gl}}_{d}$ be the *-structure on ${\mathfrak{gl}}_{d}$ defined by
$\pi(e_{ab})=e_{ba}\quad\mbox{for $1\leq a,b\leq d$.}$
The maps $\pi$ and ${\boldsymbol{\sigma}}$ give rise to a *-structure $(\pi,{\boldsymbol{\sigma}})$ on ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{p+m|q+n}$.
5.2. Bethe algebras of unitarizable ${\mathfrak{gl}}_{p+m|q+n}$-modules
The Weyl superalgebra $\mathcal{D}$ is a subalgebra of the endomorphism algebra ${\rm End}(\mathcal{F})$ of the polynomial superalgebra $\mathcal{F}$. For ${\bf w}\in{\mathbb{C}}^{d}$ and ${\bf z}\in{\mathbb{C}}^{p+q+m+n}$, Theorem 4.7 can be rephrased as saying that the image of $\mathcal{B}_{d}^{\bf w}$ in ${\rm End}(\mathcal{F})$ under $\phi_{\bf z}$ coincides with the image of ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ in ${\rm End}(\mathcal{F})$ under $\overline{\phi}_{\bf w}$. This makes it possible to study the action of ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on $\mathcal{F}$ via the action of $\mathcal{B}_{d}^{\bf w}$ on $\mathcal{F}$, and vice versa.
[…]
for $1\leq a\leq d$. As ${\mathfrak{gl}}_{d}$-modules,
$\mathcal{F}\cong\left(\!\bigotimes_{r=1}^{p}{\mathbb{C}}[{\bf y}_{r}]\!\right)\otimes\left(\!\bigotimes_{s=1}^{q}{\mathbb{C}}[{\boldsymbol{\zeta}}_{s}]\!\right)\otimes\left(\!\bigotimes_{i=1}^{m}{\mathbb{C}}[{\bf x}_{i}]\!\right)\otimes\left(\!\bigotimes_{j=1}^{n}{\mathbb{C}}[{\boldsymbol{\eta}}_{j}]\!\right)\!,$ (5.1)
and as ${\mathfrak{gl}}_{p+m|q+n}$-modules,
$\mathcal{F}\cong{\mathbb{C}}[{\boldsymbol{\Sigma}}^{1}]\otimes\cdots\otimes{\mathbb{C}}[{\boldsymbol{\Sigma}}^{d}].$ (5.2)
The Howe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ asserts that $\mathcal{F}$ decomposes into a direct sum of unitarizable ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{p+m|q+n}$-modules (with respect to $(\pi,{\boldsymbol{\sigma}})$). We refer the reader to [CLZ, Theorem 3.3] for further details.
For $\xi\in{\mathfrak{h}}_{{p+m|q+n}}^{*}$, we denote by $L_{p+m|q+n}(\xi)$ the irreducible highest weight ${\mathfrak{gl}}_{p+m|q+n}$-module with highest weight $\xi$ with respect to the Borel subalgebra $\mathfrak{b}_{{p+m|q+n}}$. Let
$\mathbbm{1}_{p|q}=\sum_{i=1}^{p}\epsilon_{i}-\sum_{j=1}^{q}\epsilon_{p+j}\qquad\text{and}\qquad\mathbbm{1}_{p}=\mathbbm{1}_{p|0}.$
Suppose $p\not=0\not=m$. For $r\in{\mathbb{Z}}$, let
$V_{r}=\begin{cases}\displaystyle{L_{{p+m|q+n}}\big{(}r\epsilon_{p+q+1}-\mathbbm{1}_{p|q}\big{)}}&\quad\mbox{if $r\geq 0$,}\\ \displaystyle{L_{{p+m|q+n}}\left(\!-\sum_{i=q+r+1}^{q}\epsilon_{p+i}-\mathbbm{1}_{p|q}\!\right)}&\quad\mbox{if $-q\leq r<0$ and $q\not=0$,}\\ \displaystyle{L_{{p+m|q+n}}\left(\!(r+q)\epsilon_{p}-\sum_{i=1}^{q}\epsilon_{p+i}-\mathbbm{1}_{p|q}\!\right)}&\quad\mbox{if $r<-q$ and $q\not=0$,}\\ \displaystyle{L_{{p+m|q+n}}\big{(}r\epsilon_{p}-\mathbbm{1}_{p}\big{)}}&\quad\mbox{if $r<0$ and $q=0$}.\end{cases}$
Note that $V_{r}$ is an infinite-dimensional unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-module (with respect to ${\boldsymbol{\sigma}}$), and that the highest weight of $V_{r}$ corresponds to a generalized partition of depth 1; see [CCL2, Section 2.5] and also [CLZ]. (Such a partition is called a generalized partition of length 1 in [CLZ].)
For any weight $\mu$ of a ${\mathfrak{gl}}_{p+m|q+n}$-module $M$, we denote by $M_{\mu}$ the $\mu$-weight space of $M$. For ${\mathfrak{gl}}_{p+m|q+n}$-modules $M_{1},\ldots,M_{d}$ and ${\bf w}=(w_{1},\ldots,w_{d})\in{\mathbb{C}}^{d}$, we write
$\underline{M\mkern-4.0mu}\mkern 4.0mu=M_{1}\otimes\cdots\otimes M_{d}\qquad\text{and}\qquad\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})=M_{1}(w_{1})\otimes\cdots\otimes M_{d}(w_{d}).$
Analogous to (4.16), we write
$L_{p+m}(\xi)=L_{p+m|0+0}(\xi)\qquad\text{and}\qquad L_{m}(\xi)=L_{0+m|0+0}(\xi).$
Theorem 5.1.
Suppose $p\not=0\not=m$. For $a=1,\ldots,d$, let $M_{a}=V_{\gamma_{a}}$ for some $\gamma_{a}\in{\mathbb{Z}}$. For each weight $\mu$ of $\underline{M\mkern-4.0mu}\mkern 4.0mu$, ${\bf w}\in{\boldsymbol{X}}_{\!d}$ and ${\bf z}\in{\boldsymbol{X}}_{\!p+q+m+n}$, we have:
  1. (i)
    $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu}$ is a cyclic ${\mathcal{B}_{p+m|q+n}^{\bf z}}$-module.
  2. (ii)
[…]
  3. (iii)
[…]
  4. (iv)
[…]
Proof.
Fix $1\leq a\leq d$. If $\gamma_{a}\geq 0$, then we may identify $M_{a}$ with the ${\mathfrak{gl}}_{p+m|q+n}$-submodule of the polynomial superalgebra ${\mathbb{C}}[{\boldsymbol{\Sigma}}^{a}]$ generated by the highest weight vector $(x^{a}_{1})^{\gamma_{a}}$. If $\gamma_{a}<0$, then we may identify $M_{a}$ with the ${\mathfrak{gl}}_{p+m|q+n}$-submodule of ${\mathbb{C}}[{\boldsymbol{\Sigma}}^{a}]$ generated by the highest weight vector
$\begin{cases}\zeta^{a}_{q+\gamma_{a}+1}\ldots\zeta^{a}_{q}&\quad\text{if }\mbox{$\gamma_{a}\geq-q$ and $q\not=0$,}\\ (y^{a}_{p})^{-\gamma_{a}-q}\zeta^{a}_{1}\ldots\zeta^{a}_{q}&\quad\text{if }\mbox{$\gamma_{a}<-q$ and $q\not=0$,}\\ (y^{a}_{p})^{-\gamma_{a}}&\quad\text{if }q=0.\end{cases}$
Under these identifications, the tensor product $\underline{M\mkern-4.0mu}\mkern 4.0mu$ is a ${\mathfrak{gl}}_{p+m|q+n}$-submodule of ${\mathbb{C}}[{\boldsymbol{\Sigma}}^{1}]\otimes\cdots\otimes{\mathbb{C}}[{\boldsymbol{\Sigma}}^{d}]$ and hence a ${\mathfrak{gl}}_{p+m|q+n}$-submodule of $\mathcal{F}$ via (5.2). We will see in a moment that for each weight $\mu:=\sum_{i=1}^{p+q+m+n}\mu_{i}\epsilon_{i}$ of $\underline{M\mkern-4.0mu}\mkern 4.0mu$, the $\mu$-weight space $\underline{M\mkern-4.0mu}\mkern 4.0mu_{\!\mu}$ is also a weight space of a tensor product of finite-dimensional irreducible ${\mathfrak{gl}}_{d}$-modules.
For $1\leq i\leq p$ (resp., $p+1\leq i\leq p+q$, $p+q+1\leq i\leq p+q+m$ and $p+q+m+1\leq i\leq p+q+m+n$), let $N_{i}$ denote the ${\mathfrak{gl}}_{d}$-submodule of ${\mathbb{C}}[y^{1}_{i},\cdots,y^{d}_{i}]$ (resp., ${\mathbb{C}}[\zeta^{1}_{i-p},\cdots,\zeta^{d}_{i-p}]$, ${\mathbb{C}}[x^{1}_{i-p-q},\cdots,x^{d}_{i-p-q}]$ and ${\mathbb{C}}[\eta^{1}_{i-p-q-m},\cdots,\eta^{d}_{i-p-q-m}]$) generated by the highest weight vector $(y^{d}_{i})^{-\mu_{i}-d}$ (resp., $\zeta^{\mu_{i}+1}_{i-p}\cdots\zeta^{d}_{i-p}$, $(x^{1}_{i-p-q})^{\mu_{i}}$ and $\eta^{1}_{i-p-q-m}\cdots\eta^{\mu_{i}}_{i-p-q-m}$). Then
$N_{i}\cong\begin{cases}L_{d}\big{(}(\mu_{i}+d)\epsilon_{d}\big{)}&\quad\text{if }1\leq i\leq p,\\ \displaystyle{L_{d}\left(\!-\sum_{a=\mu_{i}+1}^{d}\epsilon_{a}\!\right)}&\quad\text{if }p+1\leq i\leq p+q,\\ L_{d}(\mu_{i}\epsilon_{1})&\quad\text{if }p+q+1\leq i\leq p+q+m,\\ \displaystyle{L_{d}\left(\sum_{a=1}^{\mu_{i}}\epsilon_{a}\right)}&\quad\text{if }p+q+m+1\leq i\leq p+q+m+n,\end{cases}$
and so $N_{i}$ is an irreducible highest weight ${\mathfrak{gl}}_{d}$-module with a dominant integral highest weight. Write ${\bf z}=(z_{1},\ldots,z_{p+q+m+n})$. Let
$\underline{N\mkern-4.0mu}\mkern 4.0mu=N_{1}\otimes\cdots\otimes N_{p+q+m+n}\quad\text{and}\quad\underline{N\mkern-4.0mu}\mkern 4.0mu({\bf z})=N_{1}(z_{1})\otimes\cdots\otimes N_{p+q+m+n}(z_{p+q+m+n}).$
The tensor product $\underline{N\mkern-4.0mu}\mkern 4.0mu$ is a ${\mathfrak{gl}}_{d}$-submodule of $\mathcal{F}$ via (5.1). By [CLZ, Theorem 3.3], we have
$\underline{M\mkern-4.0mu}\mkern 4.0mu_{\mu}\subset\underline{N\mkern-4.0mu}\mkern 4.0mu\qquad\hbox{and}\qquad\underline{N\mkern-4.0mu}\mkern 4.0mu_{\gamma}\subset\underline{M\mkern-4.0mu}\mkern 4.0mu,$
and no nonzero element in the $\nu$-weight space $\underline{N\mkern-4.0mu}\mkern 4.0mu_{\nu}$ is contained in $\underline{M\mkern-4.0mu}\mkern 4.0mu$ for all $\nu\not=\gamma$, where $\gamma:=\sum_{a=1}^{d}\gamma_{a}\epsilon_{a}\in{\mathfrak{h}}_{d}^{*}$. We deduce that $\underline{M\mkern-4.0mu}\mkern 4.0mu_{\mu}\subseteq\underline{N\mkern-4.0mu}\mkern 4.0mu_{\gamma}$. Similarly, $\underline{N\mkern-4.0mu}\mkern 4.0mu_{\gamma}\subseteq\underline{M\mkern-4.0mu}\mkern 4.0mu_{\mu}$. Thus we have
$\underline{M\mkern-4.0mu}\mkern 4.0mu_{\mu}=\underline{N\mkern-4.0mu}\mkern 4.0mu_{\gamma}.$
This gives
[…]
in view of Theorem 4.7. By Theorem 2.1, $\underline{N\mkern-4.0mu}\mkern 4.0mu({\bf z})_{\gamma}$ is a cyclic $\mathcal{B}_{d}^{\bf w}$-module as $\mathcal{B}_{d}^{\bf w}$ preserves weight spaces, and by Theorem 3.8, $(\mathcal{B}_{d}^{\bf w})_{\underline{N\mkern-4.0mu}\mkern 4.0mu({\bf z})_{\gamma}}$ is a Frobenius algebra. Hence, (i) and (ii) follow from (5.3). Also, (iii) and (iv) are immediate consequences of [Lu, Lemma 1.3] (cf. Remark 3.10). The last statement follows from (5.3) and Theorem 2.2. ∎
Remark 5.2.
A similar statement corresponding to $p=0$ or $m=0$ can be formulated. The details will be left to the reader.
Suppose $p\not=0\not=m$. For $r\in{\mathbb{Z}}$, let
$V_{r}=\begin{cases}L_{p+m}(r\epsilon_{p+1}-\mathbbm{1}_{p})&\quad\mbox{if $r\geq 0$,}\\ L_{p+m}(r\epsilon_{p}-\mathbbm{1}_{p})&\quad\mbox{if $r<0$}.\end{cases}$
We have the following result.
Corollary 5.3.
For $a=1,\ldots,d$, let $M_{a}=V_{\gamma_{a}}$ for some $\gamma_{a}\in{\mathbb{Z}}$. For each weight $\mu$ of $\underline{M\mkern-4.0mu}\mkern 4.0mu$, ${\bf w}\in{\boldsymbol{X}}_{\!d}$ and ${\bf z}\in{\boldsymbol{X}}_{\!p+m}$, we have:
  1. (i)
    $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu}$ is a cyclic $\mathcal{B}_{p+m}^{\bf z}$-module.
  2. (ii)
    $(\mathcal{B}_{p+m}^{\bf z})_{\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu}}$ is a Frobenius algebra.
  3. (iii)
    $(\mathcal{B}_{p+m}^{\bf z})_{\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu}}$ is a maximal commutative subalgebra of ${\rm End}(\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu})$ of dimension $\dim\!\left(\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu}\right)$.
  4. (iv)
    Every eigenspace of $(\mathcal{B}_{p+m}^{\bf z})_{\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu}}$ is one-dimensional.
Moreover, $(\mathcal{B}_{p+m}^{\bf z})_{\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu}}$ is diagonalizable with a simple spectrum for generic ${\bf w}$ and generic ${\bf z}$.
Proof.
This is an immediate consequence of Theorem 5.1 by taking $q=n=0$. ∎
The structures of the Bethe algebras of finite-dimensional modules over a general linear Lie algebra have been well studied (see [FFRy, MTV2]). However, very little is known about the Bethe algebras of infinite-dimensional modules. Corollary 5.3, which concerns the action of the Bethe algebra of ${\mathfrak{gl}}_{p+m}$ on a tensor product of infinite-dimensional unitarizable highest weight ${\mathfrak{gl}}_{p+m}$-modules, makes good progress toward the infinite-dimensional cases.
We expect that Theorem 2.1 and Theorem 2.2 should have unitarizable analogs. More generally, we have the following conjecture.
Conjecture 5.4.
Let $d\in{\mathbb{N}}$ and $p,q,m,n\in{\mathbb{Z}}_{+}$ that are not all zero. For $a=1,\ldots,d$, let $M_{a}$ be a unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-module. For each weight $\mu$ of $\underline{M\mkern-4.0mu}\mkern 4.0mu$, ${\bf w}\in{\boldsymbol{X}}_{\!d}$ and ${\bf z}\in{\boldsymbol{X}}_{\!p+q+m+n}$, parts (i)–(iv) of Theorem 5.1 hold for the Bethe algebra $({\mathcal{B}_{p+m|q+n}^{\bf z}})_{\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu}}$. Moreover, $({\mathcal{B}_{p+m|q+n}^{\bf z}})_{\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})_{\mu}}$ is diagonalizable with a simple spectrum for generic ${\bf w}$ and generic ${\bf z}$.
Remark 5.5.
An analog of the conjecture when the entries of ${\bf z}$ are all zero will be studied in [CCL2].
Acknowledgments. The first author was partially supported by NSTC grant 113-2115-M-006-010 of Taiwan. The second author was partially supported by NSTC grant 112-2115-M-006-015-MY2 of Taiwan.