### v2

# Bethe algebras for unitarizable modules over classical Lie (super)algebras and a duality

Abstract.
Let $\mathfrak{g}$} denote the classical Lie algebra ${\mathfrak{gl}}_{d}$, $\mathfrak{sp}_{2d}$, or $\mathfrak{so}_{2d}$ with a fixed *-structure $\bm{\sigma}$}. Let $M_{1},\ldots,M_{\ell}$ be unitarizable $\mathfrak{g}$}-modules (with respect to $\bm{\sigma}$}), and let ${\bf z}=(z_{1},\ldots,z_{\ell})\in{\mathbb{C}}^{\ell}$. We investigate the action of the Bethe algebra $\mathcal{B}_{\mathfrak{g}}^{\mu}$ for $\mathfrak{g}$} with respect to $\mu\in{\mathfrak{g}}^{*}$ on the tensor product $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z}):=M_{1}(z_{1})\otimes\cdots\otimes M_{\ell}(z_{\ell})$ of evaluation ${\mathfrak{g}}[t]$-modules. We show that if $\mu\circ{\bm{\sigma}}$ equals the complex conjugation of $\mu$, then $\mathcal{B}_{\mathfrak{g}}^{\mu}$ is diagonalizable on any finite-dimensional $\mathcal{B}_{\mathfrak{g}}^{\mu}$-submodule of $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})$ for $\bf z$$\in{\mathbb{R}}^{\ell}$. This, together with the result derived from the duality of Bethe algebras (see below), suggests that a simple spectrum conjecture for $\mathcal{B}_{\mathfrak{g}}^{\mu}$ should hold.
We establish a duality of Bethe algebras for the general linear Lie (super)algebras ${\mathfrak{gl}}_{d}$ and ${\mathfrak{gl}}_{p+m|q+n}$. As an application, we show that under a generic condition, the Bethe algebra for ${\mathfrak{gl}}_{p+m|q+n}$ with respect to ${\bf z}\in{\mathbb{C}}^{p+q+m+n}$ is diagonalizable with a simple spectrum on any weight space of $L_{1}(w_{1})\otimes\cdots\otimes L_{d}(w_{d})$, where the $L_{i}$ are (infinite-dimensional) unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-modules corresponding to generalized partitions of depth 1, and $w_{1},\ldots,w_{d}\in{\mathbb{C}}$. We also obtain the corresponding result for ${\mathfrak{gl}}_{p+m}$ by setting $q=n=0$.
1. Introduction
Let $\mathfrak{g}$} be a finite-dimensional semisimple Lie algebra over $\mathbb{C}$}. The Bethe algebra for $\mathfrak{g}$} is a commutative subalgebra of $U({\mathfrak{g}}[t])$ induced by the Feigin–Frenkel center (see [FFR, FFTL, Ry1]), where $U({\mathfrak{g}}[t])$ denotes the universal enveloping algebra of the current algebra ${\mathfrak{g}}[t]$. It produces (higher) Gaudin Hamiltonians, including the quadratic ones originally introduced by Gaudin [G1, G2] to describe a completely integrable quantum spin chain model. Also, the notion of the Bethe algebra can be extended to the general linear Lie (super)algebras (see [HM, MR, MTV1, Ta]).
Let $d$ be a positive integer, and let $p$, $q$, $m$, and $n$ be non-negative integers that are not all zero. The aim of the paper is two-fold. Firstly, we investigate the action of the Bethe algebra for $\mathfrak{g}$} on unitarizable $\mathfrak{g}$}-modules, where $\mathfrak{g}$} is the classical Lie algebra of type $\mathfrak{a,c,d}$ of rank $d$ (i.e., ${\mathfrak{g}}={\mathfrak{gl}}_{d}$, $\mathfrak{sp}_{2d}$, or $\mathfrak{so}_{2d}$). Secondly, we establish a duality of Bethe algebras for the general linear Lie (super)algebras ${\mathfrak{gl}}_{d}$ and ${\mathfrak{gl}}_{p+m|q+n}$ and use it to study the action of the Bethe algebra for ${\mathfrak{gl}}_{p+m|q+n}$ on some infinite-dimensional unitarizable ${\mathfrak{gl}}_{p+m|q+n}$-modules.
1.1. Unitarizable modules and diagonalization of Bethe algebras
Let $\mathfrak{g}$} be a simple Lie algebra or the general linear Lie algebra over $\mathbb{C}$}. For $\mu\in{\mathfrak{g}}^{*}$, let $\mathcal{B}_{\mathfrak{g}}^{\mu}$ be the Bethe algebra for $\mathfrak{g}$} with respect to $\mu$; see Section 3.2. The action of $\mathcal{B}_{\mathfrak{g}}^{\mu}$ on a tensor product of finite-dimensional $\mathfrak{g}$}-modules has been extensively studied (see [FFRy, MTV1, MTV2, MTV4, MTV5, Ry3]). Let ${\bf z}=(z_{1},\ldots,z_{\ell})\in{\mathbb{C}}^{\ell}$, and let $V_{1},\ldots,V_{\ell}$ be finite-dimensional irreducible $\mathfrak{g}$}-modules. Feigin, Frenkel, and Rybnikov [FFRy] show that the algebra $\mathcal{B}_{\mathfrak{g}}^{\mu}$ acts cyclically on the tensor product $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z}):=V_{1}(z_{1})\otimes\cdots\otimes V_{\ell}(z_{\ell})$ of evaluation ${\mathfrak{g}}[t]$-modules for any regular $\mu$ and any pairwise distinct $z_{1},\ldots,z_{\ell}$. They also show that $\mathcal{B}_{\mathfrak{g}}^{\mu}$ is diagonalizable with a simple spectrum on $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$ for generic $\mu$ and generic $\bf z$}.
Now let ${\mathfrak{g}}={\mathfrak{gl}}_{d}$, $\mathfrak{sp}_{2d}$, or $\mathfrak{so}_{2d}$. It possesses a *-structure ${\bm{\sigma}}:{\mathfrak{g}}\longrightarrow{\mathfrak{g}}$, i.e., $\bm{\sigma}$} is an even anti-linear anti-involution (see Section 3.3). Let $M_{1},\ldots,M_{\ell}$ be unitarizable $\mathfrak{g}$}-modules with respect to $\bm{\sigma}$}, and let $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})=M_{1}(z_{1})\otimes\cdots\otimes M_{\ell}(z_{\ell})$. Define ${\mathfrak{g}}^{*}_{{\bm{\sigma}}}=\left\{\,\mu\in{\mathfrak{g}}^{*}\;\middle|\;\mu\circ{\bm{\sigma}}=\overline{\mu}\,\right\}$, where $\overline{\mu}\in{\mathfrak{g}}^{*}$ is the complex conjugation of $\mu$. Using the properties of *-structures and Feigin–Frenkel centers, we obtain the following.
Theorem 1.1 (Theorem 3.10).
The Bethe algebra $\mathcal{B}_{\mathfrak{g}}^{\mu}$ is diagonalizable on any finite-dimensional $\mathcal{B}_{\mathfrak{g}}^{\mu}$-submodule of $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})$ for any $\mu\in{\mathfrak{g}}^{*}_{{\bm{\sigma}}}$ and $\bf z$$\in{\mathbb{R}}^{\ell}$.
For ${\bf w}:=(w_{1},\ldots,w_{d})\in{\mathbb{C}}^{d}$, let $\mu_{\bf w}\in{\mathfrak{g}}^{*}$ be the element defined by (3.12) and (3.13). In other words, $\mu_{\bf w}$ corresponds to a diagonal matrix. We write $\mathcal{B}_{\mathfrak{g}}^{\bf w}:=\mathcal{B}_{\mathfrak{g}}^{\mu_{\bf w}}$.
Let $L_{1},\ldots,L_{\ell}$ be unitarizable highest weight $\mathfrak{g}$}-modules with respect to $\bm{\sigma}$}. Set $\underline{L}=L_{1}\otimes\cdots\otimes L_{\ell}$ and $\underline{L}({\bf z})=L_{1}(z_{1})\otimes\cdots\otimes L_{\ell}(z_{\ell})$. Let $\mathfrak{h}$} denote the standard Cartan subalgebra of $\mathfrak{g}$} (see Section 2.1.4). Note that for any weight $\gamma\in{\mathfrak{h}}^{*}$ of $\underline{L}$, the $\gamma$-weight space $\underline{L}({\bf z})_{\gamma}$ of $\underline{L}({\bf z})$ is a finite-dimensional $\mathcal{B}_{\mathfrak{g}}^{\bf w}$-module, and that $\mu_{\bf w}\in{\mathfrak{g}}^{*}_{{\bm{\sigma}}}$ for ${\bf w}\in{\mathbb{R}}^{d}$. We have the following as a consequence of Theorem 1.1.
Corollary 1.2 (Corollary 3.12).
Let $\gamma\in{\mathfrak{h}}^{*}$ be any weight of $\underline{L}$. Then $\mathcal{B}_{\mathfrak{g}}^{\bf w}$ is diagonalizable on $\underline{L}({\bf z})_{\gamma}$ for any ${\bf w}\in{\mathbb{R}}^{d}$ and $\bf z$$\in{\mathbb{R}}^{\ell}$.
A natural question is whether there are unitarizable analogs of the results of [FFRy] mentioned above.
Conjecture 1.3 (Conjecture 3.13).
Let $\gamma\in{\mathfrak{h}}^{*}$ be any weight of $\underline{L}$. Then the weight space $\underline{L}({\bf z})_{\gamma}$ is a cyclic $\mathcal{B}_{\mathfrak{g}}^{\bf w}$-module for any pairwise distinct $w_{1},\ldots,w_{d}\in{\mathbb{C}}$ and any pairwise distinct $z_{1},\ldots,z_{\ell}\in{\mathbb{C}}$. Moreover, $\mathcal{B}_{\mathfrak{g}}^{\bf w}$ is diagonalizable with a simple spectrum on $\underline{L}({\bf z})_{\gamma}$ for generic $\bf w$} and generic $\bf z$}.
Theorem 1.5 (see below) provides positive evidence for Conjecture 1.3 when $q$ and $n$ are set to 0.
1.2. A duality for the general linear Lie (super)algebras
For ${\bf w}:=(w_{1},\ldots,w_{d})\in{\mathbb{C}}^{d}$ and ${\bf z}:=(z_{1},\ldots,z_{p+q+m+n})\in{\mathbb{C}}^{p+q+m+n}$, we consider the Bethe algebra $\mathcal{B}_{d}^{\bf w}:=\mathcal{B}_{{\mathfrak{gl}}_{d}}^{\bf w}$ for the general linear Lie algebra ${\mathfrak{gl}}_{d}$ with respect to $\bf w$} and the Bethe algebra ${\mathcal{B}_{p+m|q+n}^{\bf z}}:={\mathcal{B}_{{\mathfrak{gl}}_{p+m|q+n}}^{\bf z}}$ for the general linear Lie superalgebra ${\mathfrak{gl}}_{p+m|q+n}$ with respect to $\bf z$}; see Section 4.2. Note that $\mathcal{B}_{d}^{\bf w}$ is a subalgebra of $U({\mathfrak{gl}}_{d}[t])$ generated by the coefficients of the series ${\rm cdet}({\mathcal{L}}_{d}^{\bf w})$ while $\mathcal{B}_{p+m|q+n}^{\bf z}$} is a subalgebra of $U({\mathfrak{gl}}_{p+m|q+n}[t])$ generated by the coefficients of the series ${\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big)$. Here $\rm cdet$} and ${\rm Ber}^{\bf s}$ are respectively the column determinant and the Berezinian of type ${\bf s}:=(0^{p},1^{q},0^{m},1^{n})$ (see Section 4.1), and ${\mathcal{L}}_{d}^{\bf w}$ and ${\mathcal{L}}_{{p+m|q+n}}^{\bf z}$ are respectively the matrices defined in (4.4) and (4.7).
Let $\mathcal{F}$ be the polynomial superalgebra generated by $x^{a}_{i}$, $\eta^{a}_{j}$, $y^{a}_{r}$, and $\zeta^{a}_{s}$, for $i=1,\ldots,m$, $j=1,\ldots,n$, $r=1,\ldots,p$, $s=1,\ldots,q$, and $a=1,\ldots,d$. Here $x^{a}_{i}$ and $y^{a}_{r}$ are even variables while $\eta^{a}_{j}$ and $\zeta^{a}_{s}$ are odd. The superalgebra $\mathcal{F}$ can be realized as the Fock space of $d(p+m)$ bosonic and $d(q+n)$ fermionic oscillators (cf. [CL, LZ]). Let $\mathcal{D}$ be the corresponding Weyl superalgebra, i.e., $\mathcal{D}$ is the associative unital superalgebra generated by $x^{a}_{i}$, $\eta^{a}_{j}$, $y^{a}_{r}$, and $\zeta^{a}_{s}$ as well as their derivatives $\frac{\partial}{\partial x_{i}^{a}}$, $\frac{\partial}{\partial\eta_{j}^{a}}$, $\frac{\partial}{\partial y_{r}^{a}}$, and $\frac{\partial}{\partial\zeta_{r}^{a}}$. There exist superalgebra homomorphisms $\phi:U({\mathfrak{gl}}_{d})\longrightarrow\mathcal{D}$ and $\varphi:U({\mathfrak{gl}}_{p+m|q+n})\longrightarrow\mathcal{D}$; see (4.9) and (4.10). They extend to superalgebra homomorphisms $\phi_{\bf z}:U({\mathfrak{gl}}_{d}[t])(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!)$ and $\varphi_{\bf w}:U({\mathfrak{gl}}_{p+m|q+n}[t])(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!)$, respectively. Here, for instance, $\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!)$ is the superalgebra of pseudo-differential operators over $\mathcal{D}$; see Section 2.3. The maps $\phi_{\bf z}$ and $\varphi_{\bf w}$ induce an action of $\mathcal{B}_{d}^{\bf w}$ and an action of $\mathcal{B}_{p+m|q+n}^{\bf z}$} on $\mathcal{F}$, respectively. The following theorem asserts that the actions are equivalent. We call it the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$.
Theorem 1.4 (Theorem 4.10).
The coefficients of $\phi_{\bf z}{\rm cdet}({\mathcal{L}}_{d}^{\bf w})$ are determined by those of $\varphi_{\bf w}{\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big)$, and vice versa. Consequently, $\phi_{\bf z}\big(\mathcal{B}_{d}^{\bf w}\big)=\varphi_{\bf w}\big({\mathcal{B}_{p+m|q+n}^{\bf z}}\big)$.
Theorem 1.4 recovers several dualities established earlier. After specializing to $p=q=0$, it recovers the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m|n})$ due to Huang and Mukhin [HM]. If we assume further that $n=0$ (resp., $m=0$), Theorem 4.10 gives the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m})$ obtained by Mukhin, Tarasov, and Varchenko [MTV3] (resp., a variant of the duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{n})$ obtained by Tarasov and Uvarov [TU]).
Also, another duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m})$ is obtained by Vicedo and Young [VY] for Gaudin models with irregular singularities. It is extended by the authors [ChL3] to a duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$, which specializes to a variant of Theorem 1.4.
Using the properties of the actions of $\mathcal{B}_{d}^{\bf w}$ on tensor products of finite-dimensional ${\mathfrak{gl}}_{d}$-modules, Theorem 1.4 enables us to better understand the actions of $\mathcal{B}_{p+m|q+n}^{\bf z}$} on tensor products of infinite-dimensional unitarizable ${\mathfrak{gl}}_{p+m|q+n}$-modules, as shown in the following theorem.
Theorem 1.5 (Theorem 4.13).
Let $\underline{L}=L_{1}\otimes\cdots\otimes L_{d}$ and $\underline{L}({\bf w})=L_{1}(w_{1})\otimes\cdots\otimes L_{d}(w_{d})$, where the $L_{i}$ are (infinite-dimensional) unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-modules corresponding to generalized partitions of depth 1 (see (4.17)). Let $\mu\in{\mathfrak{h}}_{{p+m|q+n}}^{*}$ be any weight of $\underline{L}$ (where ${\mathfrak{h}}_{{p+m|q+n}}$ is the Cartan subalgebra of ${\mathfrak{gl}}_{p+m|q+n}$ defined in Section 2.1.1). Then the weight space $\underline{L}({\bf w})_{\mu}$ is a cyclic $\mathcal{B}_{p+m|q+n}^{\bf z}$}-module for any pairwise distinct $z_{1},\ldots,z_{p+q+m+n}\in{\mathbb{C}}$ and any pairwise distinct $w_{1},\ldots,w_{d}\in{\mathbb{C}}$. Moreover, $\mathcal{B}_{p+m|q+n}^{\bf z}$} is diagonalizable with a simple spectrum on $\underline{L}({\bf w})_{\mu}$ for generic $\bf z$} and generic $\bf w$}.
1.3. Outline
We organize the paper as follows. Section 2 is devoted to reviewing the classical Lie (super)algebras of types $\mathfrak{a,c,d}$ and some background material needed in this paper. In Section 3, we investigate the action of the Bethe algebra for $\mathfrak{g}$} on a tensor product of evaluation ${\mathfrak{g}}[t]$-modules, where $\mathfrak{g}$} is the classical Lie algebra of type $\mathfrak{a,c,d}$. We prove Theorem 1.1 and obtain Corollary 1.2. In Section 4, we establish Theorem 1.4, i.e., the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$. We give an application of the duality and demonstrate Theorem 1.5, which provides positive evidence for Conjecture 1.3. In Appendix A, we prove Proposition 4.7, which gives an equality relating $\phi_{\bf z}{\rm cdet}({\mathcal{L}}_{d}^{\bf w})$ to $\varphi_{\bf w}{\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big)$.
Notations. Throughout the paper, the symbol $\mathbb{Z}$} (resp., $\mathbb{N}$} and $\mathbb{Z}$$_{+}$) stands for the set of all (resp., positive and non-negative) integers, the symbol $\mathbb{C}$} (resp., $\mathbb{R}$}) for the field of complex (resp., real) numbers, and the symbol ${\mathbb{Z}}_{2}:=\{\bar{0},\bar{1}\}$ for the field of integers modulo 2. Unless otherwise stated, all vector spaces, algebras, tensor products, etc., are over $\mathbb{C}$}. We fix $d\in{\mathbb{N}}$ and $p,q,m,n\in{\mathbb{Z}}_{+}$.
2. Preliminaries
In this section, we review the classical Lie (super)algebras of types $\mathfrak{a,c,d}$, evaluation modules, and pseudo-differential operators.
2.1. Classical Lie (super)algebras
2.1.1. The general linear Lie (super)algebra
For $p,q,m,n\in{\mathbb{Z}}_{+}$ that are not all zero, let
$\mathbb{I}=\{i\in{\mathbb{N}}\,|\,1\leq i\leq p+q+m+n\}.$
For $i\in\mathbb{I}$, define $|i|\in{\mathbb{Z}}_{2}$ by
$|i|=\begin{cases}\,\bar{0}&\,\,\mbox{if}\ \ i\in\{1,\ldots,p\}\cup\{p+q+1,\ldots,p+q+m\},\\ \,\bar{1}&\,\,\mbox{otherwise}.\end{cases}$
Let $\{e_{i}\,|\,i\in\mathbb{I}\}$ be a basis for the superspace ${\mathbb{C}}^{p|q}\oplus{\mathbb{C}}^{m|n}$ such that $\{e_{i}\,|\,1\leq i\leq p+q\}$ and $\{e_{p+q+i}\,|\,1\leq i\leq m+n\}$ are respectively the standard homogeneous bases for ${\mathbb{C}}^{p|q}$ and ${\mathbb{C}}^{m|n}$. In other words, the parity of $e_{i}$ is given by $|e_{i}|=|i|$ for $i\in\mathbb{I}$.
For any $i,j\in\mathbb{I}$, let $E^{i}_{j}$ denote the $\mathbb{C}$}-linear endomorphism on ${\mathbb{C}}^{p|q}\oplus{\mathbb{C}}^{m|n}$ defined by
$E^{i}_{j}(e_{k})=\delta_{j,k}e_{i}\quad\mbox{ for $k\in\mathbb{I}$,}$
where $\delta$ is the Kronecker delta. The parity of $E^{i}_{j}$ is given by $|E^{i}_{j}|=|i|+|j|$. The superspace of $\mathbb{C}$}-linear endomorphisms on ${\mathbb{C}}^{p|q}\oplus{\mathbb{C}}^{m|n}$ has a natural structure of a Lie superalgebra, called the general linear Lie (super)algebra and denoted by ${\mathfrak{gl}}_{p+m|q+n}$. The commutation relations of ${\mathfrak{gl}}_{p+m|q+n}$ are given by
$E^{i}_{j},E^{k}_{l}]=\delta_{j,k}E^{i}_{l}-(-1)^{(|i|+|j|)(|k|+|l|)}\delta_{i,l}E^{k}_{j}\qquad\mbox{for $i,j,k,l\in\mathbb{I}$.}$
Note that $\{E^{i}_{j}\,|\,i,j\in\mathbb{I}\}$ is a homogeneous basis for ${\mathfrak{gl}}_{p+m|q+n}$.
Let $\mathfrak{b}_{{p+m|q+n}}=\bigoplus_{\begin{subarray}{c}i,j\in\mathbb{I},i\leq j\end{subarray}}{\mathbb{C}}E^{i}_{j}$ be a Borel subalgebra of ${\mathfrak{gl}}_{p+m|q+n}$. The corresponding Cartan subalgebra ${\mathfrak{h}}_{{p+m|q+n}}$ has a basis $\{E^{i}_{i}\,|\,i\in\mathbb{I}\}$. We denote the dual basis in ${\mathfrak{h}}_{{p+m|q+n}}^{*}$ by $\{\epsilon_{i}\,|\,i\in\mathbb{I}\}$. The parity of $\epsilon_{i}$ is given by $|\epsilon_{i}|=|i|$.
For $m=d$ and $p=q=n=0$, we define
${\mathfrak{gl}}_{d}:={\mathfrak{gl}}_{d|0}\qquad\text{and}\qquad e^{\mathfrak{a}}_{ij}:=e_{ij}:=E^{i}_{j}\qquad\mbox{for $i,j=1,\ldots,d$.}$ (2.1)
2.1.2. The symplectic Lie algebra
The subalgebra of ${\mathfrak{gl}}_{2d}$ that preserves the nondegenerate skew-symmetric bilinear form on ${\mathbb{C}}^{2d}$ corresponding to the matrix
$\left[\begin{matrix}0&J_{d}\\ -J_{d}&0\end{matrix}\right]$
is called the symplectic Lie algebra and is denoted by $\mathfrak{sp}_{2d}$. Hereafter, $J_{d}$ is the $d\times d$ reversion matrix defined by $(J_{d})_{ij}=1$ if $i+j=d+1$ and $(J_{d})_{ij}=0$ otherwise.
The Lie algebra $\mathfrak{sp}_{2d}$ is simple and is spanned by
$e^{\mathfrak{c}}_{ij}:=e_{ij}-\varepsilon_{i}\varepsilon_{j}e_{j^{\prime}i^{\prime}}$ (2.2)
for $i,j=1,\ldots,2d$, where $i^{\prime}:=2d-i+1$, and $\varepsilon_{i}:=1$ if $1\leq i\leq d$ and $\varepsilon_{i}:=-1$ if $d+1\leq i\leq 2d$.
2.1.3. The orthogonal Lie algebra
The subalgebra of ${\mathfrak{gl}}_{2d}$ that preserves the nondegenerate symmetric bilinear form on ${\mathbb{C}}^{2d}$ corresponding to the matrix
$\left[\begin{matrix}0&J_{d}\\ J_{d}&0\end{matrix}\right]$
is called the orthogonal Lie algebra and is denoted by $\mathfrak{so}_{2d}$. It is spanned by
$e^{\mathfrak{d}}_{ij}:=e_{ij}-e_{j^{\prime}i^{\prime}}$ (2.3)
for $i,j=1,\ldots,2d$, where $i^{\prime}:=2d-i+1$, and is a simple Lie algebra if $d\geq 3$.
2.1.4. Triangular decompositions
The general linear Lie (super)algebras ${\mathfrak{gl}}_{p+m|q+n}$ and ${\mathfrak{gl}}_{d}$ are called the classical Lie (super)algebras of types $\mathfrak{a}$ while the Lie algebras $\mathfrak{sp}_{2d}$ and $\mathfrak{so}_{2d}$ are called the classical Lie algebras of types $\mathfrak{c}$ and $\mathfrak{d}$, respectively. Let
${\mathfrak{g}}^{\mathfrak{a}}={\mathfrak{gl}}_{d},\qquad{\mathfrak{g}}^{\mathfrak{c}}=\mathfrak{sp}_{2d},\qquad{\mathfrak{g}}^{\mathfrak{d}}=\mathfrak{so}_{2d}.$
For ${\mathfrak{x}}=\mathfrak{a,c,d}$, the classical Lie algebra ${\mathfrak{g}}^{\mathfrak{x}}$ has the triangular decomposition
${\mathfrak{g}}^{\mathfrak{x}}=\mathfrak{n}_{-}^{\mathfrak{x}}\oplus{\mathfrak{h}}^{\mathfrak{x}}\oplus\mathfrak{n}_{+}^{\mathfrak{x}},$ (2.4)
where ${\mathfrak{h}}^{\mathfrak{x}}:=\bigoplus_{i=1}^{d}{\mathbb{C}}e^{\mathfrak{x}}_{ii}$ is the standard Cartan subalgebra of ${\mathfrak{g}}^{\mathfrak{x}}$, and $\mathfrak{n}_{+}^{\mathfrak{x}}:=\sum_{\begin{subarray}{c}i<j\end{subarray}}e^{\mathfrak{x}}_{ij}$ and $\mathfrak{n}_{-}^{\mathfrak{x}}:=\sum_{\begin{subarray}{c}i<j\end{subarray}}{\mathbb{C}}e^{\mathfrak{x}}_{ji}$ are nilpotent subalgebras of ${\mathfrak{g}}^{\mathfrak{x}}$. We will suppress the superscript $\mathfrak{x}$} if there is no danger of confusion.
2.2. Evaluation modules
For any Lie superalgebra $\mathfrak{g}$}, we denote by $U({\mathfrak{g}})$ the universal enveloping algebra of $\mathfrak{g}$}. For any even variable $t$, the loop algebra ${\mathfrak{g}}[t,t^{-1}]:={\mathfrak{g}}\otimes{\mathbb{C}}[t,t^{-1}]$ is defined to be the Lie superalgebra with commutation relations
$\hskip 28.45274pt\big[A\otimes t^{r},B\otimes t^{s}\big]=[A,B]\otimes t^{r+s}\qquad\mbox{for $A,B\in{\mathfrak{g}}\,$ and $\,r,s\in{\mathbb{Z}}$.}$
Here $[A,B]$ is the supercommutator of $A$ and $B$. We will use the notation
$\hskip 28.45274ptA[r]:=A\otimes t^{r}\qquad\mbox{for $A\in{\mathfrak{g}}\,$ and $\,r\in{\mathbb{Z}}$.}$
The current algebra ${\mathfrak{g}}[t]:={\mathfrak{g}}\otimes{\mathbb{C}}[t]$ and the superalgebra $t^{-1}{\mathfrak{g}}[t^{-1}]:={\mathfrak{g}}\otimes t^{-1}{\mathbb{C}}[t^{-1}]$ are subalgebras of ${\mathfrak{g}}[t,t^{-1}]$ defined in an obvious way. We identify $\mathfrak{g}$} with the subalgebra ${\mathfrak{g}}\otimes 1$ of constant polynomials in ${\mathfrak{g}}[t]$ and hence $U({\mathfrak{g}})\subseteq U({\mathfrak{g}}[t])$.
For $\alpha\in{\mathbb{C}}$, we have an evaluation homomorphism ${\rm ev}_{\alpha}:U({\mathfrak{g}}[t])\longrightarrow U({\mathfrak{g}})$ given by
${\rm ev}_{\alpha}(A[r])=\alpha^{r}A\qquad\mbox{for $A\in{\mathfrak{g}}\,$ and $\,r\in{\mathbb{Z}}_{+}$}.$
Any $\mathfrak{g}$}-module $M$ is a ${\mathfrak{g}}[t]$-module via ${\rm ev}_{\alpha}$, called the evaluation module and denoted by $M(\alpha)$. For any $A\in{\mathfrak{g}}$ and any even variable $z$, we set
$A(z)=\sum_{r=0}^{\infty}A[r]z^{-r-1}.$
It is a formal power series in $z^{-1}$ with coefficients in ${\mathfrak{g}}[t]$ and acts on $M(\alpha)$ by
$A(z)v=\frac{Av}{z-\alpha}\qquad\mbox{for $v\in M(\alpha)$.}$
Here each rational function in $z$ represents its power series expression at $\infty$.
Let $\ell\in{\mathbb{N}}$ and ${\bf z}=(z_{1},\ldots,z_{\ell})\in{\mathbb{C}}^{\ell}$. Let $\Delta^{(\ell-1)}:U({\mathfrak{g}}[t])\longrightarrow U({\mathfrak{g}}[t])^{\otimes\ell}$ be the $(\ell-1)$-fold coproduct on $U({\mathfrak{g}}[t])$. We have the homomorphism
$\underline{{\rm ev}}_{\bf z}:=({\rm ev}_{z_{1}}\otimes\ldots\otimes{\rm ev}_{z_{\ell}})\circ\Delta^{(\ell-1)}:U({\mathfrak{g}}[t])\longrightarrow U({\mathfrak{g}})^{\otimes{\ell}},$ (2.5)
given by
$\underline{{\rm ev}}_{\bf z}(A[r])=\sum_{i=1}^{\ell}z_{i}^{r}A^{(i)}\qquad\mbox{for $A\in{\mathfrak{g}}\,$ and $\,r\in{\mathbb{Z}}_{+}$}.$
Here $A^{(i)}:=\underbrace{1\otimes\cdots\otimes 1\otimes\stackrel{{\scriptstyle i}}{{A}}\otimes 1\otimes\cdots\otimes 1}_{\ell}$ for $i=1,\ldots,\ell$. We call $\underline{{\rm ev}}_{\bf z}$ the evaluation map at $\bf z$}.
For any $\mathfrak{g}$}-modules $M_{1},\ldots,M_{\ell}$, let
$\underline{M\mkern-4.0mu}\mkern 4.0mu=M_{1}\otimes\cdots\otimes M_{\ell}.$ (2.6)
The tensor product $\underline{M\mkern-4.0mu}\mkern 4.0mu$ is a ${\mathfrak{g}}[t]$-module via $\underline{{\rm ev}}_{\bf z}$, which coincides with the tensor product of evaluation ${\mathfrak{g}}[t]$-modules
$\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z}):=M_{1}(z_{1})\otimes\cdots\otimes M_{\ell}(z_{\ell}).$ (2.7)
As $\mathfrak{g}$}-modules, $\underline{M\mkern-4.0mu}\mkern 4.0mu$ and $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})$ are isomorphic via the identity map. For any $A\in{\mathfrak{g}}$, the series $A(z)$ acts on $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})$ by
$\hskip 42.67912ptA(z)v=\sum_{i=1}^{\ell}\frac{A^{(i)}v}{z-z_{i}}\qquad\mbox{for $v\in\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})$}.$
2.3. Pseudo-differential operators
Let $\mathcal{A}$ be an associative unital superalgebra and $z$ an even variable. We denote by $\mathcal{A}[\kern-1.1pt[z]\kern-1.1pt]$ (resp., $\mathcal{A}(\!(z)\!)$) the superalgebra of formal power series (resp., formal Laurent series) in $z$ with coefficients in $\mathcal{A}$. Let $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ denote the set of all formal series of the form
$\sum_{j=-\infty}^{s}\sum_{i=-\infty}^{r}a_{ij}z^{i}\partial_{z}^{j},$
where $r,s\in{\mathbb{Z}}$ and $a_{ij}\in\mathcal{A}$. We may endow $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ with a superalgebra structure using the rules:
$\partial_{z}\partial_{z}^{-1}=\partial_{z}^{-1}\partial_{z}=1,\quad\partial_{z}^{i}z^{j}=\sum_{k=0}^{\infty}\binom{i}{k}\binom{j}{k}k!\,z^{j-k}\,\partial_{z}^{i-k},\quad\mbox{for $i,j\in{\mathbb{Z}}$.}$ (2.8)
Here, e.g., ${\binom{i}{k}:=\frac{i(i-1)\ldots(i-k+1)}{k!}}$. The superalgebra $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ is called the superalgebra of pseudo-differential operators over $\mathcal{A}$.
Let $\omega$ be an anti-involution on $\mathcal{A}$. In particular, $\omega^{2}=1$ and $\omega(ab)=\omega(b)\omega(a)$ for $a,b\in\mathcal{A}$. We readily see that $\omega$ extends to a $\mathbb{C}$}-linear automorphism $\omega:\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ defined by
$\omega(az^{i}\partial_{z}^{j})=\omega(a)z^{j}\partial_{z}^{i}$ (2.9)
for $a\in\mathcal{A}$ and $i,j\in{\mathbb{Z}}$. It is straightforward to prove the following.
Proposition 2.1.
The map $\omega$ is an anti-involution on $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$.
3. Bethe algebras for classical Lie algebras
In this section, we examine the behavior of the Feigin–Frenkel center and the Bethe algebra for the classical Lie algebra $\mathfrak{g}$} of type $\mathfrak{a,c,d}$ under a *-structure. Our primary purpose is to study the diagonalization of the Bethe algebra for $\mathfrak{g}$} on a tensor product of unitarizable $\mathfrak{g}$}-modules.
3.1. Feigin–Frenkel centers
Let $t$ be an even variable, and let $\mathfrak{g}$} be a simple Lie algebra or the general linear Lie (super)algebra. We denote by $V_{{\rm crit}}({\mathfrak{g}})$ the vacuum module at the critical level over the central extension $\widehat{{\mathfrak{g}}}:={\mathfrak{g}}[t,t^{-1}]\oplus{\mathbb{C}}K$ of the loop algebra ${\mathfrak{g}}[t,t^{-1}]$. By the Poincaré–Birkhoff–Witt theorem, $V_{{\rm crit}}({\mathfrak{g}})$ can be identified with $U(t^{-1}{\mathfrak{g}}[t^{-1}])$ as (super)spaces. Moreover, the $\widehat{{\mathfrak{g}}}-$module structure on $V_{{\rm crit}}({\mathfrak{g}})$ allows us to define a vertex algebra structure on $V_{{\rm crit}}({\mathfrak{g}})$, called the universal affine vertex algebra at the critical level (see [Fr, FBZ, Mo2, MR] for details).
[…]
${\mathfrak{z}}(\widehat{{\mathfrak{g}}})=\left\{\,\!v\in V_{{\rm crit}}({\mathfrak{g}})\;\middle|\;{\mathfrak{g}}[t]v=0\!\,\right\}.$
[…]
${\rm T}(1)=0\qquad\text{and}\qquad{\rm T}(A[-r])=rA[-r-1]$ (3.1)
[…]
The following result is a famous theorem of Feigin and Frenkel.
Theorem 3.1 ([FF]).
[…]
Remark 3.2.
Theorem 3.1 holds as well for the general linear Lie algebra (see [CM, Theorem 3.1] (and also [CT]).
Let ${\mathfrak{g}}={\mathfrak{gl}}_{d}$, $\mathfrak{sp}_{2d}$, or $\mathfrak{so}_{2d}$, which has rank $d$. For ${\mathfrak{x}}=\mathfrak{a,c,d}$, recall the elements $e^{{\mathfrak{x}}}_{ij}$ defined in (2.1), (2.2) and (2.3), respectively. We denote by $\mathfrak{g}$$_{\mathbb{R}}$ the $\mathbb{R}$}-subspace of $\mathfrak{g}$} spanned by all the $e^{\mathfrak{x}}_{ij}$ and by $\mathfrak{h}$$_{\mathbb{R}}$ the $\mathbb{R}$}-subspace of $\mathfrak{h}$} spanned by all the $e^{\mathfrak{x}}_{ii}$.
Let $U({\mathfrak{g}}[t])_{\mathbb{R}}$ be the $\mathbb{R}$}-subalgebra of $U({\mathfrak{g}}[t])$ generated by $A[r]$ for $A\in{\mathfrak{g}}_{\mathbb{R}}$ and $r\in{\mathbb{Z}}_{+}$, and let $U(t^{-1}{\mathfrak{g}}[t^{-1}])_{\mathbb{R}}$ be the $\mathbb{R}$}-subalgebra of $U(t^{-1}{\mathfrak{g}}[t^{-1}])$ generated by A[r]A[-r] for $A\in{\mathfrak{g}}_{\mathbb{R}}$ and $r\in{\mathbb{N}}$. The $\mathbb{R}$}-algebras $U({\mathfrak{h}}[t])_{\mathbb{R}}$ and $U(t^{-1}{\mathfrak{h}}[t^{-1}])_{\mathbb{R}}$ are defined analogously.
By [CM, Theorem 3.1] for type $\mathfrak{a}$ and [Mo3, Theorem 2.3] for type $\mathfrak{c,d}$ (see also [Mo1, Ya]), there exist
[…]
3.2. Bethe algebras
Let $\mathfrak{g}$} be a simple Lie algebra or the general linear Lie algebra. For any $\mu\in{\mathfrak{g}}^{*}$, there is an algebra homomorphism
$\Psi^{\mu}:U(t^{-1}{\mathfrak{g}}[t^{-1}])\longrightarrow U({\mathfrak{g}}[t])[\kern-1.1pt[z^{-1}]\kern-1.1pt]$
given by
$\Psi^{\mu}(A[-r])=A\otimes(t-z)^{-r}+\delta_{r,1}\mu(A),\qquad\mbox{for $A\in{\mathfrak{g}}\,$ and $\,r\in{\mathbb{N}}$.}$ (3.3)
[…]
$\underline{{\rm ev}}_{\bf z}:U({\mathfrak{g}}[t])[\kern-1.1pt[z^{-1}]\kern-1.1pt]\longrightarrow U({\mathfrak{g}})^{\otimes\ell}[\kern-1.1pt[z^{-1}]\kern-1.1pt],$
such that $\underline{{\rm ev}}_{\bf z}(z^{-1})=z^{-1}$. We have the commutative diagram
$U(t^{-1}{\mathfrak{g}}[t^{-1}])$$U({\mathfrak{g}}[t])[\kern-1.1pt[z^{-1}]\kern-1.1pt]$$U({\mathfrak{g}})^{\otimes\ell}[\kern-1.1pt[z^{-1}]\kern-1.1pt]$$\scriptstyle{\Psi^{\mu}_{\bf z}}$$\scriptstyle{\Psi^{\mu}}$$\scriptstyle{\underline{{\rm ev}}_{\bf z}}$
where $\Psi^{\mu}_{\bf z}$ is defined by
$\Psi^{\mu}_{\bf z}(A[-r])=\sum_{i=1}^{\ell}\frac{A^{(i)}}{(z_{i}-z)^{r}}+\delta_{r,1}\mu(A),\qquad\mbox{for $A\in{\mathfrak{g}}\,$ and $\,r\in{\mathbb{N}}$.}$
[…]
$\big(\mathcal{A}_{{\mathfrak{g}}}^{\mu}({\bf z})\big)_{\underline{M\mkern-4.0mu}\mkern 4.0mu}=\big(\mathcal{B}_{\mathfrak{g}}^{\mu}\big)_{\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})}.$
In this paper, we will restrict our attention to $\mathcal{B}_{\mathfrak{g}}^{\mu}$.
For $\mu\in{\mathfrak{g}}^{*}$, we denote by ${\mathfrak{g}}^{\mu}$ the centralizer of $\mu$ in $\mathfrak{g}$}. Moreover, $\mu$ is called regular if ${\mathfrak{g}}^{\mu}$ has dimension equal to the rank of $\mathfrak{g}$}. Set
${\bm{X}}_{\!\ell}=\left\{\,\!(z_{1},\ldots,z_{\ell})\in{\mathbb{C}}^{\ell}\;\middle|\;z_{i}\not=z_{j}\,\,\mbox{for any $i\not=j$}\!\,\right\},$
called the configuration space of $\ell$ distinct points on $\mathbb{C}$$^{\ell}$.
Let $M_{1},\ldots,M_{\ell}$ be $\mathfrak{g}$}-modules, and let ${\bf z}\in{\bm{X}}_{\!\ell}$. If ${\mathfrak{h}}\subseteq{\mathfrak{g}}^{\mu}$, then $\mathcal{B}_{\mathfrak{g}}^{\mu}$ commutes with $\mathfrak{h}$} and hence acts on the $\gamma$-weight space $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})_{\gamma}$ of $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})$ for any $\gamma\in{\mathfrak{h}}^{*}$ ([Ry1, Proposition 4]).
Let $V_{1},\ldots,V_{\ell}$ be finite-dimensional irreducible $\mathfrak{g}$}-modules. We have the following remarkable results about the action of $\mathcal{B}_{\mathfrak{g}}^{\mu}$ on $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$.
Theorem 3.3 ([FFRy, Corollary 5]).
The space $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$ is a cyclic $\mathcal{B}_{\mathfrak{g}}^{\mu}$-module for any regular $\mu\in{\mathfrak{g}}^{*}$ and ${\bf z}\in{\bm{X}}_{\!\ell}$.
Theorem 3.4 ([FFRy, Corollary 6]).
The Bethe algebra $\mathcal{B}_{\mathfrak{g}}^{\mu}$ is diagonalizable with a simple spectrum on $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$ for generic $\mu\in{\mathfrak{g}}^{*}$ and generic ${\bf z}\in{\bm{X}}_{\!\ell}$.
$3.$ *-structures
A *-superalgebra is a pair $(\mathcal{A},{\bm{\sigma}})$, where $\mathcal{A}$ is an associative superalgebra, and ${\bm{\sigma}}:\mathcal{A}\longrightarrow\mathcal{A}$ is an even anti-linear anti-involution, called a *-structure on $\mathcal{A}$. A homomorphism $f:(\mathcal{A},{\bm{\sigma}})\longrightarrow(\mathcal{A}^{\prime},{\bm{\sigma}}^{\prime})$ of *-superalgebras is a homomorphism of superalgebras satisfying ${\bm{\sigma}}^{\prime}\circ f=f\circ{\bm{\sigma}}$. Let $(\mathcal{A},{\bm{\sigma}})$ be a *-superalgebra, and let $V$ be an $\mathcal{A}$-module. A Hermitian form $\langle\cdot|\cdot\rangle$ on $V$ is said to be contravariant if $\langle av_{1}|v_{2}\rangle=\langle v_{1}|{\bm{\sigma}}(a)v_{2}\rangle$ for all $a\in\mathcal{A}$ and $v_{1},v_{2}\in V$. An $\mathcal{A}$-module equipped with a positive definite contravariant Hermitian form is called a unitarizable $\mathcal{A}$-module.
A Lie superalgebra with *-structure is a pair $(\mathcal{G},{\bm{\sigma}})$, where $\mathcal{G}$ is a Lie superalgebra, and ${\bm{\sigma}}:\mathcal{G}\longrightarrow\mathcal{G}$ is an even anti-linear anti-involution, called a *-structure on $\mathcal{G}$. Note that $\bm{\sigma}$} is a *-structure on $\mathcal{G}$ if and only if the natural extension of $\bm{\sigma}$} to $U(\mathcal{G})$ is a *-structure on $U(\mathcal{G})$. A homomorphism $f:(\mathcal{G},{\bm{\sigma}})\longrightarrow(\mathcal{G}^{\prime},{\bm{\sigma}}^{\prime})$ of Lie superalgebras with *-structures is a homomorphism of Lie superalgebras satisfying ${\bm{\sigma}}^{\prime}\circ f=f\circ{\bm{\sigma}}$.
Let $(\mathcal{G},{\bm{\sigma}})$ be a Lie superalgebra with *-structure, and let $V$ be a $\mathcal{G}$-module. A Hermitian form $\langle\cdot|\cdot\rangle$ on $V$ is said to be contravariant if $\langle Av_{1}|v_{2}\rangle=\langle v_{1}|{\bm{\sigma}}(A)v_{1}\rangle$ for all $A\in\mathcal{G}$ and $v_{1},v_{2}\in V$. A $\mathcal{G}$-module equipped with a positive definite contravariant Hermitian form is called a unitarizable $\mathcal{G}$-module. It is clear that all unitarizable $\mathcal{G}$-modules are unitarizable $U(\mathcal{G})$-modules, and vice versa, and that all unitarizable $\mathcal{G}$-modules are completely reducible. We refer the reader to [ChL1, CLZ, LZ] for related discussions.
The Lie superalgebra ${\mathfrak{gl}}_{p+m|q+n}$ admits a *-structure ${\bm{\sigma}}_{p}:{\mathfrak{gl}}_{p+m|q+n}\longrightarrow{\mathfrak{gl}}_{p+m|q+n}$ (see [CLZ, Section 3.2]), defined by
[…]
where $\bar{c}_{ij}$ denotes the complex conjugate of $c_{ij}\in{\mathbb{C}}$, and
$i]:=\begin{cases}0&\mbox{if}\ \ 1\leq i\leq p,\\ 1&\mbox{otherwise}.\end{cases}$
After specializing to $q=n=0$, we obtain a *-structure
${\bm{\sigma}}_{p}:{\mathfrak{gl}}_{p+m}\longrightarrow{\mathfrak{gl}}_{p+m}.$ (3.5)
Taking $p=m=d$, we obtain a *-structure ${\bm{\sigma}}_{d}$ on ${\mathfrak{gl}}_{2d}$. Evidently, the map ${\bm{\sigma}}_{d}$ restricts to *-structures
${\bm{\sigma}}_{d}:\mathfrak{sp}_{2d}\longrightarrow\mathfrak{sp}_{2d}\qquad\text{and}\qquad{\bm{\sigma}}_{d}:\mathfrak{so}_{2d}\longrightarrow\mathfrak{so}_{2d}.$ (3.6)
We remark that the Lie algebras with *-structures $({\mathfrak{gl}}_{p+m},{\bm{\sigma}}_{p})$, $(\mathfrak{sp}_{2d},{\bm{\sigma}}_{d})$, and $(\mathfrak{so}_{2d},{\bm{\sigma}}_{d})$ correspond to the Hermitian symmetric spaces of the noncompact type referred to as HS.1, HS.3, and HS.5, respectively, in [ES, p. 37].
For the rest of this section, we let ${\mathfrak{g}}={\mathfrak{gl}}_{p+m}$, $\mathfrak{sp}_{2d}$, or $\mathfrak{so}_{2d}$. Set
${\bm{\sigma}}=\begin{cases}{\bm{\sigma}}_{p}&\quad\mbox{if ${\mathfrak{g}}={\mathfrak{gl}}_{p+m}$,}\\ {\bm{\sigma}}_{d}&\quad\mbox{if ${\mathfrak{g}}=\mathfrak{sp}_{2d}$ or $\mathfrak{so}_{2d}$}.\end{cases}$ (3.7)
For ${\mathfrak{x}}=\mathfrak{a,c,d}$, we clearly have ${\bm{\sigma}}(e^{{\mathfrak{x}}}_{ij})=(-1)^{[i]+[j]}e^{{\mathfrak{x}}}_{ji}$ for all $i,j$. In particular, ${\bm{\sigma}}(e^{{\mathfrak{x}}}_{ii})=e^{{\mathfrak{x}}}_{ii}$ for all $i$. Note that every non-trivial unitarizable highest weight $\mathfrak{g}$}-module with respect to $\bm{\sigma}$} is infinite-dimensional (see [EHW]).
Let $k$ denote the dimension of $\mathfrak{g}$}. Consider the Segal–Sugawara vector
[…]
where $\{v_{1},\ldots,v_{k}\}$ is a basis for $\mathfrak{g}$}, and $\{v^{1},\ldots,v^{k}\}$ is the dual basis with respect to a nondegenerate invariant symmetric bilinear form on $\mathfrak{g}$}.
Theorem 3.5 ([Ry2, Theorem 1]).
[…]
We denote by $U(t^{-1}{\mathfrak{g}}[t^{-1}])^{\mathfrak{h}}$ the $\mathfrak{h}$}-centralizer of $U(t^{-1}{\mathfrak{g}}[t^{-1}])$. It admits a direct sum decomposition
$U(t^{-1}{\mathfrak{g}}[t^{-1}])^{\mathfrak{h}}=U(t^{-1}{\mathfrak{h}}[t^{-1}])\oplus\mathcal{N},$ (3.8)
where
$\mathcal{N}$ $:=$ $U(t^{-1}{\mathfrak{g}}[t^{-1}])^{\mathfrak{h}}\cap U(t^{-1}{\mathfrak{g}}[t^{-1}])t^{-1}\mathfrak{n}_{-}[t^{-1}]$ (3.9)
$\hskip 2.84526pt=$ $U(t^{-1}{\mathfrak{g}}[t^{-1}])^{\mathfrak{h}}\cap t^{-1}\mathfrak{n}_{+}[t^{-1}]U(t^{-1}{\mathfrak{g}}[t^{-1}]).$ (3.10)
The projection of $U(t^{-1}{\mathfrak{g}}[t^{-1}])^{\mathfrak{h}}$ onto $U(t^{-1}{\mathfrak{h}}[t^{-1}])$ yields an affine version of the Harish-Chandra homomorphism
$\mathfrak{f}:U(t^{-1}{\mathfrak{g}}[t^{-1}])^{\mathfrak{h}}\longrightarrow U(t^{-1}{\mathfrak{h}}[t^{-1}]).$
[…]
$\mathfrak{f}:{\mathfrak{z}}(\widehat{{\mathfrak{g}}})\longrightarrow\mathcal{W}(^{L}{\mathfrak{g}})$
is an isomorphism (see [Fr, Theorem 8.1.5] and [Mo2, Theorem 13.1.1]). It is called the Feigin–Frenkel isomorphism.
The *-structure $\bm{\sigma}$} on $\mathfrak{g}$}, defined in (3.7), extends naturally to a *-structure ${\bm{\sigma}}:U({\mathfrak{g}})\longrightarrow U({\mathfrak{g}})$, which gives rise to a *-structure ${\bm{\sigma}}:U({\mathfrak{g}}[t,t^{-1}])\longrightarrow U({\mathfrak{g}}[t,t^{-1}])$, defined by
$\hskip 42.67912pt{\bm{\sigma}}(A[r])={\bm{\sigma}}(A)[r]\qquad\mbox{for $A\in{\mathfrak{g}}$ and $r\in{\mathbb{Z}}$}.$
Evidently, $\bm{\sigma}$} restricts to *-structures on $U(t^{-1}{\mathfrak{g}}[t^{-1}])$ and on $U({\mathfrak{g}}[t])$.
[…]
We have the following proposition, which is proved using the same strategy as the proof of [Lu, Proposition 2.4].
Proposition 3.6.
The *-structure ${\bm{\sigma}}:U(t^{-1}{\mathfrak{g}}[t^{-1}])\longrightarrow U(t^{-1}{\mathfrak{g}}[t^{-1}])$ fixes $S_{1},\ldots,S_{d}$.
Proof.
Let $S\in\{S_{1},\ldots,S_{d}\}$. Since ${\bm{\sigma}}\big(S_{0}\big)=S_{0}$, we have
${\bm{\sigma}}(S)S_{0}={\bm{\sigma}}(S){\bm{\sigma}}(S_{0})={\bm{\sigma}}(S_{0}S)={\bm{\sigma}}(SS_{0})={\bm{\sigma}}(S_{0}){\bm{\sigma}}(S)=S_{0}{\bm{\sigma}}(S).$
[…]
Recall that $S_{1},\ldots,S_{d}\in U(t^{-1}{\mathfrak{g}}[t^{-1}])_{\mathbb{R}}$. According to (3.8), $S=H+N$ for some $H\in U(t^{-1}{\mathfrak{h}}[t^{-1}])_{\mathbb{R}}$ and $N\in\mathcal{N}$. Clearly, ${\bm{\sigma}}(H)=H$. By (3.9) and (3.10), ${\bm{\sigma}}(N)\in\mathcal{N}$. Thus, $\mathfrak{f}({\bm{\sigma}}(S))=H$, and hence $\mathfrak{f}({\bm{\sigma}}(S))=\mathfrak{f}(S)$. We have ${\bm{\sigma}}(S)=S$ by the injectivity of $\mathfrak{f}:{\mathfrak{z}}(\widehat{{\mathfrak{g}}})\longrightarrow\mathcal{W}(^{L}{\mathfrak{g}})$. ∎
Recall the map $\Psi^{\mu}:U(t^{-1}{\mathfrak{g}}[t^{-1}])\longrightarrow U({\mathfrak{g}}[t])[\kern-1.1pt[z^{-1}]\kern-1.1pt]$ defined by (3.3). For $\mu\in{\mathfrak{g}}^{*}$, let
$S^{\mu}_{i}(z)=\Psi^{\mu}(S_{i})\qquad\mbox{for $i=1,\ldots,d$.}$ (3.11)
We define $\overline{\mu}\in{\mathfrak{g}}^{*}$ by $\overline{\mu}(A)=\overline{\mu(A)}$ for $A\in{\mathfrak{g}}$. That is, $\overline{\mu}$ is the complex conjugation of $\mu$. Set
${\mathfrak{g}}^{*}_{{\bm{\sigma}}}=\left\{\,\mu\in{\mathfrak{g}}^{*}\;\middle|\;\mu\circ{\bm{\sigma}}=\overline{\mu}\,\right\}.$
Proposition 3.7.
For any $\mu\in{\mathfrak{g}}^{*}_{{\bm{\sigma}}}$, the *-structure ${\bm{\sigma}}:U({\mathfrak{g}}[t])\longrightarrow U({\mathfrak{g}}[t])$ fixes all coefficients of the series $S^{\mu}_{1}(z),\ldots,S^{\mu}_{d}(z)$.
Proof.
We may extend ${\bm{\sigma}}:U({\mathfrak{g}}[t])\longrightarrow U({\mathfrak{g}}[t])$ to a *-structure ${\bm{\sigma}}:U({\mathfrak{g}}[t])[\kern-1.1pt[z^{-1}]\kern-1.1pt]\longrightarrow U({\mathfrak{g}}[t])[\kern-1.1pt[z^{-1}]\kern-1.1pt]$ with ${\bm{\sigma}}(z^{-1})=z^{-1}$. For any $\mu\in{\mathfrak{g}}^{*}_{{\bm{\sigma}}}$, the diagram
[…]
commutes. This, together with Proposition 3.6, yields ${\bm{\sigma}}\big(S^{\mu}_{i}(z)\big)=S^{\mu}_{i}(z)$ for $i=1,\ldots,d$. Thus, the proposition follows. ∎
3.4. Unitarizable modules and diagonalization
[…]
Proposition 3.8.
For any $\mu\in{\mathfrak{g}}^{*}$, the Bethe algebra $\mathcal{B}_{\mathfrak{g}}^{\mu}$ equals the subalgebra of $U({\mathfrak{g}}[t])$ generated by the coefficients of $S^{\mu}_{i}(z)$ for $i=1,\ldots,d$.
Let $M_{1},\ldots,M_{\ell}$ be unitarizable $\mathfrak{g}$}-modules (with respect to the *-structure $\bm{\sigma}$}). By definition, $M_{i}$ possesses a positive definite contravariant Hermitian form $\langle\cdot|\cdot\rangle_{i}$ for each $i=1,\ldots,\ell$. Let $\langle\cdot|\cdot\rangle$ denote the tensor form of $\langle\cdot|\cdot\rangle_{1},\ldots,\langle\cdot|\cdot\rangle_{\ell}$. It is given by
$\big\langle v_{1}\otimes\ldots\otimes v_{\ell}\,|\,v^{\prime}_{1}\otimes\ldots\otimes v^{\prime}_{\ell}\big\rangle=\prod_{i=1}^{\ell}\big\langle v_{i}\,|\,v^{\prime}_{i}\big\rangle_{i}$
for $v_{i},v^{\prime}_{i}\in M_{i}$, $1\leq i\leq\ell$.
Lemma 3.9.
Let $\mu\in{\mathfrak{g}}^{*}_{{\bm{\sigma}}}$ and $\bf z$$\in{\mathbb{R}}^{\ell}$. For $i=1,\ldots,d$, let $\bm{b}$} be any coefficient of the series $S^{\mu}_{i}(z)$. Then $\bm{b}$} is a Hermitian operator on $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})$ with respect to $\langle\cdot|\cdot\rangle$.
Proof.
Using the contravariance of $\langle\cdot|\cdot\rangle_{1},\ldots,\langle\cdot|\cdot\rangle_{\ell}$ and the hypothesis that $\bf z$$\in{\mathbb{R}}^{\ell}$, we can check directly that
$\big\langle A_{1}[r_{1}]\ldots A_{k}[r_{k}]v\,|\,v^{\prime}\big\rangle\hskip-5.69054pt$ $=$ $\hskip-5.69054pt\big\langle v\,|\,{\bm{\sigma}}(A_{k})[r_{k}]\ldots{\bm{\sigma}}(A_{1})[r_{1}]v^{\prime}\big\rangle$
$=$ $\hskip-5.69054pt\big\langle v\,|\,{\bm{\sigma}}\big(A_{1}[r_{1}]\ldots A_{k}[r_{k}]\big)v^{\prime}\big\rangle$
for $v,v^{\prime}\in\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})$, $A_{1},\ldots,A_{k}\in{\mathfrak{g}}$, $r_{1},\ldots,r_{k}\in{\mathbb{Z}}_{+}$, and $k\in{\mathbb{N}}$. This gives
$\big\langle{\bm{b}}v\,|\,v^{\prime}\big\rangle=\big\langle v\,|\,{\bm{\sigma}}({\bm{b}})v^{\prime}\big\rangle=\big\langle v\,|\,{\bm{b}}v^{\prime}\big\rangle,$
where the first equality is due to the sesquilinearity of $\langle\cdot|\cdot\rangle$ and the anti-linearity of $\bm{\sigma}$}, and the second equality follows from Proposition 3.7. This proves the lemma. ∎
Theorem 3.10.
For any finite-dimensional $\mathcal{B}_{\mathfrak{g}}^{\mu}$-submodule $V$ of $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})$, the Bethe algebra $\mathcal{B}_{\mathfrak{g}}^{\mu}$ is diagonalizable on $V$ for any $\mu\in{\mathfrak{g}}^{*}_{{\bm{\sigma}}}$ and $\bf z$$\in{\mathbb{R}}^{\ell}$.
Proof.
For $i=1,\ldots,\ell$, let $\bm{b}$} be any coefficient of the series $S^{\mu}_{i}(z)$. By Lemma 3.9, $\bm{b}$} is a Hermitian operator on $V$ and hence is diagonalizable on $V$. This proves the theorem in view of Proposition 3.8. ∎
We expect the following conjecture to be true.
Conjecture 3.11.
For any finite-dimensional $\mathcal{B}_{\mathfrak{g}}^{\mu}$-submodule $V$ of $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})$, the Bethe algebra $\mathcal{B}_{\mathfrak{g}}^{\mu}$ is diagonalizable on $V$ for generic $\mu\in{\mathfrak{g}}^{*}$ and generic $\bf z$$\in{\mathbb{C}}^{\ell}$.
Let ${\mathfrak{g}}=\mathfrak{n}_{-}\oplus{\mathfrak{h}}\oplus\mathfrak{n}_{+}$ be the triangular decomposition in (2.4). For any ${\bf w}:=(w_{1},\ldots,w_{d})\in{\mathbb{C}}^{d}$, we define $\mu_{\bf w}\in{\mathfrak{g}}^{*}$ by
$\mu_{\bf w}(e^{{\mathfrak{x}}}_{ii})=-w_{i},\quad\mbox{for $1\leq i\leq d$;}$ (3.12)
$\mu_{\bf w}$ vanishes on $\mathfrak{n}_{-}\oplus\mathfrak{n}_{+}$. (3.13)
That is, $\mu_{\bf w}$ corresponds to a diagonal matrix. The choice of the minus sign will be clear in Section 4.2. Set
$\mathcal{B}_{\mathfrak{g}}^{\bf w}=\mathcal{B}_{\mathfrak{g}}^{\mu_{\bf w}}.$
It is evident that ${\mathfrak{h}}\subseteq{\mathfrak{g}}^{\mu_{\bf w}}$. Thus for any weight $\gamma\in{\mathfrak{h}}^{*}$ of $\underline{M\mkern-4.0mu}\mkern 4.0mu$, the weight space $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})_{\gamma}$ is a $\mathcal{B}_{\mathfrak{g}}^{\mu}$-submodule of $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})$.
Corollary 3.12.
Let $L_{1},\ldots,L_{\ell}$ be unitarizable highest weight $\mathfrak{g}$}-modules (with respect to $\bm{\sigma}$}), and let $\gamma\in{\mathfrak{h}}^{*}$ be any weight of $\underline{L}$. Then $\mathcal{B}_{\mathfrak{g}}^{\bf w}$ is diagonalizable on $\underline{L}({\bf z})_{\gamma}$ for any ${\bf w}\in{\mathbb{R}}^{d}$ and $\bf z$$\in{\mathbb{R}}^{\ell}$.
Proof.
Since $\underline{L}({\bf z})_{\gamma}$ is finite-dimensional and $\mu_{\bf w}\in{\mathfrak{g}}^{*}_{{\bm{\sigma}}}$ for ${\bf w}\in{\mathbb{R}}^{d}$, the corollary follows from Theorem 3.10. ∎
Theorem 3.3 and Theorem 3.4 are statements for finite-dimensional irreducible $\mathfrak{g}$}-modules. Their analogs for unitarizable highest weight $\mathfrak{g}$}-modules (with respect to $\bm{\sigma}$}) are conjectured to be true.
Conjecture 3.13.
Let $\underline{L}$ and $\gamma$ be as in Corollary 3.12. Then the weight space $\underline{L}({\bf z})_{\gamma}$ is a cyclic $\mathcal{B}_{\mathfrak{g}}^{\bf w}$-module for any ${\bf w}\in{\bm{X}}_{\!d}$ and ${\bf z}\in{\bm{X}}_{\!\ell}$. Moreover, $\mathcal{B}_{\mathfrak{g}}^{\bf w}$ is diagonalizable with a simple spectrum on $\underline{L}({\bf z})_{\gamma}$ for generic $\bf w$} and generic $\bf z$}.
Positive evidence for Conjecture 3.13 (and hence for Conjecture 3.11), in the case where $\mathfrak{g}$} is the general linear Lie algebra, will be given in Corollary 4.17.
3.5. Frobenius algebras
A Frobenius algebra is a finite-dimensional commutative associative unital algebra $\mathcal{A}$ which admits a nondegenerate symmetric bilinear form $(\cdot,\cdot)$ such that
$\hskip 28.45274pt(a_{1}a_{2},a_{3})=(a_{1},a_{2}a_{3})\quad\mbox{for all $a_{1},a_{2},a_{3}\in\mathcal{A}$.}$
Let $V_{1},\ldots,V_{\ell}$ be finite-dimensional irreducible $\mathfrak{g}$}-modules, and let ${\bf z}\in{\bm{X}}_{\!\ell}$. We will give a brief account of the fact that $(\mathcal{B}_{\mathfrak{g}}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is a Frobenius algebra for ${\bf w}\in{\bm{X}}_{\!d}$ (cf. [MTV1, Lu]).
For ${\mathfrak{x}}=\mathfrak{a,c,d}$, we define an anti-involution $\varpi:{\mathfrak{g}}\longrightarrow{\mathfrak{g}}$ by
$\varpi(e^{\mathfrak{x}}_{ij})=e^{\mathfrak{x}}_{ji}\qquad\mbox{for all $i,j$.}$ (3.14)
The map $\varpi$ is called the Cartan anti-involution on $\mathfrak{g}$}.
For $i=1,\ldots,\ell$, $V_{i}$ is an irreducible highest weight $\mathfrak{g}$}-module with a dominant integral highest weight. Let $v_{i}$ be a highest weight vector of $V_{i}$. There is a unique nondegenerate symmetric bilinear form $(\cdot,\cdot)_{i}$ on $V_{i}$ defined by
$(v_{i},v_{i})_{i}=1\quad\text{and}\quad(Av,v^{\prime})_{i}=(v,\varpi(A)v^{\prime})_{i}$
for all $v,v^{\prime}\in V_{i}$ and $A\in{\mathfrak{g}}$ ([Sh]). The bilinear form $(\cdot,\cdot)_{i}$ is called the Shapovalov form on $V_{i}$. It is easy to see that if $v$ and $v^{\prime}$ are weight vectors of $V_{i}$ of distinct weights, then $(v,v^{\prime})_{i}=0$.
Let $(\cdot,\cdot)$ denote the tensor form of the bilinear forms $(\cdot,\cdot)_{1},\ldots,(\cdot,\cdot)_{\ell}$, which is called the tensor Shapovalov form. It is a nondegenerate symmetric bilinear form on $\underline{V\mkern-4.0mu}\mkern 4.0mu$. Let ${\bf w}\in{\mathbb{C}}^{d}$. By [MTV1, Theorem 9.1] and [Lu, Lemma 2.6], the Bethe algebra $\mathcal{B}_{\mathfrak{g}}^{\bf w}$ is symmetric with respect to $(\cdot,\cdot)$, i.e.,
$\hskip 42.67912pt({\bm{b}}v,v^{\prime})=(v,{\bm{b}}v^{\prime})\quad\mbox{for ${\bm{b}}\in\mathcal{B}_{\mathfrak{g}}^{\bf w}\,$ and $\,v,v^{\prime}\in\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$}.$
Theorem 3.14.
For any ${\bf w}\in{\bm{X}}_{\!d}$ and ${\bf z}\in{\bm{X}}_{\!\ell}$, the Bethe algebra $(\mathcal{B}_{\mathfrak{g}}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is a Frobenius algebra. Moreover, $(\mathcal{B}_{\mathfrak{g}}^{\bf w})_{{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}_{\gamma}}$ is a Frobenius algebra for any weight $\gamma$ of $\underline{V\mkern-4.0mu}\mkern 4.0mu$.
Proof.
The first statement follows from the above discussion on the tensor Shapovalov form $(\cdot,\cdot)$ on $\underline{V\mkern-4.0mu}\mkern 4.0mu$ in view of Theorem 3.3 and [Lu, Lemma 2.7]. As $(\cdot,\cdot)$ restricts to a nondegenerate symmetric bilinear form on ${\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}_{\!\gamma}$ for any weight $\gamma$ of $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$, the second statement follows. ∎
The following is a consequence of [Lu, Lemma 1.3].
Theorem 3.15.
For any ${\bf w}\in{\bm{X}}_{\!d}$ and ${\bf z}\in{\bm{X}}_{\!\ell}$, the following properties hold:
  1. (i)
    The Bethe algebra $(\mathcal{B}_{\mathfrak{g}}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is a maximal commutative subalgebra of ${\rm End}(\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z}))$ of dimension $\dim\!\left({\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}\right)$.
  2. (ii)
    Every eigenspace of $(\mathcal{B}_{\mathfrak{g}}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is one-dimensional, and the set of eigenspaces of $(\mathcal{B}_{\mathfrak{g}}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is in bijective correspondence with the set of maximal ideals of $(\mathcal{B}_{\mathfrak{g}}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$.
  3. (iii)
    Every generalized eigenspace of $(\mathcal{B}_{\mathfrak{g}}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is a cyclic $\mathcal{B}_{\mathfrak{g}}^{\bf w}$-module.
Remark 3.16.
Each statement of Theorem 3.15 holds if $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$ is replaced with ${\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}_{\gamma}$ for any weight $\mu$ of $\underline{V\mkern-4.0mu}\mkern 4.0mu$.
4. A Duality of Bethe algebras and an application
In this section, we establish a duality of Bethe algebras for the general linear Lie (super)algebras ${\mathfrak{gl}}_{d}$ and ${\mathfrak{gl}}_{p+m|q+n}$. We apply it to study the action of the Bethe algebra for ${\mathfrak{gl}}_{p+m|q+n}$ on a tensor product of infinite-dimensional unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-modules and obtain evidence for Conjecture 3.13 if $q$ and $n$ are set to 0.
4.1. Column determinants and Berezinians
Let us review the notions of column determinants and Berezinians.
Let $\mathcal{A}$ be an associative unital superalgebra. The parity of a homogeneous element $a\in\mathcal{A}$ is denoted by $|a|$, which lies in ${\mathbb{Z}}_{2}$. Fix $k\in{\mathbb{N}}$. For any $k\times k$ matrix $A=\big[a_{i,j}\big]_{i,j=1,\ldots,k}$ over $\mathcal{A}$, the column determinant of $A$ is defined to be
${\rm cdet}(A)=\sum_{\sigma\in\mathfrak{S}_{k}}(-1)^{l(\sigma)}\,a_{\sigma(1),1}\ldots a_{\sigma(k),k}.$
Here $\mathfrak{S}_{k}$ denotes the symmetric group on $\{1,\ldots,k\}$ and $l(\sigma)$ denotes the length of $\sigma$.
Assume that $A$ has a two-sided inverse $A^{-1}=\big[\widetilde{a}_{i,j}\big]$. For all $i,j=1,\ldots,k$, the $(i,j)$th quasideterminant of $A$ is defined to be $|A|_{ij}:=\widetilde{a}_{j,i}^{-1}$ provided that $\widetilde{a}_{j,i}$ has an inverse in $\mathcal{A}$. Following the notation of [GGRW], it is convenient to write
$|A|_{ij}=\begin{vmatrix}a_{1,1}&\ldots&a_{1,j}&\ldots&a_{1,k}\\ \ldots&\ldots&\ldots&\ldots&\ldots\\ a_{i,1}&\ldots&\framebox{$a_{i,j}$}&\ldots&a_{i,k}\\ \ldots&\ldots&\ldots&\ldots&\ldots\\ a_{k,1}&\ldots&a_{k,j}&\ldots&a_{k,k}\end{vmatrix}$.
For $i=1,\ldots,k$, we define
$d_{i}(A)=\begin{vmatrix}a_{1,1}&\ldots&a_{1,i}\\ \ldots&\ldots&\ldots\\ a_{i,1}&\ldots&\framebox{$a_{i,i}$}\end{vmatrix}$,\
which are called the principal quasiminors of $A$.
Let $A=\big[a_{i,j}\big]_{i,j=1,\ldots,k}$ be a $k\times k$ matrix over $\mathcal{A}$. For any nonempty subset $P=\{i_{1}<\ldots<i_{p}\}$ of $\{1,\ldots,k\}$, the matrix $A_{P}:=\big[a_{i,j}\big]_{i,j\in P}$ is called a standard submatrix of $A$. We say that $A$ is sufficiently invertible if every principal quasiminor of $A$ is well defined, and that $A$ is amply invertible if each of its standard submatrices is sufficiently invertible.
A sequence $(s_{1},\ldots,s_{m+n})$ of 0’s and 1’s is called a $0^{m}1^{n}$-sequence if exactly $m$ of the $s_{i}$’s are 0 and the others are 1. Every $0^{m}1^{n}$-sequence can be written in the form $(0^{m_{1}},1^{n_{1}},\ldots,0^{m_{r}},1^{n_{r}}),$ where the sequence begins with $m_{1}$ copies of 0’s, followed by $n_{1}$ copies of $1$’s, and so on. The set of all $0^{m}1^{n}$-sequences is denoted by $\mathcal{S}_{{m|n}}$.
Fix ${\bf s}:=(s_{1},\ldots,s_{m+n})\in\mathcal{S}_{{m|n}}$. For any $\sigma\in\mathfrak{S}_{m+n}$ and any $(m+n)\times(m+n)$ matrix $A=[a_{i,j}]_{i=1,\ldots,m+n}$ over $\mathcal{A}$, we define ${\bf s}^{\sigma}=\left(s_{\sigma^{-1}(1)},s_{\sigma^{-1}(2)},\ldots,s_{\sigma^{-1}(m+n)}\right)$ and $A^{\sigma}=\big[a_{\sigma^{-1}(i),\sigma^{-1}(j)}\big]_{i=1,\ldots,m+n}$. We say that $A$ is of type $\bf s$} if $a_{i,j}$ is homogeneous of parity $|a_{i,j}|=\bar{s}_{i}+\bar{s}_{j}$ for any $i,j=1,\ldots,m+n$.
For any sufficiently invertible matrix $A$ of type $\bf s$} over $\mathcal{A}$, the Berezinian of type $\bf s$} of $A$ is defined to be
${\rm Ber}^{\bf s}(A)=d_{1}(A)^{\hat{s}_{1}}\ldots d_{m+n}(A)^{\hat{s}_{m+n}},$
where $\hat{s}_{i}:=(-1)^{s_{i}}$ for $i=1,\ldots,m+n$ (see [HM, (3.3)]). We refer the reader to [Ber, Na, MR] for earlier definitions of Berezinians. The following proposition is useful.
Proposition 4.1 (cf. [HM, Proposition 3.5]).
Let $A$ be an amply invertible matrix of type $\bf s$} over $\mathcal{A}$. Fix $k\in\{1,\ldots,m+n-1\}$. We write
$A=\begin{bmatrix}W&X\\ Y&Z\end{bmatrix}$, (4.1)
where $W,X,Y,Z$ are respectively $k\times k$, $k\times(m+n-k)$, $(m+n-k)\times k$, and $(m+n-k)\times(m+n-k)$ matrices. Then $W$ and $Z-YW^{-1}X$ are sufficiently invertible matrices of types ${\bf s}^{\prime}:=(s_{1},\ldots,s_{k})$ and ${\bf s}^{\prime\prime}:=(s_{k+1},\ldots,s_{m+n})$, respectively. Moreover,
${\rm Ber}^{\bf s}(A)={\rm Ber}^{{\bf s}^{\prime}}(W)\cdot{\rm Ber}^{{\bf s}^{\prime\prime}}\big(Z-YW^{-1}X\big).$
A matrix $A=[a_{i,j}]_{i=1,\ldots,m+n}$ over $\mathcal{A}$ is called a Manin matrix of type $\bf s$} if $A$ is a matrix of type $\bf s$} satisfying the following relations
$a_{i,j},a_{k,l}]=(-1)^{s_{i}s_{j}+s_{i}s_{k}+s_{j}s_{k}}[a_{k,j},a_{i,l}]\qquad\text{for all}\quad i,j,k,l=1,\ldots,m+n,$
where $a,b]:=ab-(-1)^{|a||b|}ba$ for any homogeneous elements $a,b\in\mathcal{A}$. We recall some basic facts about Manin matrices.
Proposition 4.2 (cf. [HM, Section 3]).
Let $A$ be a Manin matrix of type $\bf s$} over $\mathcal{A}$. Then
  1. (i)
    If $P=\{i_{1}<\ldots<i_{p}\}$ is a nonempty subset of $\{1,\ldots,m+n\}$, then the standard submatrix $A_{P}$ is a Manin matrix of type $(s_{i_{1}},\ldots s_{i_{p}})$.
  2. (ii)
    For any $\sigma\in\mathfrak{S}_{m+n}$, $A^{\sigma}$ is a Manin matrix of type ${\bf s}^{\sigma}$.
The following proposition asserts that the Berezinian of a Manin matrix is invariant under any permutation.
Proposition 4.3 ([HM, Proposition 3.6]).
Let $A$ be an amply invertible Manin matrix of type $\bf s$} over $\mathcal{A}$. Then
$\hskip 28.45274pt{\rm Ber}^{{\bf s}^{\sigma}}(A^{\sigma})={\rm Ber}^{\bf s}(A)\qquad\mbox{for any $\sigma\in\mathfrak{S}_{m+n}$.}$
We may decompose ${\rm Ber}^{\bf s}(A)$ in Proposition 4.1 in a different way if $A$ is a Manin matrix.
Proposition 4.4.
Let $A$ be an amply invertible Manin matrix of type $\bf s$} over $\mathcal{A}$. Write $A$ as in (4.1). Then $W-XZ^{-1}Y$ and $Z$ are sufficiently invertible matrices of types ${\bf s}^{\prime}:=(s_{1},\ldots,s_{k})$ and ${\bf s}^{\prime\prime}:=(s_{k+1},\ldots,s_{m+n})$, respectively. Moreover,
${\rm Ber}^{\bf s}(A)={\rm Ber}^{{\bf s}^{\prime\prime}}(Z)\cdot{\rm Ber}^{{\bf s}^{\prime}}\big(W-XZ^{-1}Y\big).$ (4.2)
Proof.
The first assertion is clear. Let $\sigma\in\mathfrak{S}_{m+n}$ be the permutation defined by $\sigma(k+i)=i$ for $i=1,\ldots,m+n-k$ and $\sigma(j)=m+n-k+j$ for $j=1,\ldots,k$. Then
$A^{\sigma}=\begin{bmatrix}Z&Y\\ X&W\end{bmatrix}$.
The equality (4.2) is now a direct consequence of Proposition 4.1 and Proposition 4.3. ∎
We also have the following.
Proposition 4.5 ([CFR, Lemma 8]).
Let $A$ be a sufficiently invertible Manin matrix of type ${\bf s}_{0}:=(0^{m})$ over $\mathcal{A}$. Then ${\rm Ber}^{{\bf s}_{0}}(A)={\rm cdet}(A)$.
4.2. The Bethe algebras for ${\mathfrak{gl}}_{d}$ and ${\mathfrak{gl}}_{p+m|q+n}$
Let $t$ be an even variable and
$\tau=-\partial_{t}.$
Here $t$ and $\partial_{t}$ satisfy the rules similar to (2.8). Let ${\bf w}=(w_{1},\ldots,w_{d})\in{\mathbb{C}}^{d}$, and recall $\mu_{\bf w}\in{\mathfrak{gl}}_{d}^{*}$ defined by (3.12) and (3.13). Write
$\Psi^{\bf w}=\Psi^{\mu_{\bf w}}.$
The map $\Psi^{\bf w}$ extends to an algebra homomorphism
$\Psi^{\bf w}:U(t^{-1}{\mathfrak{gl}}_{d}[t^{-1}])[\tau]\longrightarrow U({\mathfrak{gl}}_{d}[t])[\kern-1.1pt[z^{-1}]\kern-1.1pt][\partial_{z}]$
such that $\Psi^{\bf w}(\tau)=\partial_{z}$.
Let
$\mathcal{T}_{d}=\Big[\delta_{a,b}\tau+e_{ab}[-1]\Big]_{a,b=1,\ldots,d},$
which is a Manin matrix of type $(0^{d})$. The vectors $S_{1},\ldots,S_{d}\in{\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d})$, mentioned in (3.2), are the coefficients of ${\rm cdet}(\mathcal{T}_{d})\in U(t^{-1}{\mathfrak{gl}}_{d}[t^{-1}])[\tau]$; more precisely,
${\rm cdet}(\mathcal{T}_{d})=\tau^{d}+\sum_{i=1}^{d}S_{i}\tau^{i-1}$ (4.3)
(see [CM, Theorem 3.1]). Now consider the Manin matrix
${\mathcal{L}}_{d}^{\bf w}:=\Big[\Psi^{\bf w}\big(\delta_{a,b}\tau+e_{ab}[-1]\big)\Big]_{a,b=1,\ldots,d}=\Big[\delta_{a,b}(\partial_{z}-w_{a})-e_{ab}(z)\Big]_{a,b=1,\ldots,d}.$ (4.4)
Then in view of (4.3),
${\rm cdet}({\mathcal{L}}_{d}^{\bf w})=\partial_{z}^{d}+\sum_{i=1}^{d}S_{i}^{{\bf w}}(z)\partial_{z}^{i-1},$ (4.5)
where $S_{i}^{{\bf w}}(z):=\Psi^{\bf w}(S_{i})$. Write
$\mathcal{B}_{d}^{\bf w}=\mathcal{B}_{{\mathfrak{gl}}_{d}}^{\bf w}.$
Recall from Proposition 3.8 that $\mathcal{B}_{d}^{\bf w}$ is equal to the subalgebra of $U({\mathfrak{gl}}_{d}[t])$ generated by the coefficients of $S_{i}^{{\bf w}}(z)$ for $i=1,\ldots,d$.
Fix ${\bf s}:=(0^{p},1^{q},0^{m},1^{n})\in\mathcal{S}_{p+m|q+n}$. Let
$\mathcal{T}_{p+m|q+n}=\Big[\delta_{i,j}\tau+(-1)^{|i|}E^{i}_{j}[-1]\Big]_{i,j\in\mathbb{I}}.$
It is straightforward to verify that $\mathcal{T}_{p+m|q+n}$ is an amply invertible Manin matrix of type $\bf s$} over $U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])(\!(\tau^{-1})\!)$ (cf. [MR, Lemma 3.1]). We have the following expansion:
${\rm Ber}^{\bf s}\!\left(\mathcal{T}_{p+m|q+n}\right)=\sum_{i=-\infty}^{p+m-q-n}b_{i}\tau^{i},$ (4.6)
for some $b_{i}\in U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])$.
Let $\widehat{{\mathfrak{z}}}_{p+m|q+n}$ be the subalgebra of $U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])$ generated by
$\{{\rm T}^{r}b_{i}\,|\,i\leq p+m-q-n,\,i\in{\mathbb{Z}},\,r\in{\mathbb{Z}}_{+}\},$
where $\rm T$} is defined by (3.1). According to [MR, Corollary 3.3] (see also [ChL2, Proposition 3.3]), $b_{i}\in{\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{p+m|q+n})$ for all $i\in{\mathbb{Z}}$ with $i\leq p+m-q-n$. Since ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{p+m|q+n})$ is $\rm T$}-invariant, we see that $\widehat{{\mathfrak{z}}}_{p+m|q+n}$ is a commutative subalgebra of ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{p+m|q+n})$.
There is a conjecture asserting that ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{p+m|q+n})=\widehat{{\mathfrak{z}}}_{p+m|q+n}$ (see [MR]). The conjecture holds if $q=n=0$ (see [CM, Theorem 3.1] and also Theorem 3.1). It is also true that ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{1|1})=\widehat{{\mathfrak{z}}}_{1|1}$ ([MM]) and ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{2|1})=\widehat{{\mathfrak{z}}}_{2|1}$ ([AN]). However, it is still unknown whether the conjecture is valid in general.
Now let ${\bf z}=(z_{1},\ldots,z_{p+q+m+n})\in{\mathbb{C}}^{p+q+m+n}$. Define $\mu_{\bf z}\in{\mathfrak{gl}}_{p+m|q+n}^{*}$ by
$\hskip 28.45274pt\mu_{\bf z}(E^{i}_{j})=(-1)^{|i|+1}\delta_{i,j}z_{i}\qquad\mbox{for $i,j\in\mathbb{I}$.}$
Similar to (3.3), we have a superalgebra homomorphism
$\Psi^{\bf z}:U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])\longrightarrow U({\mathfrak{gl}}_{p+m|q+n}[t])(\!(z^{-1})\!)$
given by
$\Psi^{\bf z}(A[-r])=A\otimes(t-z)^{-r}+\delta_{r,1}\mu_{\bf z}(A),\quad\mbox{for $A\in{\mathfrak{gl}}_{p+m|q+n}\,$ and $\,r\in{\mathbb{N}}$.}$
The map $\Psi^{\bf z}$ extends to a superalgebra homomorphism
$\Psi^{\bf z}:U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])(\!(\tau^{-1})\!)\longrightarrow U({\mathfrak{gl}}_{p+m|q+n}[t])(\!(z^{-1},\partial_{z}^{-1})\!)$
given by
$\Psi^{\bf z}\left(\sum_{i=-\infty}^{s}a_{i}\tau^{i}\right)=\sum_{i=-\infty}^{s}\Psi^{\bf z}(a_{i})\partial_{z}^{i}$
for $a_{i}\in U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])$ and $s\in{\mathbb{Z}}$. Consider the matrix
${\mathcal{L}}_{{p+m|q+n}}^{\bf z}:=\Big[\Psi^{\bf z}\big(\delta_{i,j}\tau+(-1)^{|i|}E^{i}_{j}[-1]\big)\Big]_{i,j\in\mathbb{I}}=\Big[\delta_{i,j}(\partial_{z}-z_{i})-(-1)^{|i|}E^{i}_{j}(z)\Big]_{i,j\in\mathbb{I}}.$ (4.7)
It is an amply invertible Manin matrix of type $\bf s$} over $U({\mathfrak{gl}}_{p+m|q+n}[t])(\!(z^{-1},\partial_{z}^{-1})\!)$. The following is an immediate consequence of Proposition 4.3.
Proposition 4.6.
For each $\sigma\in\mathfrak{S}_{p+q+m+n}$,
${\rm Ber}^{{\bf s}^{\sigma}}\big(\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big)^{\sigma}\big)={\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big).$
By (4.6), we have
${\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big)=\sum_{i=-\infty}^{p+m-q-n}b_{i}^{{\bf z}}(z)\partial_{z}^{i},$ (4.8)
where $b_{i}^{{\bf z}}(z):=\Psi^{\bf z}(b_{i})$. Let ${\mathcal{B}_{p+m|q+n}^{\bf z}}:={\mathcal{B}_{{\mathfrak{gl}}_{p+m|q+n}}^{\bf z}}$ be the subalgebra of $U({\mathfrak{gl}}_{p+m|q+n}[t])$ generated by the coefficients of the series $b_{i}^{{\bf z}}(z)$, for $i\in{\mathbb{Z}}$ with $i\leq p+m-q-n$. It is called the Bethe algebra for ${\mathfrak{gl}}_{p+m|q+n}$ with respect to $\bf z$}. We remark that $\mathcal{B}_{p+m|q+n}^{\bf z}$} coincides with the subalgebra of $U({\mathfrak{gl}}_{p+m|q+n}[t])$ generated by the coefficients of $\Psi^{\bf z}(S)$ for $S\in\widehat{{\mathfrak{z}}}_{p+m|q+n}$ (as $\Psi^{\bf z}({\rm T}(S))=\frac{d}{dz}(\Psi^{\bf z}(S))$). If we take $q=n=0$, we recover the Bethe algebra for ${\mathfrak{gl}}_{p+m}$ by virtue of Proposition 4.5.
According to [MR, Corallary 3.7], the algebra $\mathcal{B}_{p+m|q+n}^{\bf z}$} is a commutative subalgebra of $U({\mathfrak{gl}}_{p+m|q+n}[t])$. It also commutes with the Cartan subalgebra ${\mathfrak{h}}_{{p+m|q+n}}$ of ${\mathfrak{gl}}_{p+m|q+n}$ and hence acts on any weight space of a ${\mathfrak{gl}}_{p+m|q+n}$-module.
For any $\mathcal{B}_{p+m|q+n}^{\bf z}$}-module $V$, let $\big({\mathcal{B}_{p+m|q+n}^{\bf z}}\big)_{V}$ denote the image of the Bethe algebra $\mathcal{B}_{p+m|q+n}^{\bf z}$} in ${\rm End}(V)$. We call $\big({\mathcal{B}_{p+m|q+n}^{\bf z}}\big)_{V}$ the Bethe algebra for $V$ with respect to $\bf z$}.
4.3. The Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$
Let $\mathcal{F}:={\mathbb{C}}[{\bf x},{\bf y},{\bm{\eta}},{\bm{\zeta}}]$ be the polynomial superalgebra generated by $x^{a}_{i}$, $\eta^{a}_{j}$, $y^{a}_{r}$, and $\zeta^{a}_{s}$, for $i=1,\ldots,m$, $j=1,\ldots,n$, $r=1,\ldots,p$, $s=1,\ldots,q$, and $a=1,\ldots,d$. Here $x^{a}_{i}$ and $y^{a}_{r}$ are even variables while $\eta^{a}_{j}$ and $\zeta^{a}_{s}$ are odd. With an appropriate inner product, $\mathcal{F}$ may be regarded as the Fock space of $d(p+m)$ bosonic and $d(q+n)$ fermionic oscillators (cf. [CL, LZ]). Let $\mathcal{D}:={\mathbb{D}}[{\bf x},{\bf y},{\bm{\eta}},{\bm{\zeta}}]$ be the corresponding Weyl superalgebra, i.e., $\mathcal{D}$ is the associative unital superalgebra generated by $x^{a}_{i},\,\eta^{a}_{j},\,y^{a}_{r},\,\zeta^{a}_{s}$ as well as their derivatives
$\partial_{x^{a}_{i}}:=\frac{\partial}{\partial x_{i}^{a}},\quad\partial_{\eta^{a}_{j}}:=\frac{\partial}{\partial\eta_{j}^{a}},\quad\partial_{y^{a}_{r}}:=\frac{\partial}{\partial y_{r}^{a}},\quad\partial_{\zeta^{a}_{s}}:=\frac{\partial}{\partial\zeta_{r}^{a}}$
for $1\leq i\leq m$, $1\leq j\leq n$, $1\leq r\leq p$, $1\leq s\leq q$, and $1\leq a\leq d$. The superalgebra $\mathcal{D}$ acts naturally on $\mathcal{F}$, and so it is a subalgebra of ${\rm End}(\mathcal{F})$.
According to [CLZ, pp. 789–790], there is a superalgebra homomorphism $\phi:U({\mathfrak{gl}}_{d})\longrightarrow\mathcal{D}$, given by
$e_{ab}\mapsto\sum_{i=1}^{m}x_{i}^{a}\partial_{x^{b}_{i}}+\sum_{j=1}^{n}\eta_{j}^{a}\partial_{\eta^{b}_{j}}-\sum_{r=1}^{p}y_{r}^{b}\partial_{y^{a}_{r}}-\sum_{s=1}^{q}\zeta_{s}^{b}\partial_{\zeta^{a}_{s}},$ (4.9)
for $1\leq a,b\leq d$, and a superalgebra homomorphism $\varphi:U({\mathfrak{gl}}_{p+m|q+n})\longrightarrow\mathcal{D}$, given by
$\begin{array}[]{cc}E^{r}_{r^{\prime}}\mapsto-\sum_{a=1}^{d}\partial_{y^{a}_{r}}y_{r^{\prime}}^{a},\quad&E^{r}_{p+s}\mapsto\sum_{a=1}^{d}\partial_{y^{a}_{r}}\zeta_{s}^{a},\\ E^{p+s}_{r}\mapsto-\sum_{a=1}^{d}\partial_{\zeta^{a}_{s}}y_{r}^{a},\quad&E^{p+s}_{p+s^{\prime}}\mapsto\sum_{a=1}^{d}\partial_{\zeta^{a}_{s}}\zeta_{s^{\prime}}^{a},\\ E^{r}_{p+q+i}\mapsto\sum_{a=1}^{d}\partial_{y^{a}_{r}}\partial_{x^{a}_{i}},\quad&E^{r}_{p+q+m+j}\mapsto\sum_{a=1}^{d}\partial_{y^{a}_{r}}\partial_{\eta^{a}_{j}},\\ E^{p+s}_{p+q+i}\mapsto\sum_{a=1}^{d}\partial_{\zeta^{a}_{s}}\partial_{x^{a}_{i}},\quad&E^{p+s}_{p+q+m+j}\mapsto\sum_{a=1}^{d}\partial_{\zeta^{a}_{s}}\partial_{\eta^{a}_{j}},\\ E^{p+q+i}_{r}\mapsto-\sum_{a=1}^{d}x^{a}_{i}y_{r}^{a},\quad&E^{p+q+i}_{p+s}\mapsto\sum_{a=1}^{d}x^{a}_{i}\zeta_{s}^{a},\\ E^{p+q+m+j}_{r}\mapsto-\sum_{a=1}^{d}\eta^{a}_{j}y_{r}^{a},\quad&E^{p+q+m+j}_{p+s}\mapsto\sum_{a=1}^{d}\eta^{a}_{j}\zeta_{s}^{a},\quad\\ E^{p+q+i}_{p+q+i^{\prime}}\mapsto\sum_{a=1}^{d}x_{i}^{a}\partial_{x^{a}_{i^{\prime}}},\quad&E^{p+q+i}_{p+q+m+j}\mapsto\sum_{a=1}^{d}x_{i}^{a}\partial_{\eta^{a}_{j}},\\ E^{p+q+m+j}_{p+q+i}\mapsto\sum_{a=1}^{d}\eta_{j}^{a}\partial_{x^{a}_{i}},\quad&E^{p+q+m+j}_{p+q+m+j^{\prime}}\mapsto\sum_{a=1}^{d}\eta_{j}^{a}\partial_{\eta^{a}_{j^{\prime}}}.\end{array}$} (4.10)
where $1\leq r,r^{\prime}\leq p$, $1\leq s,s^{\prime}\leq q$, $1\leq i,i^{\prime}\leq m$, and $1\leq j,j^{\prime}\leq n$. These maps induce a ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{p+m|q+n}$-action on $\mathcal{F}$.
Let ${\bm{\pi}}:{\mathfrak{gl}}_{d}\longrightarrow{\mathfrak{gl}}_{d}$ be the *-structure on ${\mathfrak{gl}}_{d}$ defined by
${\bm{\pi}}(e_{ab})=e_{ba}\qquad\mbox{for $1\leq a,b\leq d$.}$
Note that any unitarizable highest weight ${\mathfrak{gl}}_{d}$-module with respect to $\bm{\pi}$} is finite-dimensional. Recall the *-structure ${\bm{\sigma}}_{p}$ on ${\mathfrak{gl}}_{p+m|q+n}$ defined by (3.4). The maps $\bm{\pi}$} and ${\bm{\sigma}}_{p}$ give rise to a *-structure $({\bm{\pi}},{\bm{\sigma}}_{p})$ on ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{p+m|q+n}$. It is well-known that $\mathcal{F}$ decomposes into a direct sum of unitarizable ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{p+m|q+n}$-modules (with respect to $({\bm{\pi}},{\bm{\sigma}}_{p})$). This is called the Howe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ ([CLZ]). Such unitarizable modules are, in general, infinite-dimensional. Taking $p=q=0$, the maps $\phi$ and $\varphi$ specialize to the superalgebra homomorphisms $U({\mathfrak{gl}}_{d})\longrightarrow{\mathbb{D}}[{\bf x},{\bm{\eta}}]$ and $U({\mathfrak{gl}}_{m|n})\longrightarrow{\mathbb{D}}[{\bf x},{\bm{\eta}}]$, respectively, and the duality reduces to the Howe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m|n})$, which yields a decomposition of ${\mathbb{C}}[{\bf x},{\bm{\eta}}]$ into a direct sum of finite-dimensional ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{m|n}$-modules ([CW, Se1, Se2]). Assuming further that $n=0$, it gives the Howe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m})$ ([H]).
We simplify notations for the rest of this subsection. Let
$x^{a}_{m+j}=\eta^{a}_{j},\quad\partial_{x^{a}_{m+j}}=\partial_{\eta^{a}_{j}},\quad y^{a}_{p+s}=\xi^{a}_{s},\quad\partial_{y^{a}_{p+s}}=\partial_{\zeta^{a}_{s}},$
where $1\leq j\leq n$ and $1\leq s\leq q$. Fix two sequences
${\bf w}:=(w_{1},\ldots,w_{d})\in{\mathbb{C}}^{d}\qquad\text{and}\qquad{\bf z}:=(z_{1},\ldots,z_{p+q+m+n})\in{\mathbb{C}}^{p+q+m+n}.$
The map $\phi$ induces a superalgebra homomorphism $\phi_{\bf z}:U({\mathfrak{gl}}_{d}[t])\longrightarrow\mathcal{D}$ defined by
$e_{ab}[k]\mapsto-\sum_{r=1}^{p+q}z_{r}^{k}y^{b}_{r}\partial_{y^{a}_{r}}+\sum_{i=1}^{m+n}z_{p+q+i}^{k}x^{a}_{i}\partial_{x^{b}_{i}},$ (4.11)
for $1\leq a,b\leq d$ and $k\in{\mathbb{Z}}_{+}$, and $\varphi$ induces a superalgebra homomorphism $\varphi_{\bf w}:U({\mathfrak{gl}}_{p+m|q+n}[t])\longrightarrow\mathcal{D}$ defined by
$\begin{array}[]{cll}E^{r}_{s}[k]&\mapsto\sum_{a=1}^{d}(-1)^{|s|+1}w_{a}^{k}\partial_{y^{a}_{r}}y^{a}_{s},\quad E^{r}_{p+q+j}[k]&\mapsto\sum_{a=1}^{d}w_{a}^{k}\partial_{y^{a}_{r}}\partial_{x^{a}_{j}},\\ E^{p+q+i}_{s}[k]&\mapsto\sum_{a=1}^{d}(-1)^{|s|+1}w_{a}^{k}x^{a}_{i}y^{a}_{s},\quad E^{p+q+i}_{p+q+j}[k]&\mapsto\sum_{a=1}^{d}w_{a}^{k}x^{a}_{i}\partial_{x^{a}_{j}},\end{array}$} (4.12)
for $1\leq r,s\leq p+q$, $1\leq i,j\leq m+n$, and $k\in{\mathbb{Z}}_{+}$. Since the Bethe algebras $\mathcal{B}_{d}^{\bf w}$ and $\mathcal{B}_{p+m|q+n}^{\bf z}$} are subalgebras of $U({\mathfrak{gl}}_{d}[t])$ and $U({\mathfrak{gl}}_{p+m|q+n}[t])$, respectively, the maps $\phi_{\bf z}$ and $\varphi_{\bf w}$ induce an action of $\mathcal{B}_{d}^{\bf w}$ on $\mathcal{F}$ and an action of $\mathcal{B}_{p+m|q+n}^{\bf z}$} on $\mathcal{F}$, respectively.
In view of the Howe duality, it is natural to ask if any duality exists in the context of Bethe algebras as well. When $p=q=n=0$, a duality phenomenon is noticed in [TL] (see also [TV]). In [MTV3], the precise statement is made and proved. So far, various generalizations have emerged (see [HM, TU, VY]). We will show that the desired duality exists between $\mathcal{B}_{d}^{\bf w}$ and $\mathcal{B}_{p+m|q+n}^{\bf z}$} in a moment.
The maps $\phi_{\bf z}$ and $\varphi_{\bf w}$ extend naturally to the superalgebra homomorphisms
$\phi_{\bf z}:U({\mathfrak{gl}}_{d}[t])(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!)$
and
$\varphi_{\bf w}:U({\mathfrak{gl}}_{p+m|q+n}[t])(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!),$
respectively. Furthermore, there is an anti-involution $\omega:\mathcal{D}\longrightarrow\mathcal{D}$ defined by
$x^{a}_{i}\mapsto\partial_{x^{a}_{i}},\quad\partial_{x^{a}_{i}}\mapsto x^{a}_{i},\quad y^{a}_{r}\mapsto\partial_{y^{a}_{r}},\quad\partial_{y^{a}_{r}}\mapsto y^{a}_{r},$ (4.13)
for $1\leq i\leq m+n$, $1\leq r\leq p+q$, and $1\leq a\leq d$. In light of Proposition 2.1, the map $\omega$ extends naturally to an anti-involution
$\omega:\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!).$
For $u=z$ or $\partial_{z}$, let
$R_{p|q}(u)=\frac{(u-z_{1})\ldots(u-z_{p})}{(u-z_{p+1})\ldots(u-z_{p+q})}$
and
$R^{\prime}_{m|n}(u)=\frac{(u-z_{p+q+1})\ldots(u-z_{p+q+m})}{(u-z_{p+q+m+1})\ldots(u-z_{p+q+m+n})}.$
We have the following proposition. We defer its proof until Appendix A to avoid disrupting the flow of the paper.
Proposition 4.7.
Let ${\bf s}=(0^{p},1^{q},0^{m},1^{n})\in\mathcal{S}_{p+m|q+n}$. Then
$(z-w_{1})\ldots(z-w_{d})\,\varphi_{\bf w}{\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big)=\omega\left(\!R^{\prime}_{m|n}(z)\cdot\phi_{\bf z}{\rm cdet}({\mathcal{L}}_{d}^{\bf w})\cdot R_{p|q}(z)\!\right).$
The Cartan anti-involution $\varpi$ on ${\mathfrak{gl}}_{d}$, defined by (3.14), gives rise to an anti-involution $\varpi:U({\mathfrak{gl}}_{d}[t])\longrightarrow U({\mathfrak{gl}}_{d}[t])$, defined by
$\hskip 42.67912pt\varpi(A[k])=\varpi(A)[k]\qquad\mbox{for $A\in{\mathfrak{gl}}_{d}$ and $k\in{\mathbb{Z}}_{+}$}.$
Proposition 4.8 ([MTV1, Proposition 8.4]).
The anti-involution $\varpi:U({\mathfrak{gl}}_{d}[t])\longrightarrow U({\mathfrak{gl}}_{d}[t])$ fixes all elements of $\mathcal{B}_{d}^{\bf w}$.
Proposition 4.9.
For any ${\bm{b}}\in\mathcal{B}_{d}^{\bf w}$, $\omega\big(\phi_{\bf z}({\bm{b}})\big)=\phi_{\bf z}({\bm{b}})$.
Proof.
It is straightforward to verify that the diagram
[…]
commutes. This, together with Proposition 4.8, proves the proposition. ∎
The following theorem is called the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$. It not only relates the coefficients of $\phi_{\bf z}{\rm cdet}({\mathcal{L}}_{d}^{\bf w})$ to those of $\varphi_{\bf w}{\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big)$, but also gives an equality of algebras.
Theorem 4.10.
Let ${\bf s}=(0^{p},1^{q},0^{m},1^{n})\in\mathcal{S}_{p+m|q+n}$. We may write
$R_{p|q}(z)\cdot R^{\prime}_{m|n}(z)\cdot\phi_{\bf z}{\rm cdet}({\mathcal{L}}_{d}^{\bf w})=\sum_{\ell=0}^{d}\sum_{k=-\infty}^{p+m-q-n}b_{k,\ell}^{{\bf w},{\bf z}}z^{k}\partial_{z}^{\ell},$
where $b_{k,\ell}^{{\bf w},{\bf z}}:=\phi_{\bf z}(b_{k,\ell}^{\bf w})$ for some $b_{k,\ell}^{\bf w}\in\mathcal{B}_{d}^{\bf w}$. Then
$(z-w_{1})\ldots(z-w_{d})\ \varphi_{\bf w}{\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big)$
$\hskip 85.35826pt=R_{p|q}(\partial_{z})\left(\sum_{k=-\infty}^{p+m-q-n}\sum_{\ell=0}^{d}b_{k,\ell}^{{\bf w},{\bf z}}z^{\ell}\partial_{z}^{k}\right)R_{p|q}(\partial_{z})^{-1}.$
Moreover, $\phi_{\bf z}(\mathcal{B}_{d}^{\bf w})=\varphi_{\bf w}\big({\mathcal{B}_{p+m|q+n}^{\bf z}}\big)$.
Proof.
By Proposition 4.7,
$(z-w_{1})\ldots(z-w_{d})\,\varphi_{\bf w}{\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big)$
$\hskip 42.67912pt=R_{p|q}(\partial_{z})\cdot\omega\left(R^{\prime}_{m|n}(z)\cdot\phi_{\bf z}{\rm cdet}({\mathcal{L}}_{d}^{\bf w})\right)\!$
$\hskip 42.67912pt=R_{p|q}(\partial_{z})\cdot\omega\left(R_{p|q}(z)\cdot R^{\prime}_{m|n}(z)\cdot\phi_{\bf z}{\rm cdet}({\mathcal{L}}_{d}^{\bf w})\right)\cdot R_{p|q}(\partial_{z})^{-1}$
$\hskip 42.67912pt=R_{p|q}(\partial_{z})\left(\sum_{k=-\infty}^{p+m-q-n}\sum_{\ell=0}^{d}\omega\big(b_{k,\ell}^{{\bf w},{\bf z}}\big)z^{\ell}\partial_{z}^{k}\right)R_{p|q}(\partial_{z})^{-1}.$
By Proposition 4.9, $\omega\big(b_{k,\ell}^{{\bf w},{\bf z}}\big)=b_{k,\ell}^{{\bf w},{\bf z}}$ for all $k\leq p+m-q-n$ and $0\leq\ell\leq d$. This proves the first assertion. The last assertion clearly follows from the first one. ∎
Remark 4.11.
We have learned from an anonymous expert that the last assertion of Theorem 4.10 can be obtained directly from the results of [HM]. An outline of his approach is given as follows. Let $\pi$ be an automorphism on $\mathcal{D}$ defined by
$x^{a}_{i}\mapsto x^{a}_{i},\quad\partial_{x^{a}_{i}}\mapsto\partial_{x^{a}_{i}},\quad y^{a}_{r}\mapsto(-1)^{|r|+1}\partial_{y^{a}_{r}},\quad\partial_{y^{a}_{r}}\mapsto y^{a}_{r},$
for $1\leq i\leq m+n$, $1\leq r\leq p+q$, and $1\leq a\leq d$, and let $\vartheta$ be an automorphism on $U({\mathfrak{gl}}_{d}[t])$ defined by
$e_{ab}[k]\mapsto e_{ab}[k]-\delta_{a,b}\sum_{r=1}^{p+q}(-1)^{|r|}z_{r}^{k}$
for $1\leq a,b\leq d$ and $k\in{\mathbb{Z}}_{+}$. Note that $\vartheta$ restricts to an automorphism on $\mathcal{B}_{d}^{\bf w}$. As a consequence of [HM, Theorem 5.2],
$(\pi\circ\phi_{\bf z}\circ\vartheta)(\mathcal{B}_{d}^{\bf w})=(\pi\circ\varphi_{\bf w})({\mathcal{B}_{p+m|q+n}^{\bf z}}),$
which is equivalent to the last assertion of Theorem 4.10.
4.4. An application of the Bethe duality
The Weyl superalgebra $\mathcal{D}$ is a subalgebra of ${\rm End}(\mathcal{F})$. For ${\bf w}\in{\mathbb{C}}^{d}$ and ${\bf z}\in{\mathbb{C}}^{p+q+m+n}$, Theorem 4.10 implies that the image of $\mathcal{B}_{d}^{\bf w}$ in ${\rm End}(\mathcal{F})$ under $\phi_{\bf z}$ coincides with the image of $\mathcal{B}_{p+m|q+n}^{\bf z}$} in ${\rm End}(\mathcal{F})$ under $\varphi_{\bf w}$. This makes it possible to study the action of $\mathcal{B}_{p+m|q+n}^{\bf z}$} on $\mathcal{F}$ via the action of $\mathcal{B}_{d}^{\bf w}$ on $\mathcal{F}$, and vice versa.
Let
${\bf d}_{\mathcal{F}}=\sum_{a=1}^{d}\left(\!\sum_{i=1}^{m}x_{i}^{a}\partial_{x^{a}_{i}}+\sum_{j=1}^{n}\eta_{j}^{a}\partial_{\eta^{a}_{j}}-\sum_{r=1}^{p}y_{r}^{a}\partial_{y^{a}_{r}}-\sum_{s=1}^{q}\zeta_{s}^{a}\partial_{\zeta^{a}_{s}}\!\right)\!.$
For any monomial $f\in\mathcal{F}$, we have ${\bf d}_{\mathcal{F}}(f)=kf$ for some $k\in{\mathbb{Z}}$. We define the degree of $f$ to be $k$. In particular, the degrees of $x^{a}_{i}$} and $\eta^{a}_{j}$} are $1$ while the degree of $y^{a}_{r}$ and $\zeta^{a}_{s}$ are 1-1. We call $\bf d$$_{\mathcal{F}}$ the degree operator on $\mathcal{F}$.
For $1\leq r\leq p$, $1\leq s\leq q$, $1\leq i\leq m$, and $1\leq j\leq n$, let $\mathcal{F}_{(r)}={\mathbb{C}}\big[y^{1}_{r},\ldots,y^{d}_{r}\big]$, $\mathcal{F}_{(p+s)}={\mathbb{C}}\big[\zeta^{1}_{s},\ldots,\zeta^{d}_{s}\big]$, $\mathcal{F}_{(p+q+i)}={\mathbb{C}}\big[x^{1}_{i},\ldots,x^{d}_{i}\big]$, and $\mathcal{F}_{(p+q+m+j)}={\mathbb{C}}\big[\eta^{1}_{j},\ldots,\eta^{d}_{j}\big]$. It is evident that as ${\mathfrak{gl}}_{d}$-modules,
$\mathcal{F}=\mathcal{F}_{(1)}\otimes\cdots\otimes\mathcal{F}_{(p+q+m+n)}.$ (4.14)
For $1\leq a\leq d$, let $\mathcal{F}^{(a)}={\mathbb{C}}\big[x^{a}_{i},\eta^{a}_{j},y^{a}_{r},\zeta^{a}_{s}\,\big|\,1\leq i\leq m,1\leq j\leq n,1\leq r\leq p,1\leq s\leq q\big]$. As ${\mathfrak{gl}}_{p+m|q+n}$-modules,
$\mathcal{F}=\mathcal{F}^{(1)}\otimes\cdots\otimes\mathcal{F}^{(d)}.$ (4.15)
For any weight $\mu\in{\mathfrak{h}}_{{p+m|q+n}}^{*}$ of a ${\mathfrak{gl}}_{p+m|q+n}$-module $M$, we denote by $M_{\mu}$ the $\mu$-weight space of $M$. The Cartan subalgebra ${\mathfrak{h}}_{{p+m|q+n}}$ of ${\mathfrak{gl}}_{p+m|q+n}$ acts on $\mathcal{F}$ as follows (see (4.10)): $E^{i}_{i}$ acts on $\mathcal{F}$ by $\sum_{a=1}^{d}y^{a}_{i}\partial_{y^{a}_{i}}-d,$ for $1\leq i\leq p$, and by $\sum_{a=1}^{d}\zeta^{a}_{i-p}\partial_{\zeta^{a}_{i-p}}+d,$ for $p+1\leq i\leq q$. Also, $E^{p+q+i}_{p+q+i}$ acts on $\mathcal{F}$ by $\sum_{a=1}^{d}x^{a}_{i}\partial_{x^{a}_{i}},$ for $1\leq i\leq m$, and by $\sum_{a=1}^{d}\eta^{a}_{i-m}\partial_{\eta^{a}_{i-m}},$ for $m+1\leq i\leq m+n$. Clearly, there is a weight space decomposition
$\mathcal{F}=\bigoplus_{\mu\in{\mathfrak{h}}_{{p+m|q+n}}^{*}}\mathcal{F}_{\mu}.$ (4.16)
For $\xi\in{\mathfrak{h}}_{{p+m|q+n}}^{*}$, we denote by $L_{p+m|q+n}(\xi)$ the irreducible highest weight ${\mathfrak{gl}}_{p+m|q+n}$-module with highest weight $\xi$ with respect to the Borel subalgebra $\mathfrak{b}_{{p+m|q+n}}$. Let
$\mathbbm{1}_{p|q}=\sum_{i=1}^{p}\epsilon_{i}-\sum_{j=1}^{q}\epsilon_{p+j}\qquad\text{and}\qquad\mathbbm{1}_{p}=\mathbbm{1}_{p|0}.$
Suppose $p\not=0\not=m$. For $r\in{\mathbb{Z}}$, let
[…]
The space $V_{r}$ is an infinite-dimensional unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-module (with respect to ${\bm{\sigma}}_{p}$), and the highest weight of $V_{r}$, appearing in (4.17), corresponds to a generalized partition of depth 1; see [CCL1, Section 2.5] and also [CLZ]. (Such a partition is called a generalized partition of length 1 in [CLZ].)
For $a=1,\ldots,d$, let $\gamma_{a}\in{\mathbb{Z}}$. If $\gamma_{a}\geq 0$, then we may identify $V_{\gamma_{a}}$ with the ${\mathfrak{gl}}_{p+m|q+n}$-submodule of the polynomial superalgebra $\mathcal{F}^{(a)}$ generated by the highest weight vector $(x^{a}_{1})^{\gamma_{a}}$. If $\gamma_{a}<0$, then we may identify $V_{\gamma_{a}}$ with the ${\mathfrak{gl}}_{p+m|q+n}$-submodule of $\mathcal{F}^{(a)}$ generated by the highest weight vector
$\begin{cases}\zeta^{a}_{q+\gamma_{a}+1}\ldots\zeta^{a}_{q}&\quad\text{if }\mbox{$\gamma_{a}\geq-q$ and $q\not=0$,}\\ (y^{a}_{p})^{-\gamma_{a}-q}\zeta^{a}_{1}\ldots\zeta^{a}_{q}&\quad\text{if }\mbox{$\gamma_{a}<-q$ and $q\not=0$,}\\ (y^{a}_{p})^{-\gamma_{a}}&\quad\text{if }q=0.\end{cases}$
In any case, $V_{\gamma_{a}}$ is the subspace of $\mathcal{F}^{(a)}$ spanned by all monomials of degree $\gamma_{a}$. Via (4.15), the tensor product $V_{\gamma_{1}}\otimes\cdots\otimes V_{\gamma_{d}}$ is a ${\mathfrak{gl}}_{p+m|q+n}$-submodule of $\mathcal{F}$. We thus obtain a direct sum decomposition of ${\mathfrak{gl}}_{p+m|q+n}$-modules:
$\mathcal{F}=\bigoplus_{\gamma_{1},\ldots,\gamma_{d}\in{\mathbb{Z}}}V_{\gamma_{1}}\otimes\cdots\otimes V_{\gamma_{d}}.$ (4.18)
Let ${\mathfrak{h}}_{d}$ denote the standard Cartan subalgebra of ${\mathfrak{gl}}_{d}$. The action of ${\mathfrak{h}}_{d}$ on $\mathcal{F}$ is determined by the action of $e_{aa}$ on $\mathcal{F}$, which is given by
$\sum_{i=1}^{m}x_{i}^{a}\partial_{x^{a}_{i}}+\sum_{j=1}^{n}\eta_{j}^{a}\partial_{\eta^{a}_{j}}-\sum_{r=1}^{p}y_{r}^{a}\partial_{y^{a}_{r}}-\sum_{s=1}^{q}\zeta_{s}^{a}\partial_{\zeta^{a}_{s}},$
for $1\leq a\leq d$ (see (4.9)). There is a weight space decomposition
$\mathcal{F}=\bigoplus_{\gamma\in{\mathfrak{h}}_{d}^{*}}\mathcal{F}_{\gamma}.$ (4.19)
For $1\leq i\leq p+q$ and $k\in-{\mathbb{Z}}_{+}$, let $\mathcal{F}_{(i)}^{k}$ be the subspace of $\mathcal{F}_{(i)}$ spanned by all monomials of degree $k$. For $1\leq j\leq m+n$ and $l\in{\mathbb{Z}}_{+}$, let $\mathcal{F}_{(p+q+j)}^{l}$ be the subspace of $\mathcal{F}_{(p+q+j)}$ spanned by all monomials of degree $l$. Evidently, $\mathcal{F}_{(i)}^{k}$ and $\mathcal{F}_{(p+q+j)}^{l}$ are finite-dimensional ${\mathfrak{gl}}_{d}$-modules. We have the following direct sum decompositions of ${\mathfrak{gl}}_{d}$-modules:
$\mathcal{F}_{(i)}=\bigoplus_{k\in-{\mathbb{Z}}_{+}}\mathcal{F}_{(i)}^{k}\qquad\text{and}\qquad\mathcal{F}_{(p+q+j)}=\bigoplus_{l\in{\mathbb{Z}}_{+}}\mathcal{F}_{(p+q+j)}^{l}.$
It is well known that $\mathcal{F}_{(i)}^{k}$ and $\mathcal{F}_{(p+q+j)}^{l}$ are irreducible ${\mathfrak{gl}}_{d}$-modules for $k\in-{\mathbb{Z}}_{+}$ and $l\in{\mathbb{Z}}_{+}$ (see, for instance, [CLZ, Theorem 3.3] by setting exactly one of $p,q,m,n$ (given there) to 1 and the others to 0). By (4.14), $\mathcal{F}$ decomposes into a direct sum of ${\mathfrak{gl}}_{d}$-modules
$\mathcal{F}=\bigoplus_{\begin{subarray}{c}k_{1},\ldots,k_{p+q}\in-{\mathbb{Z}}_{+},\\ k_{p+q+1},\ldots,k_{p+q+m+n}\in{\mathbb{Z}}_{+}\end{subarray}}\mathcal{F}_{(1)}^{k_{1}}\otimes\cdots\otimes\mathcal{F}_{({p+q+m+n})}^{k_{p+q+m+n}}.$ (4.20)
Proposition 4.12.
For $\gamma_{1},\ldots,\gamma_{d}\in{\mathbb{Z}}$, let $\gamma=\sum_{a=1}^{d}\gamma_{a}\epsilon_{a}\in{\mathfrak{h}}_{d}^{*}$. Let $\mu\in{\mathfrak{h}}_{{p+m|q+n}}^{*}$ be any weight of $V_{\gamma_{1}}\otimes\cdots\otimes V_{\gamma_{d}}$. Then
$(V_{\gamma_{1}}\otimes\cdots\otimes V_{\gamma_{d}})_{\mu}=\big(\mathcal{F}_{(1)}^{\overline{\mu}_{1}}\otimes\cdots\otimes\mathcal{F}_{({p+q+m+n})}^{\overline{\mu}_{p+q+m+n}}\big)_{\gamma},$
where
$\overline{\mu}_{i}=\begin{cases}\,\mu(E^{i}_{i})+(-1)^{|i|}d&\mbox{if}\ \ 1\leq i\leq p+q;\\ \,\mu(E^{i}_{i})\ \ &\mbox{if}\ \ p+q+1\leq i\leq p+q+m+n.\end{cases}$ (4.21)
Proof.
By (4.18) and (4.19), we have
$\mathcal{F}_{\gamma}=V_{\gamma_{1}}\otimes\cdots\otimes V_{\gamma_{d}}.$
Also, (4.16) and (4.20) yield
$\mathcal{F}_{\mu}=\mathcal{F}_{(1)}^{\overline{\mu}_{1}}\otimes\cdots\otimes\mathcal{F}_{({p+q+m+n})}^{\overline{\mu}_{p+q+m+n}}.$
Consequently,
$(V_{\gamma_{1}}\otimes\cdots\otimes V_{\gamma_{d}})_{\mu}=\mathcal{F}_{\gamma}\cap\mathcal{F}_{\mu}=\big(\mathcal{F}_{(1)}^{\overline{\mu}_{1}}\otimes\cdots\otimes\mathcal{F}_{\overline{\mu}_{p+q+m+n}}^{(p+q+m+n)}\big)_{\gamma},$
as asserted. ∎
For ${\mathfrak{gl}}_{p+m|q+n}$-modules $M_{1},\ldots,M_{d}$ and ${\bf w}=(w_{1},\ldots,w_{d})\in{\mathbb{C}}^{d}$, we write
$\underline{M\mkern-4.0mu}\mkern 4.0mu=M_{1}\otimes\cdots\otimes M_{d}\qquad\text{and}\qquad\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})=M_{1}(w_{1})\otimes\cdots\otimes M_{d}(w_{d}).$
Theorem 4.13.
Suppose $p\not=0\not=m$. For $a=1,\ldots,d$, let $L_{a}=V_{\gamma_{a}}$ for some $\gamma_{a}\in{\mathbb{Z}}$. For any weight $\mu\in{\mathfrak{h}}_{{p+m|q+n}}^{*}$ of $\underline{L}$, ${\bf z}\in{\bm{X}}_{\!p+q+m+n}$, and ${\bf w}\in{\bm{X}}_{\!d}$, we have:
  1. (i)
    $\underline{L}({\bf w})_{\mu}$ is a cyclic $\mathcal{B}_{p+m|q+n}^{\bf z}$}-module.
  2. (ii)
    $\big({\mathcal{B}_{p+m|q+n}^{\bf z}}\big)_{\underline{L}({\bf w})_{\mu}}$ is a Frobenius algebra.
  3. (iii)
    $\big({\mathcal{B}_{p+m|q+n}^{\bf z}}\big)_{\underline{L}({\bf w})_{\mu}}$ is a maximal commutative subalgebra of ${\rm End}(\underline{L}({\bf w})_{\mu})$ of dimension $\dim\!\left(\underline{L}({\bf w})_{\mu}\right)$.
  4. (iv)
    Every eigenspace of $\big({\mathcal{B}_{p+m|q+n}^{\bf z}}\big)_{\underline{L}({\bf w})_{\mu}}$ is one-dimensional.
Moreover, $\mathcal{B}_{p+m|q+n}^{\bf z}$} is diagonalizable with a simple spectrum on $\underline{L}({\bf w})_{\mu}$ for generic $\bf z$} and generic $\bf w$}.
Proof.
Let $\mu\in{\mathfrak{h}}_{{p+m|q+n}}^{*}$ be a weight of $\underline{L}$. For $i\in\mathbb{I}$, let $N_{i}=\mathcal{F}_{(i)}^{\overline{\mu}_{i}}$, where $\overline{\mu}_{i}$ is defined as in (4.21). Write ${\bf z}=(z_{1},\ldots,z_{p+q+m+n})$. We set $\underline{N\mkern-4.0mu}\mkern 4.0mu=N_{1}\otimes\cdots\otimes N_{p+q+m+n}$ and $\underline{N\mkern-4.0mu}\mkern 4.0mu({\bf z})=N_{1}(z_{1})\otimes\cdots\otimes N_{p+q+m+n}(z_{p+q+m+n})$. In view of Theorem 4.10 and Proposition 4.12,
$\big({\mathcal{B}_{p+m|q+n}^{\bf z}}\big)_{\underline{L}({\bf w})_{\mu}}=(\mathcal{B}_{d}^{\bf w})_{\underline{N\mkern-4.0mu}\mkern 4.0mu({\bf z})_{\gamma}}.$ (4.22)
By Theorem 3.3, $\underline{N\mkern-4.0mu}\mkern 4.0mu({\bf z})_{\gamma}$ is a cyclic $\mathcal{B}_{d}^{\bf w}$-module as $\mathcal{B}_{d}^{\bf w}$ preserves weight spaces, and by Theorem 3.14, $(\mathcal{B}_{d}^{\bf w})_{\underline{N\mkern-4.0mu}\mkern 4.0mu({\bf z})_{\gamma}}$ is a Frobenius algebra. Hence, (i) and (ii) follow from (4.22). Also, (iii) and (iv) are immediate consequences of [Lu, Lemma 1.3] (cf. Remark 3.16). The last statement follows from (4.22) and Theorem 3.4. ∎
Remark 4.14.
A similar statement corresponding to $p=0$ or $m=0$ can be formulated. The details will be left to the reader.
Theorem 4.13 suggests that the super analog of Conjecture 3.13 should be true for the general linear Lie superalgebra.
Conjecture 4.15.
Suppose $p$, $q$, $m$, and $n$ are not all zero. For $a=1,\ldots,d$, let $L_{a}$ be a unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-module (with respect to ${\bm{\sigma}}_{p}$). For any weight $\mu\in{\mathfrak{h}}_{{p+m|q+n}}^{*}$ of $\underline{L}$, ${\bf z}\in{\bm{X}}_{\!p+q+m+n}$, and ${\bf w}\in{\bm{X}}_{\!d}$, parts (i)–(iv) of Theorem 4.13 hold for the Bethe algebra $({\mathcal{B}_{p+m|q+n}^{\bf z}})_{\underline{L}({\bf w})_{\mu}}$. Moreover, $\mathcal{B}_{p+m|q+n}^{\bf z}$} is diagonalizable with a simple spectrum on $\underline{L}({\bf w})_{\mu}$ for generic $\bf z$} and generic $\bf w$}.
Remark 4.16.
An analog of Conjecture 4.15 when the entries of $\bf z$} are all zero is studied in [CCL2].
Suppose $p\not=0\not=m$. For $r\in{\mathbb{Z}}$, let
$W_{r}=\begin{cases}L_{p+m|0}(r\epsilon_{p+1}-\mathbbm{1}_{p})&\quad\mbox{if $r\geq 0$,}\\ L_{p+m|0}(r\epsilon_{p}-\mathbbm{1}_{p})&\quad\mbox{if $r<0$}.\end{cases}$
The following provides evidence for Conjecture 3.13.
Corollary 4.17.
For $a=1,\ldots,d$, let $L_{a}=W_{\gamma_{a}}$ for some $\gamma_{a}\in{\mathbb{Z}}$. For any weight $\mu\in{\mathfrak{h}}_{p+m}^{*}$ of $\underline{L}$, ${\bf z}\in{\bm{X}}_{\!p+m}$, and ${\bf w}\in{\bm{X}}_{\!d}$, parts (i)–(iv) of Theorem 4.13 hold for the Bethe algebra $(\mathcal{B}_{p+m}^{\bf z})_{\underline{L}({\bf w})_{\mu}}$. Moreover, $\mathcal{B}_{p+m}^{\bf z}$ is diagonalizable with a simple spectrum on $\underline{L}({\bf w})_{\mu}$ for generic $\bf z$} and generic $\bf w$}.
Proof.
This is an immediate consequence of Theorem 4.13 by taking $q=n=0$. ∎
Acknowledgments. The first author was partially supported by NSTC grants 113-2115-M-006-010 and 114-2115-M-006-005. The second author was partially supported by NSTC grant 112-2115-M-006-015-MY2. The authors are grateful to an anonymous expert for valuable comments on Theorem 4.10 (see Remark 4.11).
Appendix A The proof of Proposition 4.7
The appendix is devoted to proving Proposition 4.7. We keep the notations defined in Section 4. Let
$\mathcal{C}=\left[\delta_{a,b}(z-w_{a})+\sum_{r=1}^{p+q}\frac{y^{a}_{r}\partial_{y^{b}_{r}}}{\partial_{z}-z_{r}}-\sum_{i=1}^{m+n}\frac{x^{b}_{i}\partial_{x^{a}_{i}}}{\partial_{z}-z_{p+q+i}}\right]_{a,b=1,\ldots,d}.$
Lemma A.1.
$\omega\big(\phi_{\bf z}{\rm cdet}({\mathcal{L}}_{d}^{\bf w})\big)={\rm cdet}(\mathcal{C})$.
Proof.
For $1\leq a,b\leq d$,
$\phi_{\bf z}\big(e_{ab}(z)\big)=-\sum_{r=1}^{p+q}\frac{y^{b}_{r}\partial_{y^{a}_{r}}}{z-z_{r}}+\sum_{i=1}^{m+n}\frac{x^{a}_{i}\partial_{x^{b}_{i}}}{z-z_{p+q+i}}.$ (A.1)
For $i\in\mathbb{I}\cup\{0\}$ and $1\leq a,b\leq d$, we define
$F^{i}_{a,b}=\begin{cases}\,\,\delta_{a,b}(\partial_{z}-w_{a})&\mbox{if }\,i=0,\\ \displaystyle{\,\,\frac{y^{b}_{i}\partial_{y^{a}_{i}}}{z-z_{i}}}&\mbox{if }\,1\leq i\leq p+q,\\ \displaystyle{\,\frac{-x^{a}_{i-p-q}\partial_{x^{b}_{i-p-q}}}{z-z_{i}}}&\mbox{otherwise},\end{cases}$
and
$G^{i}_{a,b}=\begin{cases}\,\,\delta_{a,b}(z-w_{a})&\mbox{if }\,i=0,\\ \displaystyle{\,\,\frac{y^{a}_{i}\partial_{y^{b}_{i}}}{\partial_{z}-z_{i}}}&\mbox{if }\,1\leq i\leq p+q,\\ \displaystyle{\,\frac{-x^{b}_{i-p-q}\partial_{x^{a}_{i-p-q}}}{\partial_{z}-z_{i}}}&\mbox{otherwise}.\end{cases}$
We see that
$\omega(F^{i}_{a,b})=G^{i}_{a,b}.$ (A.2)
Note that ${\mathcal{L}}_{d}^{\bf w}$ is an amply invertible Manin matrix of type $(0^{d})$. Let $\pi\in\mathfrak{S}_{d}$ be the permutation given by $\pi(i)=d+1-i$ for $i=1,\ldots,d$. Clearly, $\pi^{2}=1$. By (A.1), Proposition 4.5 and Proposition 4.6, we obtain
$\phi_{\bf z}{\rm cdet}({\mathcal{L}}_{d}^{\bf w})$ $=$ $\phi_{\bf z}{\rm cdet}\big(({\mathcal{L}}_{d}^{\bf w})^{\pi}\big)$
$=$ $\sum_{\sigma\in\mathfrak{S}_{d}}\sum_{i_{1},\ldots,i_{d}=0}^{p+q+m+n}(-1)^{l(\sigma)}F^{i_{d}}_{\pi(\sigma(1)),\pi(1)}\ldots F^{i_{1}}_{\pi(\sigma(d)),\pi(d)}$
$=$ $\sum_{\sigma\in\mathfrak{S}_{d}}\sum_{i_{1},\ldots,i_{d}=0}^{p+q+m+n}(-1)^{l(\sigma)}F^{i_{d}}_{\pi\sigma(1),d}\ldots F^{i_{1}}_{\pi\sigma(d),1}$
$=$ $\sum_{\sigma\in\mathfrak{S}_{d}}\sum_{i_{1},\ldots,i_{d}=0}^{p+q+m+n}(-1)^{l(\sigma)}F^{i_{d}}_{\sigma(d),d}\ldots F^{i_{1}}_{\sigma(1),1}.$
Here the last equality holds since $\pi^{-1}\mathfrak{S}_{d}\pi=\mathfrak{S}_{d}$ and $l(\pi^{-1}\sigma\pi)=l(\sigma)$ for all $\sigma\in\mathfrak{S}_{d}$. Thus,
$\omega\big(\phi_{\bf z}{\rm cdet}({\mathcal{L}}_{d}^{\bf w})\big)=\sum_{\sigma\in\mathfrak{S}_{d}}\sum_{i_{1},\ldots,i_{d}=0}^{p+q+m+n}(-1)^{l(\sigma)}\omega\!\left(F^{i_{1}}_{\sigma(1),1}\right)\ldots\omega\!\left(F^{i_{d}}_{\sigma(d),d}\right).$
In view of (A.2), we deduce that $\omega\big(\phi_{\bf z}{\rm cdet}({\mathcal{L}}_{d}^{\bf w})\big)={\rm cdet}(\mathcal{C})$. ∎
Now we are ready to prove Proposition 4.7. For $1\leq r,s\leq p+q$ and $1\leq i,j\leq m+n$, we have
$\begin{array}[]{cll}\varphi_{\bf w}\big(E^{r}_{s}(z)\big)\hskip-28.45274pt&=\sum_{a=1}^{d}\frac{(-1)^{|s|+1}\partial_{y^{a}_{r}}y^{a}_{s}}{z-w_{a}},\quad\varphi_{\bf w}\big(E^{r}_{p+q+j}(z)\big)\hskip-8.5359pt&=\sum_{a=1}^{d}\frac{\partial_{y^{a}_{r}}\partial_{x^{a}_{j}}}{z-w_{a}},\\ \varphi_{\bf w}\big(E^{p+q+i}_{s}(z)\big)\hskip-8.5359pt&=\sum_{a=1}^{d}\frac{(-1)^{|s|+1}x^{a}_{i}y^{a}_{s}}{z-w_{a}},\quad\varphi_{\bf w}\big(E^{p+q+i}_{p+q+j}(z)\big)\hskip-8.5359pt&=\sum_{a=1}^{d}\frac{x^{a}_{i}\partial_{x^{a}_{j}}}{z-w_{a}}.\end{array}$} (A.3)
Let
$D=\Big[\delta_{a,b}(z-w_{a})\Big]_{a,b=1,\ldots,d}\qquad\text{and}\qquad D^{\prime}=\Big[\delta_{i,j}(\partial_{z}-z_{i})\Big]_{i,j\in\mathbb{I}}.$
Consider the $(m+n)\times d$ matrices
$X:=\big[(-1)^{|p+q+i|}x^{a}_{i}\big]^{a=1,\ldots,d}_{i=1,\ldots,m+n}\qquad\text{and}\qquad P_{X}:=\big[\partial_{x^{a}_{i}}\big]^{a=1,\ldots,d}_{i=1,\ldots,m+n}$
and the $(p+q)\times d$ matrices
$Y:=\big[(-1)^{|r|+1}y^{a}_{r}\big]^{a=1,\ldots,d}_{r=1,\ldots,p+q}\qquad\text{and}\qquad P_{Y}:=\big[(-1)^{|r|}\partial_{y^{a}_{r}}\big]^{a=1,\ldots,d}_{r=1,\ldots,p+q}.$
Write
$T=\begin{bmatrix}Y^{t}\hskip 5.69054ptP_{X}^{t}\end{bmatrix}\qquad\text{and}\qquad T^{\prime}=\begin{bmatrix}P_{Y}\vskip 8.53581pt\\ X\end{bmatrix}$.
By (A.3), we get
$\varphi_{\bf w}{\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big)={\rm Ber}^{\bf s}(D^{\prime}-T^{\prime}D^{-1}T).$
Let
$\mathfrak{L}=\begin{bmatrix}D&T\vskip 5.69054pt\\ T^{\prime}&D^{\prime}\end{bmatrix}$.
It is straightforward to check that $\mathfrak{L}$ is an amply invertible Manin matrix of type $\hat{{\bf s}}:=(0^{d},{\bf s})\in\mathcal{S}_{d+p+m|q+n}$. By Proposition 4.1, we have
${\rm Ber}^{\hat{{\bf s}}}(\mathfrak{L})=(z-w_{1})\ldots(z-w_{d})\,\varphi_{\bf w}{\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big).$
As $D-T(D^{\prime})^{-1}T^{\prime}$ is a sufficiently invertible Manin matrix of type $(0^{d})$, Proposition 4.4 and Proposition 4.5 imply that
${\rm Ber}^{\hat{{\bf s}}}(\mathfrak{L})$ $=$ ${\rm Ber}^{\bf s}(D^{\prime})\cdot{\rm cdet}(D-T(D^{\prime})^{-1}T^{\prime})$
$=$ $R_{p|q}(\partial_{z})\cdot R^{\prime}_{m|n}(\partial_{z})\cdot{\rm cdet}(\mathcal{C}^{\prime}),$
where
$\mathcal{C}^{\prime}:=\left[\delta_{a,b}(z-w_{a})+\sum_{r=1}^{p+q}\frac{y^{a}_{r}\partial_{y^{b}_{r}}}{\partial_{z}-z_{r}}-\sum_{i=1}^{m+n}\frac{(-1)^{|p+q+i|}\partial_{x^{a}_{i}}x^{b}_{i}}{\partial_{z}-z_{p+q+i}}\right]_{a,b=1,\ldots,d}.$
We claim that
$R^{\prime}_{m|n}(\partial_{z})\cdot{\rm cdet}(\mathcal{C}^{\prime})={\rm cdet}(\mathcal{C})\cdot R^{\prime}_{m|n}(\partial_{z}).$
This yields
${\rm Ber}^{\hat{{\bf s}}}(\mathfrak{L})=R_{p|q}(\partial_{z})\cdot{\rm cdet}(\mathcal{C})\cdot R^{\prime}_{m|n}(\partial_{z}),$
which proves Proposition 4.7 by Lemma A.1.
It remains to prove the claim. We only need to show that
$R^{\prime}_{m|n}(\partial_{z})\cdot\mathcal{C}^{\prime}_{a,b}=\mathcal{C}_{a,b}\cdot R^{\prime}_{m|n}(\partial_{z}),\qquad\mbox{for $1\leq a,b\leq d$.}$ (A.4)
Here, for instance, $\mathcal{C}_{a,b}$ denotes the $(a,b)$ entry of $\mathcal{C}$. Since
$\partial_{x^{a}_{i}}x^{b}_{i}=(-1)^{|p+q+i|}x^{b}_{i}\partial_{x^{a}_{i}}+\delta_{a,b},\qquad\mbox{for $1\leq a,b\leq d$},$ (A.5)
the off-diagonal entries of $\mathcal{C}^{\prime}$ coincide with those of $\mathcal{C}$. For $1\leq i\leq m+n$, the term $\partial_{z}-z_{p+q+i}$ commutes with $\partial_{x^{a}_{j}}x^{b}_{j}$ and $y^{a}_{r}\partial_{y^{b}_{r}}$ for all $1\leq j\leq m+n$, $1\leq r\leq p+q$, and $1\leq a,b\leq d$. In particular, $R^{\prime}_{m|n}(\partial_{z})$ commutes with any off-diagonal entry of $\mathcal{C}^{\prime}$. This proves (A.4) for $a\not=b$. Moreover, for $1\leq i\leq m+n$ and $1\leq a\leq d$,
$(\partial_{z}-z_{p+q+i})(z-w_{a})=(z-w_{a})(\partial_{z}-z_{p+q+i})+1.$ (A.6)
If $1\leq i\leq m$, we see from (A.5) and (A.6) that
${(\partial_{z}-z_{p+q+i})}\left(z-w_{a}-\frac{\partial_{x^{a}_{i}}x^{a}_{i}}{\partial_{z}-z_{p+q+i}}\right)=\left(z-w_{a}-\frac{x^{a}_{i}\partial_{x^{a}_{i}}}{\partial_{z}-z_{p+q+i}}\right)(\partial_{z}-z_{p+q+i}).$
Similarly, if $m+1\leq i\leq m+n$, we have
${(\partial_{z}-z_{p+q+i})^{-1}}\left(z-w_{a}+\frac{\partial_{x^{a}_{i}}x^{a}_{i}}{\partial_{z}-z_{p+q+i}}\right)=\left(z-w_{a}-\frac{x^{a}_{i}\partial_{x^{a}_{i}}}{\partial_{z}-z_{p+q+i}}\right)(\partial_{z}-z_{p+q+i})^{-1}.$
The above discussion proves (A.4) for $a=b$. Thus, the proof of (A.4) is now complete.